### v4

# Cointegration with Occasionally Binding Constraints

Abstract
In the literature on nonlinear cointegration, a long-standing open problem relates to how a (nonlinear) vector autoregression, which provides a unified description of the short- and long-run dynamics of a vector of time series, can generate ‘nonlinear cointegration’ in the profound sense of those series sharing common nonlinear stochastic trends. We consider this problem in the setting of the censored and kinked structural VAR (CKSVAR), which provides a flexible yet tractable framework within which to model time series that are subject to threshold-type nonlinearities, such as those arising due to occasionally binding constraints, of which the zero lower bound (ZLB) on short-term nominal interest rates provides a leading example. We provide a complete characterisation of how common linear and nonlinear stochastic trends may be generated in this model, via unit roots and appropriate generalisations of the usual rank conditions, providing the first extension to date of the Granger–Johansen representation theorem to a nonlinearly cointegrated setting, and thereby giving the first successful treatment of the open problem. The limiting common trend processes include regulated, censored and kinked Brownian motions, none of which have previously appeared in the literature on cointegrated VARs. Our results and running examples illustrate that the CKSVAR is capable of supporting a far richer variety of long-run behaviour than is a linear VAR, in ways that may be particularly useful for the identification of structural parameters.
We thank B. Beare, P. Bonomolo, H. P. Boswijk, A. Bykhovskaya, G. Cavaliere, R. Engle, J. Gao, D. Harris, X. Jiao, S. Johansen, I. Kasparis, D. Kristensen, M. Kulish, O. Lieberman, O. Linton, Y. Lu, T. Magdalinos, J. Morley, U. Müller, L. Neri, B. Nielsen, A. Onatski, M. Plagborg-Møller, W.-K. Seo, A. Srakar, J. Stock, A. M. R. Taylor, and participants in seminars at Aarhus, BI Oslo, Cambridge, Cyprus, Melbourne, Monash, Oxford, St. Andrews, Stanford, Sydney, the VTSS workshop, the May 2023 Harvard conference in honour of Jim Stock and Mark Watson, and the 24th Zaragoza Workshop in Time Series Econometrics, for helpful comments on earlier drafts of this work. This research was supported by the European Research Council via Consolidator Grant 647152.
1 Introduction
Nonstationarity, in the form of highly persistent, randomly wandering time series, is ubiquitous in macroeconomics and finance. It presents both a challenge for inference (Stock, 1994; Watson, 1994) and an opportunity for the identification of dynamic causal effects (Blanchard and Quah, 1989). The canonical framework for modelling such series is as (common) stochastic trends generated by a linear vector autoregression (VAR) with unit roots, underpinned by the powerful Granger–Johansen representation theorem (Johansen, 1995, Ch. 4). However, this framework is inadequate when even one of the variables is subject to an occasionally binding constraint, such as the zero lower bound (ZLB) constraint on short-term nominal interest rates, which has recently gained particular prominence in macroeconomic policy analysis (e.g. Summers, 2014; Williams, 2014; Eggertsson et al., 2019; Kocherlakota, 2019).
In an autoregressive model, an occasionally binding constraint naturally gives rise to nonlinearity in the form of multiple endogenously switching regimes, whose presence significantly complicates the links between unit roots and stochastic trends. While it has become increasingly common to introduce stochastic trends into (linear) empirical models of monetary policy by treating the ‘natural rate of interest’ or ‘trend inflation’ as latent random walks (Laubach and Williams, 2003; Cogley and Sbordone, 2008; Del Negro et al., 2017; Andrade et al., 2019; Bauer and Rudebusch, 2020; Schmitt-Grohé and Uribe, 2022), little is understood about what would happen to the implied time series properties of the observable series, such as the actual inflation rate and the nominal rate of interest, if nonlinearities were introduced into such models via the ZLB constraint. There is thus a pressing need to extend our understanding of how stochastic trends may be modelled beyond the linear VAR framework, to more general settings capable of accommodating such nonlinearities.
This paper addresses this problem in the setting of the censored and kinked structural VAR (CKSVAR) model (Mavroeidis, 2021; Aruoba et al., 2022), which provides a flexible yet tractable framework within which to model time series that are subject to occasionally binding constraints, and more generally to threshold-type nonlinearities. In this model, which is otherwise like a linear structural VAR, one series is allowed to enter differently according to whether it lies above or below a threshold; e.g. in applications to monetary policy, this series may be taken to represent the stance of monetary policy in each period, which above zero coincides with the short-term policy rate, and below zero corresponds to the (unobserved) ‘shadow rate’. We provide a complete characterisation of how common linear and nonlinear stochastic trends may be generated in this model, via unit roots and appropriate generalisations of the usual rank conditions, providing the first extension to date of the Granger–Johansen representation theorem to a nonlinearly cointegrated setting. Our results, which describe the behaviour of both the short- and long-run components of the CKSVAR, are foundational for frequentist inference in this setting, in the presence of common stochastic trends (a treatment of which will be given elsewhere).
As in a linear VAR, unit roots are unavoidable if one wishes to apply the CKSVAR to series with stochastic trends. The usual criticism of simply estimating a model in differences – that this obliterates the identifying long-run information carried by the cointegrating relations – is here magnified by the threshold nonlinearity in the model, which dictates whether the affected variable (e.g. interest rates) should enter in levels or differences, and in turn prescribes appropriate forms for the other variables. [^1] That these other variables (which enter the model linearly) cannot generally be replaced by their first differences can be seen from the vector error correction form of the CKSVAR given in (3.1) below, from which it is evident that making this replacement (of $x_{t}$ by $\Delta x_{t}$) would amount to imposing the restriction that $\Pi^{x}=0$. To put it another way, that nonlinearity prevents series from being simply ‘differenced to stationarity’. Moreover, we show that in our setting, unit roots are not a mere technical nuisance: rather, their presence may impart significant identifying power to the low frequency behaviour of the series. For instance, the possibility that cointegrating relations between series may change as one series crosses a threshold, something accommodated by the CKSVAR, may be utilised to test hypotheses about the relative effectiveness of unconventional monetary policy (see Examples 2.1 and 2.1 below; this is a problem that has been studied econometrically by e.g. Gambacorta et al., 2014; Wu and Xia, 2016; Debortoli et al., 2020; and Ikeda et al., 2024).
In analysing the CKSVAR with unit roots, we make a major contribution to the literature on nonlinear cointegration. Here a long-standing open problem relates to how a (nonlinear) vector autoregression, which provides a unified description of the short- and long-run dynamics of a collection of time series, can generate nonlinear cointegration between those series – where ‘nonlinear cointegration’ is understood in the profound sense of those series having common nonlinear stochastic trends with possibly nonlinear cointegrating relations between those trends. As discussed in the recent review by Tjøstheim (2020), despite the voluminous literature on the subject of ‘nonlinear cointegration’, this problem has yet to be addressed at any reasonable level of generality. [^2] While Cai et al. (2017) make an important effort in this direction, their results are limited to a first-order bivariate VAR with two regimes, in which one of those regimes is delimited by a compact set, and so makes a negligible contribution to the long-run behaviour of the series generated by the model. Their results are thus markedly different from those obtained below. Within the framework of the CKSVAR, we provide a resolution of this problem, showing that the model naturally gives rise to nonlinear cointegration, generating nonlinear common trend processes – censored, regulated, and kinked Brownian motions (see Definition 3.3 below) – that have not previously appeared in multivariate settings.
To clarify how our work relates to the existing literature on ‘nonlinear cointegration’, and to explain why we have been able to make progress in an area that has previously seemed intractable, we briefly recall the two main strands of that literature. One strand (Tjøstheim, 2020, p. 657) starts from the vector error correction model (VECM) representation of a cointegrated VAR, and introduces nonlinearity into the error correction mechanism; a prototypical model is
$\Delta z_{t}=g(\beta^{\mathsf{T}}z_{t-1})+\sum_{i=1}^{k-1}\Gamma_{i}\Delta z_{t-i}+u_{t},$ (1.1)
in which the usually linear loadings $\alpha[\beta^{\mathsf{T}}z_{t-1}]=\alpha\xi_{t-1}$ on the equilibrium errors are replaced by a general nonlinear function. In the original ‘threshold cointegration’ conception of this model, due to Balke and Fomby (1997), $g$ is piecewise linear, i.e. $g(\xi_{t})=\sum_{i=1}^{m}\alpha^{(i)}\mathbf{1}\{\xi_{t}\in\Xi_{i}\}\xi_{t}$, where each of the $\alpha^{(i)}$’s correspond to different ‘regimes’. The values of $\{\Gamma_{i}\}$ may also depend on lags of $\xi_{t}$ or $\Delta z_{t}$. (For regime-switching versions, including of the smoothed variety, see e.g. Hansen and Seo, 2002; Saikkonen, 2005, 2008; and Seo, 2011; for versions in which $g$ is allowed to be a more general nonlinear function, but the $\{\Gamma_{i}\}$ matrices are fixed, see Escribano and Mira, 2002; and Kristensen and Rahbek, 2010, 2013). Notably, the nonlinearity in such models is wholly confined to the short-run dynamics: as in a linearly cointegrated VAR, there remains a globally defined cointegrating space spanned by the columns of $\beta$ (i.e. which is common to all ‘regimes’), and the limiting common trends remain a (vector) Brownian motion.
The other strand (Tjøstheim, 2020, pp. 658–666) takes as its starting point the triangular representation of a linearly cointegrated system, and introduces nonlinearity directly into the common trends, by specifying
$y_{t}$ $=f(x_{t})+\varepsilon_{yt},$ $x_{t}$ $=x_{t-1}+\varepsilon_{xt}.$ (1.2)
Here $f(x_{t})$ replaces what would ordinarily be a linear function, with the consequence that the weak limit of $Y_{n}(\lambda)\coloneqq n^{-1/2}y_{\lfloor n\lambda\rfloor}$ will now be a nonlinear transformation of the limiting Brownian motions associated with $X_{n}(\lambda)\coloneqq n^{-1/2}x_{\lfloor n\lambda\rfloor}$. The errors $\varepsilon_{t}=(\varepsilon_{yt},\varepsilon_{xt})$ may be weakly dependent and cross-correlated, permitting $\{x_{t}\}$ to be endogenous. The function $f$ is typically estimated via some sort of regression, either parametrically (Park and Phillips, 1999, 2001; Chan and Wang, 2015; Li et al., 2016) or nonparametrically (Karlsen et al., 2007; Wang and Phillips, 2009, 2016; Duffy, 2017; Duffy and Kasparis, 2021); there is also literature on specification testing in this setting (e.g. Wang and Phillips, 2012; Dong et al., 2017; Wang et al., 2018; Berenguer-Rico and Nielsen, 2020). Notable variants have used $f$ to model transitions between regimes with distinct linear cointegrating relations (Saikkonen and Choi, 2004; Gonzalo and Pitarakis, 2006), or allowed it to take the ‘functional coefficient’ form $\beta(w_{t})x_{t}$ (Xiao, 2009; Phillips and Wang, 2023).
In developing a VAR model that exhibits both nonlinearity in its short-run dynamics, as in (1.1), and in the implied (long-run) common trends, as in (1.2), this paper is the first to bridge the remarkably wide gulf that has existed between these two strands of the literature. The CKSVAR turns out to provide just enough flexibility to accommodate meaningful departures from linear cointegration, while retaining a sufficiently nice structure to be tractable. We show that depending on the rank conditions imposed on (submatrices of) the autoregressive polynomial evaluated at unity, the CKSVAR is capable of generating three distinct kinds of nonlinear cointegration, which we term: (i) regulated cointegration; (ii) kinked cointegration; and (iii) linear cointegration in a nonlinear VECM.
At a technical level, our contribution consists of identifying the alternative configurations of the model that give rise to cases (i)–(iii), which are essentially exhaustive of the possibilities here, and deriving analogues of the Granger–Johansen representation theorem in these three cases. [^3] Our analysis is exhaustive with respect to the possibilities for generating series that are integrated of order one (in a suitably extended sense of the term; see Definition 3.1) within the CKSVAR; higher orders of integration are not considered here. In analysing case (i), our work relates to that of Cavaliere (2005), Liu et al. (2011), Gao et al. (2013), and most closely to Bykhovskaya and Duffy (2024), all of whom obtain convergence to regulated Brownian motions in univariate models. Case (ii), even in the univariate setting, does not appear to have been considered by any previous literature. While cases (i) and (ii) describe phenomena that are entirely new to the literature, case (iii) holds under a configuration of the model that falls within the very broad class of nonlinear VECMs considered by Saikkonen (2008): but even here, our results regarding the ergodicity of the short-memory components of the model extend his, insofar as we are able to exploit certain properties of the CKSVAR that are not shared by all the models encompassed by his general framework.
We illustrate the economic significance of cases (i) and (ii) by demonstrating how each may arise in our running example of a stylised structural model of monetary policy in the presence of a zero lower bound, contingent on the values taken by certain model parameters. We further apply our results to determine the long-run properties of the structural model of Ascari et al. (2023), for which the presence of nonlinear transformations of potentially stochastically trending series (due to a ‘long-run Phillips curve’) precludes the application of any pre-existing version of the Granger–Johansen representation theorem.
The remainder of the paper is organised as follows. Section 2 introduces the CKSVAR model and the stylised structural macroeconomic models that we use as running examples. Section 3 develops the heuristics of the model with unit roots, outlining the tripartite classification noted above. The main results of this paper, which extend the Granger–Johansen representation theorem to the CKSVAR model, are given in Section 4. All proofs appear in the appendices.
Notation.
$e_{m,i}$ denotes the $i$th column of an $m\times m$ identity matrix; when $m$ is clear from the context, we write this simply as $e_{i}$. In a statement such as $f(a^{\pm},b^{\pm})=0$, the notation $\pm$’ signifies that both $f(a^{+},b^{+})=0$ and $f(a^{-},b^{-})=0$ hold; similarly, $a^{\pm}\in A$’ denotes that both $a^{+}$ and $a^{-}$ are elements of $A$. All limits are taken as $n\rightarrow\infty$ unless otherwise stated. $\overset{p}{\rightarrow}$ and $\rightsquigarrow$ respectively denote convergence in probability and in distribution (weak convergence). We write $X_{n}(\lambda)\rightsquigarrow X(\lambda)$ on $D_{\mathbb{R}^{m}}[0,1]$’ to denote that $\{X_{n}\}$ converges weakly to $X$, where these are considered as random elements of $D_{\mathbb{R}^{m}}[0,1]$, the space of cadlag functions $0,1]\rightarrow\mathbb{R}^{m}$, equipped with the uniform topology; we denote this as $D[0,1]$ whenever the value of $m$ is clear from the context. $\lVert\cdot\rVert$ denotes the Euclidean norm on $\mathbb{R}^{m}$, and the matrix norm that it induces. For $X$ a random vector and $p\geq 1$, $\lVert X\rVert_{p}\coloneqq(\mathbb{E}\lVert X\rVert^{p})^{1/p}$.
2 Model: the censored and kinked SVAR
We consider a VAR($k$) model in $p$ variables, in which one series, $y_{t}$, enters with coefficients that differ according to whether it is above or below a time-invariant threshold $b$, while the other p1p-1 series, collected in $x_{t}$, enter linearly. Defining
$y_{t}^{+}$ $\coloneqq\max\{y_{t},b\}$ $y_{t}^{-}$ $\coloneqq\min\{y_{t},b\},$ (2.1)
we specify that $(y_{t},x_{t})$ follow
$\phi_{0}^{+}y_{t}^{+}+\phi_{0}^{-}y_{t}^{-}+\Phi_{0}^{x}x_{t}=c+\sum_{i=1}^{k}[\phi_{i}^{+}y_{t-i}^{+}+\phi_{i}^{-}y_{t-i}^{-}+\Phi_{i}^{x}x_{t-i}]+u_{t}$ (2.2)
or, more compactly,
$\phi^{+}(L)y_{t}^{+}+\phi^{-}(L)y_{t}^{-}+\Phi^{x}(L)x_{t}=c+u_{t},$ (2.3)
where
$\phi^{\pm}(L)$ $\coloneqq\phi_{0}^{\pm}-\sum_{i=1}^{k}\phi_{i}^{\pm}L^{i}$ $\Phi^{x}(L)$ $\coloneqq\Phi_{0}^{x}-\sum_{i=1}^{k}\Phi_{i}^{x}L^{i},$
for $\phi_{i}^{\pm}\in\mathbb{R}^{p\times 1}$ and $\Phi_{i}^{x}\in\mathbb{R}^{p\times(p-1)}$, and $L$ denotes the lag operator. As in a linear SVAR, $\{u_{t}\}$ may be an i.i.d. sequence of mutually orthogonal structural shocks, but our results below also permit them to be cross-correlated or weakly dependent. Through an appropriate redefinition of $y_{t}$ and $c$, we may take $b=0$ without loss of generality, and will do so throughout the sequel. [^4] Defining $y_{b,t}\coloneqq y_{t}-b$, $y_{b,t}^{+}\coloneqq\max\{y_{b,t},0\}$, $y_{b,t}^{-}\coloneqq\min\{y_{b,t},0\}$ and $c_{b}\coloneqq c-[\phi^{+}(1)+\phi^{-}(1)]b$, we can rewrite (2.3) as $\phi^{+}(L)y_{b,t}^{+}+\phi^{-}(L)y_{b,t}^{-}+\Phi^{x}(L)x_{t}=c_{b}+u_{t}.$ In this case, $y_{t}^{+}$ and $y_{t}^{-}$ respectively equal the positive and negative parts of $y_{t}$, and $y_{t}=y_{t}^{+}+y_{t}^{-}$. (Throughout the following, the notation $a^{\pm}$’ connotes $a^{+}$ and $a^{-}$ as objects associated respectively with $y_{t}^{+}$ and $y_{t}^{-}$, or their lags. If we want to instead denote the positive and negative parts of some $a\in\mathbb{R}$, we shall do so by writing $a]_{+}\coloneqq\max\{a,0\}$ or $a]_{-}\coloneqq\min\{a,0\}$.)
Models of the form of (2.3) have previously been employed in the literature to account for the dynamic effects of censoring, occasionally binding constraints, and endogenous regime switching. Mavroeidis (2021) proposed exactly this model, which he termed the censored and kinked structural VAR (CKSVAR) model, to describe the operation of monetary policy during periods when a zero lower bound may bind on the policy rate: in our notation, $y_{t}$ corresponds to his ‘shadow rate’, expressing the central bank’s desired policy stance, and $y_{t}^{+}$ to the actual policy rate. Aruoba et al. (2022) considered a model in which one variable is subject to an occasionally binding constraint, which although in its initial formulation is somewhat more general, reduces to an instance of the CKSVAR once the conditions necessary for the model to have a unique solution (for all values of $u_{t}$) have been imposed (see their Proposition 1(i)). This version of their model – i.e. that in which the ‘private sector regression functions’ are piecewise linear and continuous – is thus accommodated by (2.3). [^5] See also Aruoba et al. (2021), for a DSGE model with an occasionally binding constraint, in which agents’ decision rules are approximated by functions with these properties.
To put some economic flesh on the bones of the representation theory developed in Sections 3 and 4 below, we here introduce the running example of a stylised structural model of monetary policy in the presence of a zero lower bound (ZLB). This model provides a simple, economically interpretable framework in which we may illustrate the various forms of novel long-run behaviour permitted by the CKSVAR, by considering alternate parametrisations of the model. Moreover, as discussed in Section 3 below, the model elucidates how that long-run behaviour may provide identifying information on the relative effectiveness of unconventional monetary policy (as compared with conventional rate-setting policy), i.e. on whether the zero lower bound really constrains the ability of a central bank to target inflation.
Example 2.1.
Consider the following stylised structural model, a simplified version of the model of Ikeda et al. (2024), consisting of a composite IS and Phillips curve (PC) equation
$\pi_{t}-\overline{\pi}_{t}$ $=\theta[i_{t}^{+}+\mu i_{t}^{-}-(r_{t}^{\ast}+\overline{\pi}_{t})]+\varepsilon_{t}$ (2.4)
and a policy reaction function (Taylor rule)
$i_{t}=(r_{t}^{\ast}+\overline{\pi}_{t})+\gamma(\pi_{t}-\overline{\pi}_{t}),$ (2.5)
where $r_{t}^{\ast}$ denotes the (real) natural rate of interest, $\overline{\pi}_{t}$ the central bank’s inflation target, $\pi_{t}$ inflation, and $\varepsilon_{t}$ a mean zero, i.i.d. innovation. $i_{t}$ measures the stance of monetary policy; thus $i_{t}^{+}\coloneqq[i_{t}]_{+}$ gives the actual policy rate (constrained to be non-negative), and $i_{t}^{-}\coloneqq[i_{t}]_{-}$ the desired stance of policy when the ZLB binds, to be effected via some form of ‘unconventional’ monetary policy, such as long-term asset purchases. We maintain that $\gamma>0$ and $\theta<0$. The parameter $\mu\in[0,1]$ reflects the relative efficacy of unconventional policy, with $\mu=1$ if this is as effective as conventional policy.
To ‘close’ the model, we consider two alternative specifications for the underlying processes followed by $\{r_{t}^{\ast}\}$ and $\{\overline{\pi}_{t}\}$. In the first of these (henceforth, Example 2.1a) the inflation target is assumed to be constant and is normalised to zero (i.e. $\overline{\pi}_{t}=\overline{\pi}=0$), while the natural real rate of interest follows a random walk AR(1) process (as in the model of Laubach and Williams, 2003),
$r_{t}^{\ast}=r_{t-1}^{\ast}+\eta_{t}$ (2.6)
where $\eta_{t}$ is an i.i.d. mean zero innovation, possibly correlated with $\varepsilon_{t}$. Substituting (2.5) into (2.4) and (2.6), we render the system as a CKSVAR for $(i_{t},\pi_{t})$ as
$\begin{bmatrix}1&1&-\gamma\\ 0&\theta(1-\mu)&1-\theta\gamma\end{bmatrix}\begin{bmatrix}i_{t}^{+}\\ i_{t}^{-}\\ \pi_{t}\end{bmatrix}=\begin{bmatrix}1&1&-\gamma\\ 0&0&0\end{bmatrix}\begin{bmatrix}i_{t-1}^{+}\\ i_{t-1}^{-}\\ \pi_{t-1}\end{bmatrix}+\begin{bmatrix}\eta_{t}\\ \varepsilon_{t}\end{bmatrix}$. (2.7)
This model will provide an illustration of the second kind of nonlinear cointegration developed in Section 3 below.
In the second variant of the model (henceforth, Example 2.1b) the natural rate is assumed to be constant and, for simplicity of exposition, normalised to zero (i.e. $r_{t}^{\ast}=r^{\ast}=0$), while the inflation target is allowed to be time-varying, according to
$\overline{\pi}_{t}=\overline{\pi}_{t-1}+\delta(\pi_{t-1}-\overline{\pi}_{t-1})+\eta_{t}$ (2.8)
where $\delta\in(-1,0]$, and $\eta_{t}$ is an i.i.d. innovation as above. When $\delta=0$, this corresponds to a model in which the inflation target follows a pure random walk, possibly reflecting the time-varying preferences of the central bank (cf. Cogley and Sbordone, 2008); when $\delta<0$, the model allows past deviations of inflation from target to feed back into the target, such that e.g. below-target inflation induces an upward revision of the inflation target. Motivation for this aspect of the model comes from the manner in which the ZLB may constrain policy to be excessively deflationary for a sustained period, something that has prompted the literature to consider the costs and benefits of adopting a higher inflation target (e.g. Blanchard et al., 2010, pp. 207f.; Coibion et al., 2012). Supposing additionally that $\gamma>1$, we may put (2.4), (2.5) and (2.8) in the form of a CKSVAR as
$\begin{bmatrix}-1&-1&\gamma\\ \varphi_{1}&\varphi_{\mu}&-\varphi_{1}\end{bmatrix}\begin{bmatrix}i_{t}^{+}\\ i_{t}^{-}\\ \pi_{t}\end{bmatrix}=\begin{bmatrix}\delta-1&\delta-1&\gamma-\delta\\ 0&0&0\end{bmatrix}\begin{bmatrix}i_{t-1}^{+}\\ i_{t-1}^{-}\\ \pi_{t-1}\end{bmatrix}+(\gamma-1)\begin{bmatrix}\eta_{t}\\ \varepsilon_{t}\end{bmatrix}$ (2.9)
where $\varphi_{\mu}\coloneqq(1-\mu\theta\gamma)-\theta(1-\mu)$ and so $\varphi_{1}=1-\theta\gamma$. Depending on the assumptions made on the model parameters (in particular $\delta$), this model is capable of generating either of the first two types of nonlinear cointegration discussed in Section 3.
While both Mavroeidis (2021) and Aruoba et al. (2022) motivate and interpret (2.3) as a structural model, empirically motivated reduced-form models of this kind have also appeared in the literature, particularly in the univariate ($p=1$) case of (2.3), which encompasses the dynamic Tobit model (Maddala 1983, p. 186; for applications, see e.g. Demiralp and Jordà, 2002; De Jong and Herrera, 2011; Brezigar-Masten et al., 2021; and Bykhovskaya, 2023).
Example 2.2 (univariate).
Consider (2.3) with $p=1$ and $\phi_{0}^{+}=\phi_{0}^{-}=1$, so that
$y_{t}=c+\sum_{i=1}^{k}(\phi_{i}^{+}y_{t-i}^{+}+\phi_{i}^{-}y_{t-i}^{-})+u_{t}.$ (2.10)
In the nomenclature of Bykhovskaya and Duffy (2024, Sec. 1), if $\phi_{i}^{-}=0$ for all $i\in\{1,\ldots,k\}$, so that only the positive part of $y_{t-i}$ enters the r.h.s., then
$y_{t}^{+}=\left[c+\sum_{i=1}^{k}\phi_{i}^{+}y_{t-i}^{+}+u_{t}\right]_{+}$ (2.11)
follows a ‘censored’ dynamic Tobit.
We follow Mavroeidis (2021) and Aruoba et al. (2022) in maintaining the following, which are necessary and sufficient to ensure that (2.3) has a unique solution for $(y_{t},x_{t})$, for all possible values of $u_{t}$. Define
$\Phi_{0}\coloneqq\begin{bmatrix}\phi_{0}^{+}&\phi_{0}^{-}&\Phi_{0}^{x}\end{bmatrix}=\begin{bmatrix}\phi_{0,yy}^{+}&\phi_{0,yy}^{-}&\phi_{0,yx}^{\mathsf{T}}\\ \phi_{0,xy}^{+}&\phi_{0,xy}^{-}&\Phi_{0,xx}\end{bmatrix}$,
$\Phi_{0}^{+}\coloneqq[\phi_{0}^{+},\Phi_{0}^{x}]$ and $\Phi_{0}^{-}\coloneqq[\phi_{0}^{-},\Phi_{0}^{x}]$.
Assumption DGP.
  1. 1.
    $\{(y_{t},x_{t})\}$ are generated according to (2.1)–(2.3) with $b=0$, with (possibly random) initial values $(y_{i},x_{i})$, for $i\in\{-k+1,\ldots,0\}$;
  2. 2.
    $\operatorname{sgn}(\det\Phi_{0}^{+})=\operatorname{sgn}(\det\Phi_{0}^{-})\neq 0$.
  3. 3.
    $\Phi_{0,xx}$ is invertible, and
    $\operatorname{sgn}\{\phi_{0,yy}^{+}-\phi_{0,yx}^{\mathsf{T}}\Phi_{0,xx}^{-1}\phi_{0,xy}^{+}\}=\operatorname{sgn}\{\phi_{0,yy}^{-}-\phi_{0,yx}^{\mathsf{T}}\Phi_{0,xx}^{-1}\phi_{0,xy}^{-}\}>0.$
For a further discussion of these conditions, including why DGP.3 may be maintained without loss of generality when DGP.2 holds, see Duffy et al. (2023, Sec. 2). As in that paper, we shall designate a CKSVAR as canonical if
$\Phi_{0}=\begin{bmatrix}1&1&0\\ 0&0&I_{p-1}\end{bmatrix}\eqqcolon I_{p}^{\ast}.$ (2.12)
While it is not always the case that the reduced form of (2.3) corresponds directly to a canonical CKSVAR, by defining the canonical variables
$\begin{bmatrix}\tilde{y}_{t}^{+}\\ \tilde{y}_{t}^{-}\\ \tilde{x}_{t}\end{bmatrix}\coloneqq\begin{bmatrix}\bar{\phi}_{0,yy}^{+}&0&0\\ 0&\bar{\phi}_{0,yy}^{-}&0\\ \phi_{0,xy}^{+}&\phi_{0,xy}^{-}&\Phi_{0,xx}\end{bmatrix}\begin{bmatrix}y_{t}^{+}\\ y_{t}^{-}\\ x_{t}\end{bmatrix}\eqqcolon P^{-1}\begin{bmatrix}y_{t}^{+}\\ y_{t}^{-}\\ x_{t}\end{bmatrix}$, (2.13)
where $\bar{\phi}_{0,yy}^{\pm}\coloneqq\phi_{0,yy}^{\pm}-\phi_{0,yx}^{\mathsf{T}}\Phi_{0,xx}^{-1}\phi_{0,xy}^{\pm}>0$ and $P^{-1}$ is invertible under DGP; and setting
$\begin{bmatrix}\tilde{\phi}^{+}(\lambda)&\tilde{\phi}^{-}(\lambda)&\tilde{\Phi}^{x}(\lambda)\end{bmatrix}\coloneqq Q\begin{bmatrix}\phi^{+}(\lambda)&\phi^{-}(\lambda)&\Phi^{x}(\lambda)\end{bmatrix}P,$ (2.14)
where
$Q\coloneqq\begin{bmatrix}1&-\phi_{0,yx}^{\mathsf{T}}\Phi_{0,xx}^{-1}\\ 0&I_{p-1}\end{bmatrix}$, (2.15)
we obtain a canonical CKSVAR for $(\tilde{y}_{t},\tilde{x}_{t})$. This is formalised by the following, which reproduces the first part of Proposition 2.1 in Duffy et al. (2023).
Proposition 2.1.
Suppose DGP holds. Then there exist $(\tilde{y}_{t},\tilde{x}_{t})$ such that (2.13)–(2.14) hold, $\tilde{y}_{t}^{+}=\max\{\tilde{y}_{t},0\}$, $\tilde{y}_{t}^{-}=\min\{\tilde{y}_{t},0\}$ and
$\tilde{\phi}^{+}(L)\tilde{y}_{t}^{+}+\tilde{\phi}^{-}(L)\tilde{y}_{t}^{-}+\tilde{\Phi}^{x}(L)\tilde{x}_{t}=\tilde{c}+\tilde{u}_{t},$ (2.16)
is a canonical CKSVAR, where $\tilde{c}=Qc$ and $\tilde{u}_{t}=Qu_{t}$.
To distinguish between a general CKSVAR in which possibly $\Phi_{0}\neq I_{p}^{\ast}$, and its associated canonical form as given by Proposition 2.1, we shall refer to the former as the ‘structural form’ of the CKSVAR. Since the time series properties of a general CKSVAR are largely inherited from its derived canonical form, we shall often work with this more convenient representation of the system, and indicate this as follows.
Assumption DGP.
$\{(y_{t},x_{t})\}$ are generated by a canonical CKSVAR, i.e. DGP holds with $\Phi_{0}=[\phi_{0}^{+},\phi_{0}^{-},\Phi^{x}]=I_{p}^{\ast}$, so that (2.2) may be equivalently written as
$\begin{bmatrix}y_{t}\\ x_{t}\end{bmatrix}=c+\sum_{i=1}^{k}\begin{bmatrix}\phi_{i}^{+}&\phi_{i}^{-}&\Phi_{i}^{x}\end{bmatrix}\begin{bmatrix}y_{t-i}^{+}\\ y_{t-i}^{-}\\ x_{t-i}\end{bmatrix}+u_{t}.$ (2.17)
3 Unit roots and nonlinear cointegration: heuristics
3.1 Nonlinearity and cointegration
It is well known that a linear VAR can faithfully replicate the high persistence, random wandering and long-run co-movement that is characteristic of a great many macroeconomic time series, via the imposition of unit autoregressive roots and the familiar rank conditions (Johansen, 1995). The question thus arises as to whether, and how, such behaviour may also be generated within a CKSVAR, so that the model might still be applied to series for which a stationary CKSVAR would be inappropriate. As we shall see, it is possible not only to accommodate linear cointegration within the CKSVAR, but also to generate a variety of nonlinear forms of cointegration, owing to the richer class of common trend processes that the model supports. Moreover, as our examples below illustrate, such departures from linear cointegration may also aid in the identification of structural parameters.
In developing the CKSVAR with unit roots, we shall find it necessary to depart from the usual classification of processes according to their orders of integration, since the nonlinearity in the model generally prevents it from generating series that are difference stationary. This is an issue commonly encountered in regime-switching cointegration models: see e.g. the discussion in Gonzalo and Pitarakis (2006, pp. 816f.), where this motivates the definition of the ‘order of summability’ of a time series, and the allied notion of ‘co-summability’ as a generalisation of linear cointegration (Berenguer-Rico and Gonzalo, 2014, pp. 335f.). While those concepts could well be applied to the CKSVAR, the following properties, which may be more easily verified, will suffice for our purposes.
Definition 3.1.
Let $\{w_{t}\}_{t\in\mathbb{N}}$ be a random sequence taking values in $\mathbb{R}^{p}$. We say that $\{w_{t}\}$ is:
  1. (i)
    $\ast$-stationary, denoted $w_{t}\sim I^{\ast}(0)$, if $\sup_{1\leq t\leq n}\lVert w_{t}\rVert=o_{p}(n^{1/2})$; or
  2. (ii)
    $\ast$-integrated (of order one), denoted $w_{t}\sim I^{\ast}(1)$, if $n^{-1/2}w_{\lfloor n\lambda\rfloor}\rightsquigarrow\ell(\lambda)$ on $D_{\mathbb{R}^{p}}[0,1]$, where $\ell$ is a non-degenerate stochastic process with continuous sample paths;
and analogously for subvectors (and individual elements) of $\{w_{t}\}$.
The preceding relates to the discussion of ‘cointegration’ in Müller and Watson (2013), who consider a model with linear cointegration in which the common trends belong to a broad class of processes satisfying a weak convergence criterion (see their eq. (2)), which includes strict $I(1)$ and local-to-unity models as special cases, so that the vector of time series generated by the model is $I^{\ast}(1)$ as defined above. The bounded unit root processes of Cavaliere (2005) also provide an example of a series that converges weakly upon standardisation by $n^{-1/2}$, and so are $I^{\ast}(1)$, but not $I(1)$.
Example 3.1 (univariate; ctd).
Within the CKSVAR framework, a simple, non-trivial example of series that are $I^{\ast}(d)$ but not $I(d)$ is provided by the censored dynamic Tobit. When $\sum_{i=1}^{k}\phi_{i}^{+}=1$, this model has a unit root, and Bykhovskaya and Duffy (2024) show that $n^{-1/2}y_{\lfloor n\lambda\rfloor}^{+}\rightsquigarrow Y^{+}(\lambda)$ on $D[0,1]$, where $Y^{+}$ is a Brownian motion regulated (at zero) from below (their Theorem 3.2; see Definition 3.3 below), and $\lVert\Delta y_{t}^{+}\rVert_{2+\delta_{u}}$ is uniformly bounded (their Lemma B.2). Thus $\Delta y_{t}^{+}\sim I^{\ast}(0)$ and $y_{t}^{+}\sim I^{\ast}(1)$, even though, due to the nonlinearity in the model, neither series are $I(d)$ for any $d$.
Here we also need an enlarged notion of ‘cointegration’ that is sufficiently general to encompass the possibilities of nonlinear cointegration accommodated by the CKSVAR model, particularly for the analysis of ‘case (ii)’ below, such as is provided by the following (cf. Gonzalo and Pitarakis, 2006, p. 817).
Definition 3.2.
Let $\mathscr{Z}^{+}\coloneqq\{(y,x)\in\mathbb{R}^{p}\mid y\geq 0\}$ and $\mathscr{Z}^{-}\coloneqq\{(y,x)\in\mathbb{R}^{p}\mid y\leq 0\}$, $r^{\pm}\in\{0,\ldots,p-1\}$, and $\beta^{\pm}\in\mathbb{R}^{p\times r^{\pm}}$ have full column rank. Suppose $z_{t}=(y_{t},x_{t}^{\mathsf{T}})^{\mathsf{T}}\sim I^{\ast}(1)$, but
$\mathbf{1}\{z_{t}\in\mathscr{Z}^{(i)}\}\theta^{\mathsf{T}}z_{t}\sim I^{\ast}(0)\iff\theta\in\operatorname{sp}\beta^{(i)}$
for $(i)\in\{+,-\}$. Then $z_{t}$ is said to be cointegrated on $\mathscr{Z}^{(i)}$, with $r^{(i)}$ the cointegrating rank on $\mathscr{Z}^{(i)}$, $\operatorname{sp}\beta^{(i)}$ the cointegrating space on $\mathscr{Z}^{(i)}$, and any (nonzero) element of $\operatorname{sp}\beta^{(i)}$ a cointegrating vector on $\mathscr{Z}^{(i)}$, for $(i)\in\{+,-\}$. If $\beta^{(i)}$ does not depend on $(i)$, we drop the ‘on $\mathscr{Z}^{(i)}$’ qualifiers.
3.2 The CKSVAR with unit roots
Our next step is to rewrite the CKSVAR, as in (2.2) or (2.3) above, in the form of a vector error-correction model (VECM). Define the autoregressive polynomials
$\Phi^{\pm}(\lambda)\coloneqq\begin{bmatrix}\phi^{\pm}(\lambda)&\Phi^{x}(\lambda)\end{bmatrix}$,
and let $\Gamma_{i}^{\pm}\coloneqq-\sum_{j=i+1}^{k}\Phi_{j}^{\pm}\eqqcolon[\gamma_{i}^{\pm},\Gamma^{x}]$ for $i\in\{1,\ldots,k-1\}$, so that $\Gamma^{\pm}(\lambda)\coloneqq\Phi_{0}^{\pm}-\sum_{i=1}^{k-1}\Gamma_{i}^{\pm}\lambda^{i}$ is such that
$\Phi^{\pm}(\lambda)=\Phi^{\pm}(1)\lambda+\Gamma^{\pm}(\lambda)(1-\lambda).$
Set $\pi^{\pm}\coloneqq-\phi^{\pm}(1)$ and $\Pi^{x}\coloneqq-\Phi^{x}(1)$. Then
$\Phi_{0}\begin{bmatrix}\Delta y_{t}^{+}\\ \Delta y_{t}^{-}\\ \Delta x_{t}\end{bmatrix}=c+\begin{bmatrix}\pi^{+}&\pi^{-}&\Pi^{x}\end{bmatrix}\begin{bmatrix}y_{t-1}^{+}\\ y_{t-1}^{-}\\ x_{t-1}\end{bmatrix}+\sum_{i=1}^{k-1}\begin{bmatrix}\gamma_{i}^{+}&\gamma_{i}^{-}&\Gamma_{i}^{x}\end{bmatrix}\begin{bmatrix}\Delta y_{t-i}^{+}\\ \Delta y_{t-i}^{-}\\ \Delta x_{t-i}\end{bmatrix}+u_{t},$ (3.1)
where $\Delta\coloneqq 1-L$ denotes the difference operator, and for clarity we note that $\Delta y_{t}^{+}=y_{t}^{+}-y_{t-1}^{+}$ (rather than being the positive part of $\Delta y_{t}$). In the case of a canonical CKSVAR, (3.1) helpfully reduces to
$\begin{bmatrix}\Delta y_{t}\\ \Delta x_{t}\end{bmatrix}=c+\begin{bmatrix}\pi^{+}&\pi^{-}&\Pi^{x}\end{bmatrix}\begin{bmatrix}y_{t-1}^{+}\\ y_{t-1}^{-}\\ x_{t-1}\end{bmatrix}+\sum_{i=1}^{k-1}\begin{bmatrix}\gamma_{i}^{+}&\gamma_{i}^{-}&\Gamma_{i}^{x}\end{bmatrix}\begin{bmatrix}\Delta y_{t-i}^{+}\\ \Delta y_{t-i}^{-}\\ \Delta x_{t-i}\end{bmatrix}+u_{t}.$ (3.2)
While our main results apply to the general CKSVAR, the reader may find it helpful to work through the remainder of this section under the supposition that $(y_{t},x_{t})$ are generated by a canonical CKSVAR.
Just as in a linear (cointegrated) VAR, which corresponds to the special case of (3.2) in which $\pi^{+}=\pi^{-}$ and $\gamma_{i}^{+}=\gamma_{i}^{-}$ for all $i\in\{1,\ldots,k-1\}$, the long-run dynamics will be governed by the matrix of coefficients on the lagged levels. More precisely, there are two such $p\times p$ matrices, $\Pi^{+}$ and $\Pi^{-}$, defined by
$\Pi^{\pm}\coloneqq[\pi^{\pm},\Pi^{x}]=-\Phi^{\pm}(1).$ (3.3)
Although the canonical CKSVAR technically has $2^{k}$ distinct autoregressive ‘regimes’ (corresponding to the possible sign patterns of $\boldsymbol{y}_{t-1}\coloneqq(y_{t-1},\ldots,y_{t-k})^{\mathsf{T}}$; see also Duffy et al., 2023, Sec. 4.2), the behaviour of the CKSVAR with unit roots depends largely on the two regimes in which the elements of $\boldsymbol{y}_{t-1}$ are all either positive or negative, which we shall loosely refer to as the ‘positive’ and ‘negative’ regimes, to which $\Pi^{+}$ and $\Pi^{-}$ correspond. This simplification occurs because whenever $y_{t}\sim I^{\ast}(1)$, it spends most of its time away from the origin, so that all elements of $\boldsymbol{y}_{t-1}$ will have the same sign almost all of the time.
Our baseline assumptions on the CKSVAR with unit roots may now be stated.
Assumption CVAR.
  1. 1.
    $\det\Phi^{\pm}(\lambda)$ has $q^{\pm}\in\{1,\ldots,p\}$ unit roots, and all others outside the unit circle; and
  2. 2.
    $\operatorname{rk}\Pi^{\pm}=r^{\pm}=p-q^{\pm}$.
Assumption DET.
$c\in\operatorname{sp}\Pi^{+}\cap\operatorname{sp}\Pi^{-}$.
Assumption ERR.
$\{u_{t}\}_{t\in\mathbb{N}}$ is a random sequence in $\mathbb{R}^{p}$, such that $\sup_{t\in\mathbb{N}}\lVert u_{t}\rVert_{2+\delta_{u}}<\infty$ for some $\delta_{u}>0$, and
$U_{n}(\lambda)\coloneqq n^{-1/2}\sum_{t=1}^{\lfloor n\lambda\rfloor}u_{t}\rightsquigarrow U(\lambda)$ (3.4)
on $D[0,1]$, where $U$ is a Brownian motion in $\mathbb{R}^{p}$ with (positive definite) variance $\Sigma$.
Remark 3.1.
By the functional martingale central limit theorem, under the stated moment condition on $\{u_{t}\}$, a sufficient but not necessary condition for (3.4) is that $\{u_{t}\}$ be a stationary and ergodic martingale difference sequence with $\Sigma\coloneqq\mathbb{E}u_{t}u_{t}^{\mathsf{T}}$. However, ERR also allows $\{u_{t}\}$ to be weakly dependent, such as if e.g. $u_{t}=\sum_{i=0}^{\infty}\Theta_{i}\eta_{t-i}$, where $\sum_{i=0}^{\infty}\lVert\Theta_{i}\rVert<\infty$ and $\{\eta_{t}\}$ is a mean zero, i.i.d. process with $\lVert\eta_{t}\rVert_{2+\delta_{u}}<\infty$.
In a linear VAR (i.e. where $\Phi^{+}(\lambda)=\Phi^{-}(\lambda)$) under CVAR the Granger–Johansen representation theorem (Johansen, 1995, Thm. 4.2 and Cor. 4.3) implies that $z_{t}\coloneqq(y_{t},x_{t}^{\mathsf{T}})^{\mathsf{T}}\sim I^{\ast}(1)$ and that there exists a full column rank matrix $\beta\in\mathbb{R}^{p\times r}$ such that $\beta^{\mathsf{T}}z_{t}\sim I^{\ast}(0)$, where $r=r^{+}=r^{-}$. Thus $\{z_{t}\}$ is cointegrated in the sense of Definition 3.2, with cointegrating space $\operatorname{sp}\beta$. Here DET specialises to $c\in\operatorname{sp}\Phi(1)$, preventing the model from generating any common deterministic trends between the series; it has the same effect more generally when $\Pi^{+}\neq\Pi^{-}$. (We shall relax this condition, so as to allow for deterministic trends, in Section 4.4 below.)
While linear cointegration may occur in a CKSVAR, other phenomena are possible, depending on the ranks of $\Pi^{+}$, $\Pi^{-}$ and $\Pi^{x}$. Within the framework of $I^{\ast}(0)$ and $I^{\ast}(1)$ processes, as delimited by CVAR, there are three possibilities, each of which generate profoundly different trajectories for $\{y_{t}\}$. These are characterised in Table 3.1; in each case $r$ (without a superscript) is defined such that it is possible to regard the system as having a cointegrating rank of at least $r$ in each ‘regime’. Since, $\Pi^{+}$ and $\Pi^{-}$ differ by only their first column, $r^{+}$ and $r^{-}$ may differ by at most one, so that the case where $r^{+}\neq r^{-}$ may be identified with case (i) in the table without loss of generality. [^6] If we were to instead take $r^{+}=r^{-}+1$ in this case, the characteristic behaviour of the series in the positive and negative regimes, with $y_{t}^{+}\sim I^{\ast}(1)$ and $y_{t}^{-}\sim I^{\ast}(0)$, would now be reversed, but the key properties of the model would otherwise be unaltered.
Case $^{+} rr^{-}$ $\operatorname{rk}\Pi^{x}$ $\pi^{+}\in\operatorname{sp}\Pi^{x}$ $\pi^{-}\in\operatorname{sp}\Pi^{x}$
(i) $r$ $r+1$ $r$ yes no
(ii) $r$ $r$ $r$ yes yes
(iii) $r$ $r$ r1r-1 no no
Table 3.1: Possible configurations of $\Pi^{\pm}=[\pi^{\pm},\Pi^{x}]$ and $r^{\pm}=\operatorname{rk}\Pi^{\pm}$
3.3 Common trends and long-run trajectories
To develop some intuition for the properties of the model in these three cases, in advance of the representation theory developed in the next section, it is helpful to regard the long-run behaviour of the processes as being characterised by a space of common trends $\mathscr{M}$, defined as the set of non-stochastic (i.e. $u_{t}\coloneqq 0$, $\forall t$) steady state solutions to (3.1):
$\mathscr{M}$ $\coloneqq\{(y,x)\in\mathbb{R}^{p}\mid\pi^{+}[y]_{+}+\pi^{-}[y]_{-}+\Pi^{x}x=0\}$
$=\{(y,x)\in\mathscr{Z}^{+}\mid\pi^{+}y+\Pi^{x}x=0\}\cup\{(y,x)\in\mathscr{Z}^{-}\mid\pi^{-}y+\Pi^{x}x=0\}$
$\eqqcolon\mathscr{M}^{+}\cup\mathscr{M}^{-}.$
This space defines the domain of the limiting processes
$Z_{n}(\lambda)\coloneqq\begin{bmatrix}Y_{n}(\lambda)\\ X_{n}(\lambda)\end{bmatrix}\coloneqq\frac{1}{n^{1/2}}\begin{bmatrix}y_{\lfloor n\lambda\rfloor}\\ x_{\lfloor n\lambda\rfloor}\end{bmatrix}\rightsquigarrow\begin{bmatrix}Y(\lambda)\\ X(\lambda)\end{bmatrix}\eqqcolon Z(\lambda).$ (3.5)
In a linear VAR, $\mathscr{M}=\ker\Pi$ is a $q$-dimensional linear subspace of $\mathbb{R}^{p}$, the orthogonal complement of which is the cointegrating space; and $Z$ is a $p$-dimensional Brownian motion with rank $q$ covariance matrix – i.e. it is a rank $q$ linear function of $U$ in (3.4) above – taking values in $\mathscr{M}$. Whereas in the CKSVAR, $\mathscr{M}$ no longer need be a linear subspace, but is instead a linear cone [^7] Recall that a set $S\subset\mathbb{R}^{p}$ is termed a linear cone if $\lambda s\in S$ for every $s\in S$ and $\lambda\geq 0$, i.e. if $S$ is closed under multiplication by non-negative scalars. formed from the union of $\mathscr{M}^{+}$ and $\mathscr{M}^{-}$, the vectors orthogonal to which are the cointegrating vectors $\beta^{+}$ and $\beta^{-}$ on the half spaces $\mathscr{Z}^{+}$ and $\mathscr{Z}^{-}$ respectively (recall Definition 3.2 above). While $Z$ remains a function of $U$, that function need not be linear: indeed, in addition to (linear) Brownian motions (BMs), any of the following nonlinear processes may also appear among the limiting stochastic trends generated by the CKSVAR. (For further discussion of regulated BMs, see Harrison, 1985, Ch. 1.)
Definition 3.3.
Let $W$ be a linear BM initialised from some $W(0)\in\mathbb{R}^{p}$. Suppose $p=1$; the scalar process $V$ is said to be a
  1. (i)
    censored BM (from below), if $V(\lambda)=\max\{W(\lambda),0\}$
  2. (ii)
    regulated BM (from below), if $V(\lambda)=W(\lambda)+\sup_{\lambda^{\prime}\leq\lambda}[-W(\lambda^{\prime})]_{+}$.
If $V$ is as in (i) or (ii), then V-V is respectively censored or regulated from above. Suppose now that $p\geq 1$, and let $G:\mathbb{R}\rightarrow\mathbb{R}^{p\times p}$ be a map that depends only on the sign of its argument, and is such that: (a) $h\coloneqq e_{1}^{\mathsf{T}}G(+1)=\mu e_{1}^{\mathsf{T}}G(-1)$ for some $\mu>0$; and (b) $w\mapsto G(h^{\mathsf{T}}w)w$ is continuous. Then the $p$-dimensional process $V$ is said to be a
  1. (iii)
    kinked BM, if $V(\lambda)=G[h^{\mathsf{T}}W(\lambda)]W(\lambda)=G[V_{1}(\lambda)]W(\lambda).$
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 3.1: Linear Brownian motion and derived nonlinear processes
Remark 3.2.
Trajectories of these processes (denoted by $V$) in the univariate case ($p=1$), together with the realisation of the standard Brownian motion $W$ used to construct them, are plotted in Figure 3.1. While both censored and regulated BMs are constrained to be positive, it is evident from panels (b) and (c) that there are important differences between them. For the former, the censoring does not feed back into the underlying dynamics, and $V$ spends long stretches at zero (while $W$ is negative); whereas for the latter, the $\sup_{\lambda^{\prime}\leq\lambda}[-W(\lambda^{\prime})]_{+}$ term continually reflects $V$ away from zero, so that $V$ spends relatively little time near zero. The kinked BM in panel (d) is constructed as
$V(\lambda)=\sigma_{-}W(\lambda)\mathbf{1}\{W(\lambda)<0\}+\sigma_{+}W(\lambda)\mathbf{1}\{W(\lambda)\geq 0\}$
with $\sigma_{-}=1$ and $\sigma_{+}=2$; hence it tracks $W$ exactly when $W(\lambda)<0$, but doubles the scale of $W$ when $W(\lambda)>0$. In other respects, in the univariate case, the trajectory of a kinked BM more closely resembles that of a linear BM than do either censored or regulated BMs. In the multivariate case, kinked BMs are linear combinations of the elements of $W$, with weights that depend on the sign of $V_{1}$. Thus when plotted individually they appear similarly to panel (d), but if $G(\pm 1)$ is rank deficient, then there will also be certain linear combinations of the elements of $V$ that will be zero, with those combinations depending on the sign of $V_{1}$.
For $p=2$, various possible shapes of $\mathscr{M}$ are illustrated graphically in Figures 3.23.4, along with matching example trajectories for $(y_{t},x_{t})$. In all figures, $(y_{t},x_{t})$ are generated by a canonical CKSVAR with $k=1$, $c=0$, $u_{t}\sim_{\textnormal{i.i.d.}}N[0,I_{2}]$, and $y_{0}=x_{0}=0$; the specification of the model is completed by specifying $\pi^{+}$, $\pi^{-}$ and $\Pi^{x}$ in (3.2), the values of which are given in each panel. The associated cointegrating vectors (on the half spaces $\mathscr{Z}^{+}$ and $\mathscr{Z}^{-}$) are the vectors orthogonal of $\mathscr{M}^{+}$ and $\mathscr{M}^{-}$, while the form of the limiting processes $Z$ is suggested by the shape of the set $\mathscr{M}$, where it concentrates. The main qualitative features of the three cases, as enumerated in Table 3.1 above, are as follows.
3.3.1 Case (i): regulated cointegration
Refer to caption Refer to caption
Figure 3.2: Case (i) configuration of $\mathscr{M}$ and trajectories of $(y_{t},x_{t})$
The distinguishing characteristic of this case is that the common trends are restricted to the region where $y\geq 0$, so that $Y$ will always be a Brownian motion regulated (from below) at zero, even though $y_{t}$ itself may take negative values. In the case that $q=1$, as e.g. when $p=2$ and $r=1$ as depicted in Figure 3.2, $X$ will also be a regulated process, i.e. we have cointegration where both processes share a common regulated stochastic trend. (When $q\geq 2$, $X$ will also depend on q1q-1 additional linear BMs.) Accordingly, a model configured as in case (i) would be most appropriate when $y_{t}$ appears to wander randomly above a threshold, and makes only brief sojourns below that threshold.
Example 3.2 (trending inflation target; ctd).
Recall that this model (with $r_{t}^{\ast}=r^{\ast}=0$) is described by:
$\pi_{t}-\overline{\pi}_{t}$ $=\theta(i_{t}^{+}+\mu i_{t}^{-}-\overline{\pi}_{t})+\varepsilon_{t}$ (2.4b)
$i_{t}$ $=\overline{\pi}_{t}+\gamma(\pi_{t}-\overline{\pi}_{t})$ (2.5b)
$\overline{\pi}_{t}$ $=\overline{\pi}_{t-1}+\delta(\pi_{t-1}-\overline{\pi}_{t-1})+\eta_{t},$
with the associated CKSVAR for $(i_{t},\pi_{t})$ being as in (2.9) above. In this first-order model, $\Pi^{\pm}=-\Phi^{\pm}(1)=\Phi_{1}^{\pm}-\Phi_{0}^{\pm}$, and thus it follows from (2.9) that
$\Pi^{+}$ $=\begin{bmatrix}\delta&-\delta\\ -\varphi_{1}&\varphi_{1}\end{bmatrix}=\begin{bmatrix}\delta\\ -\varphi_{1}\end{bmatrix}\begin{bmatrix}1&-1\end{bmatrix}$ $\Pi^{-}$ $=\begin{bmatrix}\delta&-\delta\\ -\varphi_{\mu}&\varphi_{1}\end{bmatrix}$. (3.6)
Suppose $\delta<0$, which in the context of 2.8 implies that past deviations from target indeed feed back into the central bank’s inflation target. Then unless $\mu=1$ (in which case the model is linear), $\varphi_{\mu}\neq\varphi_{1}$ and $\Pi^{-}$ has full rank. It follows that $\mathscr{M}^{-}=\{0\}$, whereas since $\operatorname{rk}\Pi^{+}=1$, $\mathscr{M}^{+}$ is the ‘half’ subspace orthogonal to $\beta\coloneqq(1,-1)^{\mathsf{T}}$, as depicted in Figure 3.2. It follows (via Theorem 4.1 below) that $i_{t}$ and $\pi_{t}$ are cointegrated, with cointegrating vector $\beta$, when $i_{t}$ is positive; but $I^{\ast}(0)$ when $i_{t}$ is negative; their common limiting stochastic trend is a regulated BM.
For the economics underlying this, note that (2.4b) and (2.5b) imply that when the solution to the model has $i_{t}>0$, i.e. when the ZLB is not binding, monetary policy is able to fully achieve its objectives, in the sense that inflation is stabilised to within an i.i.d. error of its target, as
$\pi_{t}-\overline{\pi}_{t}=(1-\gamma\theta)^{-1}\varepsilon_{t}$
As a consequence, (2.8) entails that $\overline{\pi}_{t}$ has a stochastic trend, which is inherited by both $\overline{\pi}_{t}$ and $i_{t}^{+}$ – in the latter case, because the equilibrium rate of interest implied by (2.5b) is equal to $\overline{\pi}_{t}$. Both $i_{t}^{+}$ and $\pi_{t}$ thus put the same loading on the common trend, whence $(1,-1)^{\mathsf{T}}$ is the cointegrating vector when $i_{t}>0$. On the other hand, when the solution to the model has $i_{t}<0$, i.e. when the ZLB binds, the lesser effectiveness of unconventional monetary policy ($\mu<1$) entails that policy is too contractionary, and so $\pi_{t}$ begins to drift below $\overline{\pi}_{t}$. However, via (2.8) this discrepancy raises the inflation target, and thereby raises the nominal rate of interest required to achieve target inflation. This feedback actually renders $\overline{\pi}_{t}$ as $I^{\ast}(0)$, and hence also $\pi_{t}$ and $i_{t}^{-}$. In a relatively short time, the solution to the model entails $i_{t}>0$ again, and thus the economy tends to spend relatively little time in the vicinity of the ZLB.
It will be observed that the qualitative behaviour described above is wholly contingent on $\mu<1$, i.e. on the ZLB as actually constraining the conduct of monetary policy. This manifests itself, quantitatively, as $\operatorname{rk}\Pi^{+}+1=\operatorname{rk}\Pi^{-}=2$ when $\mu<1$, as opposed to $\operatorname{rk}\Pi^{+}=\operatorname{rk}\Pi^{-}=2$ when $\mu=1$. Thus, in this setting, a test for $\Pi^{+}$ having reduced rank would amount to a test of the null that unconventional monetary policy is less effective than conventional policy, against the alternative that it is equally effective.
3.3.2 Case (ii): kinked cointegration
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 3.3: Case (ii) configurations of $\mathscr{M}$ and trajectories of $(y_{t},x_{t})$
This case is perhaps more reminiscent of linear cointegration, which it accommodates as a special case. Here $\mathscr{M}^{+}$ and $\mathscr{M}^{-}$ trace out ‘half’ subspaces of the same dimension $q$, but they need not be parallel, giving rise to a kink in $\mathscr{M}$ at the origin, with $\mathscr{M}$ itself being a linear cone. In general, $(Y,X)$ will follow a kinked Brownian motion driven by $p$ linear BMs, whose loadings, and the associated $r$ cointegrating relations, vary with the sign of $Y$. In the top panel of Figure 3.3, $y_{t}$ and $x_{t}$ are cointegrated, but with distinct cointegrating vectors $\beta^{+}=(1,-1)^{\mathsf{T}}$ or $\beta^{-}=(\tfrac{1}{2},-1)^{\mathsf{T}}$ applying on $\mathscr{Z}^{+}$ or $\mathscr{Z}^{-}$, i.e. when $y_{t}$ is positive or negative. We have a kind of ‘threshold cointegration’, with the movement of $y_{t}$ across zero causing the model to switch between distinct cointegrating spaces (cf. Saikkonen and Choi, 2004). This switch is reflected in the trajectories of $y_{t}$ and $x_{t}$: when $y_{t}\geq 0$, the two series move together approximately one-for-one; whereas if $y_{t}\leq 0$, $y_{t}$ changes by two units for every one-unit change in $x_{t}$, in the long run. Relative to case (i), the trajectories of $\{y_{t}\}$ will now much more closely resemble those of a linear unit root process, and $\{y_{t}\}$ will accordingly tend to spend long stretches in both the positive and negative regions, with no tendency to revert to one or the other.
The bottom panel of Figure 3.3 presents an important special case where the cointegrating relationships are such that $\{x_{t}\}$ behaves like an $I^{\ast}(0)$ process when $y_{t}<0$, but cointegrates with $y_{t}$ when $y_{t}>0$; the limiting process $X$ will thus be a censored Brownian motion.
Example 3.3 (trending natural rate; ctd).
This model (with $\overline{\pi}_{t}=\overline{\pi}=0$) may be rendered as
$\pi_{t}$ $=\theta(i_{t}^{+}+\mu i_{t}^{-}-r_{t}^{\ast})+\varepsilon_{t}$ (2.4a)
$i_{t}$ $=r_{t}^{\ast}+\gamma\pi_{t}$ (2.5a)
$r_{t}^{\ast}$ $=r_{t-1}^{\ast}+\eta_{t}.$ (2.6)
Then similarly to (3.6), we have from the corresponding CKSVAR for $(i_{t},\pi_{t})$, as given in (2.7) above, that
$\Pi^{+}$ $=\begin{bmatrix}0\\ -(1-\theta\gamma)\end{bmatrix}\begin{bmatrix}0&1\end{bmatrix}$ $\Pi^{-}$ $=\begin{bmatrix}0\\ -(1-\theta\gamma)\end{bmatrix}\begin{bmatrix}\gamma^{-1}\tau_{\mu}&1\end{bmatrix}$,
where $\gamma^{-1}\tau_{\mu}=\theta(1-\mu)(1-\theta\gamma)^{-1}\leq 0$, with strict inequality unless $\mu=1$. Thus $\operatorname{rk}\Pi^{+}=\operatorname{rk}\Pi^{-}=1$, and (by Theorem 4.2 below), there is a common stochastic trend present in both regimes – but when $i_{t}>0$ this loads only on $i_{t}$, and the cointegrating vector is $\beta^{+}=(0,1)^{\mathsf{T}}$. If unconventional policy is as effective as conventional policy ($\mu=1$), this holds also when $i_{t}<0$; otherwise the trend is shared by both $i_{t}$ and $\pi_{t}$, and their cointegrating vector in the negative regime is $\beta^{-}=(\gamma^{-1}\tau_{\mu},1)^{\mathsf{T}}$. Qualitatively, the behaviour of the series is as plotted for $(y_{t},x_{t})$ in the second row of Figure 3.3 (with $y_{t}=-i_{t}$ and $x_{t}=-\pi_{t}$, i.e. with the signs reversed); in the limit, $n^{-1/2}\pi_{\lfloor n\lambda\rfloor}$ will be a censored BM (from above).
To account for this in economic terms, recall that (2.4a) and (2.5a) imply that when $i_{t}>0$, the central bank is able to fully stabilise inflation, in the sense that $\pi_{t}=\overline{\pi}_{t}+(1-\gamma\theta)^{-1}\varepsilon_{t}=(1-\gamma\theta)^{-1}\varepsilon_{t}$, since $\overline{\pi}_{t}$ is constant and normalised to zero. Thus $\pi_{t}\mathbf{1}\{i_{t}>0\}\sim I^{\ast}(0)$; whereas, $i_{t}^{+}\sim I^{\ast}(1)$, since it inherits the stochastic trend in the natural rate of interest. However, if the model solution requires $i_{t}<0$, then the ZLB constraint inhibits the operation of monetary policy (if $\mu<1$), and inflation begins to drift away from its target. That drift corresponds to the stochastic trend in $r_{t}^{\ast}$, which is thus present in both $i_{t}^{-}$ and $\pi_{t}$, and hence these series cointegrate.
By contrast, if monetary policy is not effectively constrained by the ZLB, then $\pi_{t}$ would remain $I^{\ast}(0)$, irrespective of the sign of $i_{t}$. Thus the long run behaviour of $\pi_{t}$ – whether it is $I^{\ast}(0)$, or whether it follows a stochastic trend (and so is $I^{\ast}(1)$) when interest rates are at the ZLB – here provides identifying information on the relative effectiveness of unconventional monetary policy, i.e. on whether the ZLB is ever a truly binding constraint on the central bank, just as it did in Example 2.1.
Example 3.4 (trending inflation target; ctd).
Suppose that $\delta=0$ in (2.8), so the model is now described by (2.4b), (2.5b) and
$\overline{\pi}_{t}$ $=\overline{\pi}_{t-1}+\eta_{t}.$
The inflation target thus remains time-varying, but there is no longer any feedback from past failures to hit that target. Then (3.6) simplifies to
$\Pi^{+}$ $=\begin{bmatrix}0\\ -\varphi_{1}\end{bmatrix}\begin{bmatrix}1&-1\end{bmatrix}$ $\Pi^{-}$ $=\begin{bmatrix}0\\ -\varphi_{1}\end{bmatrix}\begin{bmatrix}\varphi_{1}^{-1}\varphi_{\mu}&-1\end{bmatrix}$,
where $\varphi_{1}^{-1}\varphi_{\mu}=1+\frac{\theta(\gamma-1)(1-\mu)}{1-\theta\gamma}\in(0,1]$. Thus $i_{t}$ and $\pi_{t}$ are $I^{\ast}(1)$ and cointegrated, but with cointegrating relations $\beta^{+}=(1,-1)^{\mathsf{T}}$ and $\beta^{-}=(\varphi_{1}^{-1}\varphi_{\mu},-1)^{\mathsf{T}}$ that depend on the sign of $i_{t}$, unless $\mu=1$. Even though $i_{t}^{-}$ is unobserved, we can still distinguish between the cases $\mu=1$ and $\mu<1$ on the basis that, when $\mu<1$, the long-run variance of $\pi_{t}$ will differ depending on whether $i_{t}>0$ or $i_{t}=0$, as is evident in the behaviour of $x_{t}$ in the top panel of Figure 3.3.888Theorem 4.2 below implies that $n^{-1/2}\pi_{\lfloor n\lambda\rfloor}$ converges weakly to $\sigma_{\varepsilon}(\gamma-1)[1+(\varphi_{1}^{-1}\varphi_{\mu}-1)\mathbf{1}\{W(\lambda)<0\}]W(\lambda)$, for $\sigma_{\varepsilon}^{2}\coloneqq\mathbb{E}\varepsilon_{t}^{2}$ and $W$ a standard BM.
3.3.3 Case (iii): linear cointegration in a nonlinear VECM
Refer to caption Refer to caption
Figure 3.4: Case (iii) configuration of $\mathscr{M}$ and trajectories of $(y_{t},x_{t})$
This case, which is depicted in the Figure 3.4, entails that no trends are present in $\{y_{t}\}$, which is in fact a stationary process. The common trends are loaded entirely on $\{x_{t}\}$, and the cointegrating relationships between the elements of $x_{t}$ are unaffected by the sign of $y_{t}$, exactly as in literature on regime-switching nonlinear VECM models (see the text following (1.1) above). For this reason, a model configured as in case (iii) is appropriate only when the univariate behaviour of $\{y_{t}\}$ appears stationary, in which case there will be frequent switches of ‘regime’, particularly if the mean of $y_{t}$ is near zero.
4 Representation theory and asymptotics
We now proceed to develop analogues of the Granger–Johansen representation theorem for cases (i)–(iii). These results are of interest in their own right, because by characterising the processes generated by the model, under alternative configurations of $\Pi^{\pm}$, they delimit the classes of observed time series that the model might be fruitfully applied to. Indeed, they indicate precisely the restrictions on $\Pi^{\pm}$ that might be appropriate for specific applications, according to whether $\{y_{t}\}$ is observed to either: wander randomly above a threshold, but spend only brief periods below it (case (i)); wander randomly on both sides of a threshold (case (ii)); or behave like a stationary process (case (iii)). Beyond such guidance, our results also lay the groundwork for the development of the asymptotics of likelihood-based estimators of the CKSVAR model in the presence of unit roots, by establishing the asymptotic behaviour of the (standardised) processes generated by the model.
To facilitate the exposition, we shall initially suppose the data to be generated by a canonical CKSVAR, i.e. DGP holds, and that there are no deterministic trends in cases (i) and (ii), i.e. DET holds. Theorems 4.14.3 below are stated under these assumptions: the minor modifications required when DGP is replaced by DGP are given in the subsequent remarks, while a general treatment of deterministic trends follows in Section 4.4. To avoid the need to specify how some of the quantities below should be defined when $k=1$, we shall treat this as a special case of the model with $k=2$, in which $\phi_{2}^{+}=\phi_{2}^{-}=0$ and $\Phi_{2}^{x}=0$; and thus henceforth $k\geq 2$, unless otherwise stated. Because of the overlap between the arguments used to analyse the CKSVAR in each case, it will be occasionally necessary to redefine objects that have already appeared in the discussion of another case. While we have endeavoured to keep such notational conflicts to a minimum, and explicitly indicated wherever these arise, the reader is advised to treat each of the following three subsections, and the accompanying subsections of Appendix B where the proofs of the theorems appear, as independent of each other.
Our proofs make repeated use of companion form representations of the VECM (3.2), formulated slightly differently for each of cases (i)–(iii). In this respect, the nearest analogue to our arguments, in the setting of a linear VAR, is provided by Hansen (2005).
Notation.
For $A\in\mathbb{R}^{m\times n}$ having full column rank, $A_{\perp}\in\mathbb{R}^{m\times(m-n)}$ denotes a full column rank matrix such that $A_{\perp}^{\mathsf{T}}A=0$; we refer to $A_{\perp}$ (which is unique only up to its column span) as ‘the orthocomplement of $A$’. (Note that it is not implied that the columns of $A_{\perp}$ should be orthogonal vectors; any further normalisation of $A_{\perp}$ will be noted in the text if required.) $A_{i:j}$ denotes the submatrix formed from rows $\{i,i+1,\ldots,j\}$ of $A$.
4.1 Case (i): regulated cointegration
Recalling Table 3.1, we have the familiar factorisation
$\Pi^{+}=\alpha^{+}\beta^{+\mathsf{T}},$
where $\alpha^{+},\beta^{+}\in\mathbb{R}^{p\times r}$ have rank $r$. As we show below, $\beta^{+}$ spans the cointegrating space on $\mathscr{Z}^{+}=\mathbb{R}_{+}\times\mathbb{R}^{p-1}$ (recall Definition 3.2), so that $(y_{t}^{+},x_{t})\sim I^{\ast}(1)$ but $\beta^{+\mathsf{T}}(y_{t}^{+},x_{t})\sim I^{\ast}(0)$. Moreover, since $y_{t}^{-}\sim I^{\ast}(0)$, it follows that $\beta^{+\mathsf{T}}(y_{t},x_{t})\sim I^{\ast}(0)$, so that the columns of $\beta^{+}$ are globally cointegrating vectors.
In a linear cointegrated VAR, no assumptions additional to CVAR and ERR are needed, but the nonlinearity of the CKSVAR prevents our assumptions on the roots of $\Phi^{\pm}(\lambda)$ from being sufficient to ensure that the short-memory components (i.e. the equilibrium errors and the first differences) are indeed $I^{\ast}(0)$. For this reason, two further regularity conditions are required. To state these, let
$\boldsymbol{F}_{\delta}\coloneqq\begin{bmatrix}I_{p(k-1)+r}+\boldsymbol{\beta}^{+\mathsf{T}}\boldsymbol{\alpha}^{+}&\boldsymbol{\beta}_{1:p}^{+\mathsf{T}}(\phi_{1}^{-}-\phi_{1}^{+})\delta&\boldsymbol{\beta}_{1:p}^{+\mathsf{T}}(\varphi^{-}-\varphi^{+})\\ e_{1}^{\mathsf{T}}\boldsymbol{\alpha}^{+}&[1+e_{1}^{\mathsf{T}}(\phi_{1}^{-}-\phi_{1}^{+})]\delta&e_{1}^{\mathsf{T}}(\varphi^{-}-\varphi^{+})\\ 0&\delta&0\\ 0&0&D\end{bmatrix}$ (4.1)
where $\varphi^{\pm}\coloneqq[\phi_{2}^{\pm},\ldots,\phi_{k}^{\pm}]$, $D\coloneqq[I_{k-2},0_{(k-2)\times 1}]$; $\boldsymbol{\alpha}^{+},\boldsymbol{\beta}^{+}\in\mathbb{R}^{kp\times[p(k-1)+r]}$ with
$\boldsymbol{\alpha}^{+}$ $\coloneqq\begin{bmatrix}\alpha^{+}&\Gamma_{1}^{+}&\cdots&\Gamma_{k-1}^{+}\\ &I_{p}\\ &&\ddots\\ &&&I_{p}\end{bmatrix}$, $\boldsymbol{\beta}^{+\mathsf{T}}$ $\coloneqq\begin{bmatrix}\beta^{+\mathsf{T}}\\ I_{p}&-I_{p}\\ &\ddots&\ddots\\ &&I_{p}&-I_{p}\end{bmatrix}$; (4.2)
[…]
$P_{\beta_{\perp}^{+}}\coloneqq\beta_{\perp}^{+}[\alpha_{\perp}^{+\mathsf{T}}\Gamma^{+}(1)\beta_{\perp}^{+}]^{-1}\alpha_{\perp}^{+\mathsf{T}}$ (4.3)
for $\alpha_{\perp}^{+}$ and $\beta_{\perp}^{+}$ the orthocomplements of $\alpha^{+}$ and $\beta^{+}$. Let $\rho(M)$ denote the spectral radius of $M\in\mathbb{R}^{m\times m}$, and for $\mathcal{A}\subset\mathbb{R}^{m\times m}$ a bounded collection of matrices, let
$\rho_{{\scriptstyle\mathrm{JSR}}}(\mathcal{A})\coloneqq\limsup_{t\rightarrow\infty}\sup_{B\in\mathcal{A}^{t}}\rho(B)^{1/t}$
denote its joint spectral radius (JSR; e.g. Jungers, 2009, Defn. 1.1), for $\mathcal{A}^{t}\coloneqq\{\prod_{s=1}^{t}M_{s}\mid M_{s}\in\mathcal{A}\}$ the set of $t$-fold products of matrices in $\cal A$}. [^10] For a further discussion of the JSR, and references to the literature on methods for numerically approximating it, see Duffy et al. (2023, Sec. 4.1). Let $z_{t}\coloneqq(y_{t},x_{t}^{\mathsf{T}})^{\mathsf{T}}$.
Assumption CO(i).
  1. 1.
    $r^{+}=\operatorname{rk}\Pi^{x}=r$ and $r^{-}=r+1$, for some $r\in\{0,1,\ldots,p-1\}$.
  2. 2.
    $\rho_{{\scriptstyle\mathrm{JSR}}}(\{\boldsymbol{F}_{0},\boldsymbol{F}_{1}\})<1$.
  3. 3.
    $\kappa_{1}<0$, where $\kappa_{1}$ denotes the first element of $\kappa\coloneqq P_{\beta_{\perp}^{+}}\pi^{-}$.
  4. 4.
    1. a.
      $\beta^{+\mathsf{T}}z_{t}$, $y_{t}^{-}$, and $\Delta z_{t}$ have uniformly bounded $2+\delta_{u}$ moments, for $t\in\{-k+1,\ldots,0\}$.
    2. b.
      $n^{-1/2}z_{0}\overset{p}{\rightarrow}{\cal Z}_{0}=[\begin{smallmatrix}{\cal Y}_{0}\\ {\cal X}_{0}\end{smallmatrix}]\in\mathscr{M}$, where ${\cal Z}_{0}$ is non-random.
Theorem 4.1.
Suppose DGP, ERR, CVAR and CO(i) hold. Then:
  1. (i)
    $\xi_{t}^{+}\coloneqq\beta^{+\mathsf{T}}z_{t}\sim I^{\ast}(0)$, $\Delta z_{t}\sim I^{\ast}(0)$, and $y_{t}^{-}\sim I^{\ast}(0)$;
and if additionally DET holds:
  1. (ii)
    for $U_{0}(\lambda)\coloneqq\Gamma^{+}(1){\cal Z}_{0}+U(\lambda)$, jointly with (3.4),
    $\begin{bmatrix}Y_{n}(\lambda)\\ X_{n}(\lambda)\end{bmatrix}\rightsquigarrow\begin{bmatrix}Y(\lambda)\\ X(\lambda)\end{bmatrix}=P_{\beta_{\perp}^{+}}U_{0}(\lambda)+\kappa_{1}^{-1}\kappa\sup_{\lambda^{\prime}\leq\lambda}[-e_{1}^{\mathsf{T}}P_{\beta_{\perp}^{+}}U_{0}(\lambda^{\prime})]_{+}$
    on $D[0,1]$, where in particular $Y(\lambda)=e_{1}^{\mathsf{T}}P_{\beta_{\perp}^{+}}U_{0}(\lambda)+\sup_{\lambda^{\prime}\leq\lambda}[-e_{1}^{\mathsf{T}}P_{\beta_{\perp}^{+}}U_{0}(\lambda^{\prime})]_{+}$.
Remark 4.1.
(i). If DGP is replaced by DGP, then the theorem continues to hold as stated, except that CO(i).2 should be replaced by
  1. CO(i).2
    $\rho_{{\scriptstyle\mathrm{JSR}}}(\{\tilde{\boldsymbol{F}}_{0},\tilde{\boldsymbol{F}}_{1}\})<1$;
where the tildes refer to the parameters of the canonical CKSVAR derived from the structural form via Proposition 2.1. (This obtains because the derived canonical variables then themselves follow a canonical CKSVAR that satisfies the conditions of the theorem; see Appendix D.)
(ii). The contrast with a linear cointegrated VAR is marked. $Y$ is now a regulated Brownian motion, which also enters into other components of $X$. Indeed, as noted in Section 3.3 above, some components of $X$ may themselves be regulated BMs.
(iii). Part (i) of the theorem is proved by obtaining a nonlinear VAR representation for $\xi_{t}^{+}$, $\Delta z_{t}$ and $y_{t}^{-}$, whose companion form can be expressed in terms of the matrices $\{\boldsymbol{F}_{\delta}\mid\delta\in[0,1]\}$ (Lemma B.1). Since the parameters of that VAR depend on $y_{t}^{+}$, these processes cannot be stationary, but CO(i).2 ensures that the system is sufficiently ‘constrained’ that they will be $I^{\ast}(0)$. A necessary but not sufficient condition for CO(i).2 is that $\boldsymbol{F}_{0}$ and $\boldsymbol{F}_{1}$ have all their eigenvalues inside the unit circle. (This is implied by CVAR: see Appendix E.)
(iv). It will be seen from the proof that part (ii) holds if CO(i).2 is replaced by any condition sufficient to ensure $(\xi_{t}^{+},\Delta z_{t},y_{t}^{-})\sim I^{\ast}(0)$. There may thus be some scope for relaxing this assumption, which takes essentially a worst-case approach to the behaviour of the nonlinear VAR governing the evolution of these processes. However, even in the more tractable univariate set-up of Bykhovskaya and Duffy (2024), it is far from obvious what this condition (their Assumption A4) might be replaced by.
(v). In deriving the weak limit of $n^{-1/2}y_{\lfloor n\lambda\rfloor}$, a key step is to obtain a univariate representation for $y_{t}^{+}$ as a regulated process. The proof of Theorem 4.1 shows that it is possible to write $y_{t}^{+}-\kappa_{1}y_{t}^{-}=w_{t}$ for a certain series $\{w_{t}\}$: the role of CO(i).3 is to ensure that this equation is solved uniquely by taking $y_{t}^{+}=[w_{t}]_{+}$ and $y_{t}^{-}=\kappa_{1}^{-1}[w_{t}]_{-}$. It is possible that this condition may be redundant: as shown in Appendix E, it is implied by our other assumptions if either $k=1$ or $p=1$.
Example 4.1 (univariate; ctd).
In the univariate ($p=1$) model (2.10), case (i) with $r=0$ corresponds to a model in which $c=0$, $\sum_{i=1}^{k}\phi_{i}^{+}=1$ and $\phi^{-}(\lambda)$ has all its roots outside the unit circle, so that
$y_{t}=\sum_{i=1}^{k}(\phi_{i}^{+}y_{t-i}^{+}+\phi_{i}^{-}y_{t-i}^{-})+u_{t}=(1+\pi^{+})y_{t-1}^{+}+\sum_{i=1}^{k-1}\gamma_{i}^{+}\Delta y_{t-i}^{+}+\sum_{i=1}^{k}\phi_{i}^{-}y_{t-i}^{-}+u_{t}$ (4.4)
may be loosely regarded as an autoregressive model with a unit root regime (since $\pi^{+}=0$) and a stationary regime (though the model technically has $2^{k}$ distinct autoregressive regimes). Theorem 4.1 implies that $y_{t}^{-}\sim I^{\ast}(0)$, and
$Y_{n}(\lambda)\rightsquigarrow Y(\lambda)=K(\lambda)-\sup_{\lambda^{\prime}\leq\lambda}[-K(\lambda^{\prime})]_{+}$ (4.5)
on $D[0,1]$, where $K(\lambda)=\gamma^{+}(1)^{-1}U(\lambda)$, for $\gamma^{+}(1)=1-\sum_{i=1}^{k-1}\gamma_{i}^{+}$. $Y$ is thus a regulated Brownian motion. (4.5) extends the results of Liu et al. (2011, Thm. 3.1) and Gao et al. (2013) from a first- to a higher-order autoregressive setting. It also agrees with the limit theory developed in Bykhovskaya and Duffy (2024, Thm. 3.2), when their censored dynamic Tobit model (in which $\phi_{i}^{-}=0$ for all $i\in\{1,\ldots,k\}$) is specialised to one with an exact unit root and no intercept.
4.2 Case (ii): kinked cointegration
We turn next to the case in which the cointegrating rank $r$ is the same across both the positive and negative regimes, though the cointegrating space itself need not be. Here we also suppose that $\operatorname{rk}\Pi^{x}=r$, which as discussed in Section 3.3 entails that $y_{t}\sim I^{\ast}(1)$. Under the foregoing, we must have $\pi^{\pm}\in\operatorname{sp}\Pi^{x}$, and so
$\Pi^{\pm}=\Pi^{x}\begin{bmatrix}\theta^{\pm}&I_{p-1}\end{bmatrix}=\alpha\begin{bmatrix}\beta_{y}^{\pm}&\beta_{x}^{\mathsf{T}}\end{bmatrix}\eqqcolon\alpha\beta^{\pm\mathsf{T}},$ (4.6)
where $\alpha\in\mathbb{R}^{p\times r}$, $\beta_{x}\in\mathbb{R}^{(p-1)\times r}$ and $\beta^{\pm}\in\mathbb{R}^{p\times r}$ have rank $r$, and $\theta^{\pm}\in\mathbb{R}^{p-1}$ is such that $\Pi^{x}\theta^{\pm}=\pi^{\pm}$. Let $\mathbf{1}^{+}(y)\coloneqq\mathbf{1}\{y\geq 0\}$ and $\mathbf{1}^{-}(y)\coloneqq\mathbf{1}\{y<0\}$, and set $\beta(y)\coloneqq\beta^{+}\mathbf{1}^{+}(y)+\beta^{-}\mathbf{1}^{-}(y)$. [^11] There is unavoidably some arbitrariness with respect to how such objects are defined when $y=0$, but since these only play a role in the model when multiplied by $y$, it does not matter which convention is adopted. Throughout the paper, we use the functions $\mathbf{1}^{\pm}(y)$ to ensure that all such definitions are mutually consistent. Then we can define the equilibrium errors as
$\xi_{t}\coloneqq\beta(y_{t})^{\mathsf{T}}z_{t}=\mathbf{1}^{+}(y_{t})\beta^{+\mathsf{T}}z_{t}+\mathbf{1}^{-}(y_{t})\beta^{-\mathsf{T}}z_{t}.$
Observe how (4.6) implies that the ‘loadings’ $\alpha$ of the equilibrium errors will be the same in both regimes, even though the cointegrating vectors that define those errors need not be. (Case (iii) entails the opposite, with fixed cointegrating vectors but loadings that depend on the sign of $y_{t}$: see (4.15) below.)
The theorem below establishes that $\xi_{t}\sim I^{\ast}(0)$, and that $\beta^{+}$ and $\beta^{-}$ span the cointegrating spaces on $\mathscr{Z}^{+}$ and $\mathscr{Z}^{-}$ respectively. The limiting common trends are kinked Brownian motions given by a kind of projection of $U$ onto $\mathscr{M}$, defined in terms of
$P_{\beta_{\perp}}(y)\coloneqq\beta_{\perp}(y)[\alpha_{\perp}^{\mathsf{T}}\Gamma(1;y)\beta_{\perp}(y)]^{-1}\alpha_{\perp}^{\mathsf{T}},$ (4.7)
$\begin{aligned} \beta_{\perp}(y)&\coloneqq\begin{bmatrix}1&0\\ -\theta(y)&\beta_{x,\perp}\end{bmatrix},&\qquad\Gamma(1;y)&\coloneqq\Gamma^{+}(1)\mathbf{1}^{+}(y)+\Gamma^{-}(1)\mathbf{1}^{-}(y)\end{aligned},$ (4.8)
where $\theta(y)\coloneqq\mathbf{1}^{+}(y)\theta^{+}+\mathbf{1}^{-}(y)\theta^{-}$. Such objects as $P_{\beta_{\perp}}(y)$ take only two distinct values, depending on the sign of $y$, and we routinely use the notation $P_{\beta_{\perp}}(+1)$ and $P_{\beta_{\perp}}(-1)$ to indicate these. Similarly to case (i), beyond our assumptions on the ranks of $\Pi^{\pm}$ and $\Pi^{x}$, two further regularity conditions are needed to ensure that the system is well behaved. To state these, let $\boldsymbol{\alpha},\boldsymbol{\beta}(y)\in\mathbb{R}^{[k(p+1)-1]\times[r+(k-1)(p+1)]}$ with
$\boldsymbol{\alpha}\coloneqq$ $\begin{bmatrix}\alpha&\Gamma_{1}&\Gamma_{2}&\cdots&\Gamma_{k-1}\\ &I_{p+1}\\ &&I_{p+1}\\ &&&\ddots\\ &&&&I_{p+1}\end{bmatrix}$, $\boldsymbol{\beta}(y)^{\mathsf{T}}$ $\coloneqq\begin{bmatrix}\beta(y)^{\mathsf{T}}\\ S(y)&-I_{p+1}\\ &I_{p+1}&-I_{p+1}\\ &&\ddots&\ddots\\ &&&I_{p+1}&-I_{p+1}\end{bmatrix}$, (4.9)
where $\Gamma_{i}\coloneqq[\gamma_{i}^{+},\gamma_{i}^{-},\Gamma_{i}^{x}]$ for $i\in\{1,\ldots,k-1\}$, and
$S(y)\coloneqq\begin{bmatrix}\mathbf{1}^{+}(y)&0\\ \mathbf{1}^{-}(y)&0\\ 0&I_{p-1}\end{bmatrix}$ (4.10)
so that $S(y_{t})z_{t}=(y_{t}^{+},y_{t}^{-},x_{t}^{\mathsf{T}})^{\mathsf{T}}$, where $z_{t}=(y_{t},x_{t}^{\mathsf{T}})^{\mathsf{T}}$.
Assumption CO(ii).
  1. 1.
    $r^{+}=r^{-}=\operatorname{rk}\Pi^{x}=r$, for some $r\in\{0,1,\ldots,p-1\}$.
  2. 2.
    $\rho_{{\scriptstyle\mathrm{JSR}}}(\{I+\boldsymbol{\beta}(+1)^{\mathsf{T}}\boldsymbol{\alpha},I+\boldsymbol{\beta}(-1)^{\mathsf{T}}\boldsymbol{\alpha}\})<1$.
  3. 3.
    $\operatorname{sgn}\det\alpha_{\perp}^{\mathsf{T}}\Gamma(1;+1)\beta_{\perp}(+1)=\operatorname{sgn}\det\alpha_{\perp}^{\mathsf{T}}\Gamma(1;-1)\beta_{\perp}(-1)\neq 0$.
  4. 4.
    1. a.
      $\beta(y_{t})^{\mathsf{T}}z_{t}$, and $\Delta z_{t}$ have uniformly bounded $2+\delta_{u}$ moments, for $t\in\{-k+1,\ldots,0\}$.
    2. b.
      $n^{-1/2}z_{0}\overset{p}{\rightarrow}{\cal Z}_{0}=[\begin{smallmatrix}{\cal Y}_{0}\\ {\cal X}_{0}\end{smallmatrix}]\in\mathscr{M}$, where ${\cal Z}_{0}$ is non-random.
Theorem 4.2.
Suppose DGP, ERR, CVAR and CO(ii) hold, and let $\vartheta^{\mathsf{T}}\coloneqq e_{1}^{\mathsf{T}}P_{\beta_{\perp}}(+1)$ and $U_{0}(\lambda)\coloneqq\Gamma(1;{\cal Y}_{0}){\cal Z}_{0}+U(\lambda)$. Then:
  1. (i)
    $\xi_{t}\coloneqq\beta(y_{t})^{\mathsf{T}}z_{t}\sim I^{\ast}(0)$ and $\Delta z_{t}\sim I^{\ast}(0)$;
and if additionally DET holds:
  1. (ii)
    on $D[0,1]$, jointly with (3.4),
    $\begin{bmatrix}Y_{n}(\lambda)\\ X_{n}(\lambda)\end{bmatrix}\rightsquigarrow\begin{bmatrix}Y(\lambda)\\ X(\lambda)\end{bmatrix}=P_{\beta_{\perp}}[Y(\lambda)]U_{0}(\lambda)=P_{\beta_{\perp}}[\vartheta^{\mathsf{T}}U_{0}(\lambda)]U_{0}(\lambda),$ (4.11)
    where in particular $\operatorname{sgn}Y(\lambda)=\operatorname{sgn}\vartheta^{\mathsf{T}}U_{0}(\lambda)$.
Remark 4.2.
(i). Similarly to Remark 4.14.1 above, if DGP is replaced by DGP, then the theorem continues to hold exactly as stated, except that CO(ii).2 should be modified to
  1. CO(ii).2
    $\rho_{{\scriptstyle\mathrm{JSR}}}(\{I+\tilde{\boldsymbol{\beta}}(+1)^{\mathsf{T}}\tilde{\boldsymbol{\alpha}},I+\tilde{\boldsymbol{\beta}}(-1)^{\mathsf{T}}\tilde{\boldsymbol{\alpha}}\})<1$,
where the tildes refer to the parameters of the canonical CKSVAR derived from the structural form via Proposition 2.1. (See Appendix D.)
(ii). Even when $\beta^{+}=\beta^{-}$, such that the cointegrating space is the same in both the positive and negative regimes, $(Y,X)$ will generally be a kinked Brownian motion because of the residual dependence of $P_{\beta_{\perp}}(y)=\beta_{\perp}[\alpha_{\perp}^{\mathsf{T}}\Gamma(1;y)\beta_{\perp}]^{-1}\alpha_{\perp}^{\mathsf{T}}$ on $y$ via $\Gamma(1;y)$. Indeed, in the univariate model (2.10) under case (ii) with $r=0$,
$\Delta y_{t}=\sum_{i=1}^{k-1}(\gamma_{i}^{+}\Delta y_{t-i}^{+}+\gamma_{i}^{-}\Delta y_{t-i}^{-})+u_{t},$
Theorem 4.2 entails $Y(\lambda)=\gamma[1;U_{0}(\lambda)]^{-1}U_{0}(\lambda)$ is a kinked Brownian motion, whose variance depends on the sign of $U_{0}(\lambda)$.
(iii). CO(ii).2 plays an analogous role to CO(i).2 above, ensuring that the nonlinear VAR representation (see Lemma B.2) obtained for the short-memory components $(\xi_{t},\Delta z_{t})$ is sufficiently well-behaved that these processes are $I^{\ast}(0)$. Part (ii) would continue to hold if CO(ii).2 were replaced by any other condition sufficient for $(\xi_{t},\Delta z_{t})\sim I^{\ast}(0)$. A necessary condition for CO(ii).2 is that the eigenvalues of $I+\boldsymbol{\beta}(\pm 1)^{\mathsf{T}}\boldsymbol{\alpha}$ should lie strictly inside the unit circle, which is implied by CVAR (see Lemma A.3).
(iv). The map $(y,u)\mapsto P_{\beta_{\perp}}(y)u$ is not in general continuous, a fact that could interfere with the convergence in (4.11). CO(ii).3 helps to ensure this map is continuous on a sufficiently large domain to permit (4.11) to follow via an application of the continuous mapping theorem (CMT), with $Y$ and $X$ having continuous paths.
(v). Given $\alpha,\beta\in\mathbb{R}^{p\times r}$ with full column rank, their orthocomplements $\alpha_{\perp},\beta_{\perp}\in\mathbb{R}^{p\times q}$ are unique only up to their column span. Since in a linearly cointegrated system CVAR implies that $\alpha_{\perp}^{\mathsf{T}}\Gamma(1)\beta_{\perp}$ has nonzero determinant (Lemma A.3), we may normalise $\alpha_{\perp}$ and/or $\beta_{\perp}$ so that $\det\alpha_{\perp}^{\mathsf{T}}\Gamma(1)\beta_{\perp}=1$. It should therefore be emphasised that CO(ii).3 applies when $\beta_{\perp}(+1)$ and $\beta_{\perp}(-1)$ are related via (4.8), so we are not entirely free to choose $\beta_{\perp}(\pm 1)$ such that the signs of these determinants can be brought into agreement.
Example 4.2.
To illustrate how Theorem 4.2 may be applied to derive the long-run behaviour of a structural model – even one in which the observables do not follow a (nonlinear) VAR – here we consider the model of Ascari et al. (2023). They regard a vector of observable series $w_{t}$ (the inflation rate, GDP per capita, and the nominal interest rate) as fluctuating in a stationary manner around their long-run components $\bar{w}_{t}$, as per
$A(L)(w_{t}-\bar{w}_{t})=\varepsilon_{t},$
where $A(L)$ is the lag polynomial of a stationary VAR. The long-run behaviour of $w_{t}$ will thus be governed by that of $\bar{w}_{t}$, which consists of trend inflation $\bar{\pi}_{t}$, potential output $\bar{y}_{t}$, and the trend nominal interest rate $\bar{\imath}_{t}$. The first two components are assumed, together with the trend growth rate $g_{t}$ of potential output, to evolve according to the nonlinear VAR
$\bar{\pi}_{t}$ $=\bar{\pi}_{t-1}+u_{t}^{\pi}$ (4.12a)
$\Delta\bar{y}_{t}-\delta(\bar{\pi}_{t}-\tau)$ $=-\delta(\bar{\pi}_{t-1}-\tau)+g_{t-1}+u_{t}^{y}$ (4.12b)
$g_{t}$ $=g_{t-1}+u_{t}^{g},$ (4.12c)
where $\delta(\cdot)$ is a piecewise linear function of the form $\delta(x)=\delta^{+}[x]_{-}+\delta^{-}[x]_{+}$, which captures their ‘long-run Phillips curve’; see their equations (5)–(9). (Since we can trivially rewrite the model in terms of $\bar{\pi}_{t}-\tau$ rather than $\bar{\pi}_{t}$, here we also suppose that $\tau=0$, for simplicity.) The trend in the nominal interest rate is specified to be a linear function of $\bar{\pi}_{t}$, $g_{t}$, and an additional stochastic trend $z_{t}$, as per their equations (10) and (11):
$\bar{\imath}_{t}$ $=\bar{\pi}_{t}+cg_{t}+z_{t}$ (4.13a)
$z_{t}$ $=z_{t-1}+u_{t}^{z}.$ (4.13b)
Since $\bar{\imath}_{t}$ does not enter (4.12), we can analyse the long-run behaviour of this subsystem separately, and thence deduce that of $\bar{\imath}_{t}$.
Because $\bar{\pi}_{t}$ enters (4.12) nonlinearly, the long-run properties of $(\bar{\pi},\Delta\bar{y}_{t},g_{t})$ cannot be determined on the basis of the ordinary Granger–Johansen representation theorem, but instead requires an application of the theory developed in the present work. To that end, we render the system (4.12) in CKSVAR form as
$\begin{bmatrix}1\\ -\delta^{+}\\ 0\end{bmatrix}\bar{\pi}_{t}^{+}+\begin{bmatrix}1\\ -\delta^{-}\\ 0\end{bmatrix}\bar{\pi}_{t}^{-}+\begin{bmatrix}0&0\\ 1&0\\ 0&1\end{bmatrix}\begin{bmatrix}\Delta\bar{y}_{t}\\ g_{t}\end{bmatrix}\\ =\begin{bmatrix}1\\ -\delta^{+}\\ 0\end{bmatrix}\bar{\pi}_{t-1}^{+}+\begin{bmatrix}1\\ -\delta^{-}\\ 0\end{bmatrix}\bar{\pi}_{t-1}^{-}+\begin{bmatrix}0&0\\ 0&1\\ 0&1\end{bmatrix}\begin{bmatrix}\Delta\bar{y}_{t-1}\\ g_{t-1}\end{bmatrix}+\begin{bmatrix}u_{t}^{\pi}\\ u_{t}^{y}\\ u_{t}^{g}\end{bmatrix}$. (4.14)
Direct calculation (for further details on this and the subsequent calculations, see Appendix F) yields
$\Pi^{+}=\Pi^{-}$ $=\begin{bmatrix}0&0&0\\ 0&-1&1\\ 0&0&0\end{bmatrix}=\begin{bmatrix}0\\ 1\\ 0\end{bmatrix}\begin{bmatrix}0&-1&1\end{bmatrix}$, $\Pi^{x}$ $=\begin{bmatrix}0&0\\ -1&1\\ 0&0\end{bmatrix}$
so that $\operatorname{rk}\Pi^{+}=\operatorname{rk}\Pi^{-}=\operatorname{rk}\Pi^{x}=1$, indicating that the system falls within the purview of case (ii) with $\alpha=(0,1,0)^{\mathsf{T}}$ and $\beta^{+}=\beta^{-}=(0,-1,1)^{\mathsf{T}}$. It remains to verify the conditions of Theorem 4.2. We have immediately that CO(ii).1 holds with $r=1$. By calculating $\tilde{\Pi}^{\pm}$ for the associated canonical form of the model, we obtain that $1+\tilde{\beta}^{+\mathsf{T}}\tilde{\alpha}=1+\tilde{\beta}^{-\mathsf{T}}\tilde{\alpha}=0$, so (ii) is trivially satisfied. Finally, since in this first-order model $\Gamma^{\pm}(1)=\Phi_{0}^{\pm}$, it may be shown that $\det\alpha_{\perp}^{\mathsf{T}}\Gamma^{\pm}(1)\beta_{\perp}^{\pm}=\det I_{2}=1$, and thus CO(ii).3 holds. Since
$P_{\beta_{\perp}^{\pm}}=\beta_{\perp}^{\pm}[\alpha_{\perp}^{\mathsf{T}}\Gamma^{\pm}(1)\beta_{\perp}^{\pm}]^{-1}\alpha_{\perp}^{\mathsf{T}}=\begin{bmatrix}1&0\\ 0&1\\ 0&1\end{bmatrix}\begin{bmatrix}1&0&0\\ 0&0&1\end{bmatrix}$
is regime-invariant, it follows therefore by Theorem 4.2 that (supposing that all processes are initialised at zero, for simplicity)
$n^{-1/2}\begin{bmatrix}\bar{\pi}_{\lfloor n\lambda\rfloor}\\ \Delta\bar{y}_{\lfloor n\lambda\rfloor}\\ g_{\lfloor n\lambda\rfloor}\end{bmatrix}\rightsquigarrow\begin{bmatrix}1&0\\ 0&1\\ 0&1\end{bmatrix}\begin{bmatrix}U^{\pi}(\lambda)\\ U^{g}(\lambda)\end{bmatrix}$,
where $n^{-1/2}\sum_{t=1}^{\lfloor n\lambda\rfloor}(u_{t}^{\pi},u_{t}^{g})\rightsquigarrow[U^{\pi}(\lambda),U^{g}(\lambda)]$.
There are thus two common stochastic trends in the subsystem (4.12), one of which is shared between $\Delta\bar{y}_{t}$ and $g_{t}$; there is a single (linear) cointegrating relation given by the vector $\beta^{+}=\beta^{-}=(0,-1,1)^{\mathsf{T}}$, which eliminates these trends. It follows moreover from (4.13) that
$n^{-1/2}\bar{\imath}_{\lfloor n\lambda\rfloor}=n^{-1/2}(\bar{\pi}_{\lfloor n\lambda\rfloor}+cg_{\lfloor n\lambda\rfloor}+z_{\lfloor n\lambda\rfloor})\rightsquigarrow U^{\pi}(\lambda)+cU^{g}(\lambda)+U^{z}(\lambda),$
where $n^{-1/2}\sum_{t=1}^{\lfloor n\lambda\rfloor}u_{t}^{z}\rightsquigarrow U^{z}(\lambda)$, and so $\bar{\imath}_{t}$ does not cointegrate with $(\bar{\pi},\Delta\bar{y}_{t},g_{t})$.
4.3 Case (iii): linear cointegration in a nonlinear VECM
Finally, we consider the other case with a regime-invariant cointegrating rank $r$, but in which $\operatorname{rk}\Pi^{x}=r-1$. Then we have the factorisation
$\Pi^{\pm}=\begin{bmatrix}\pi^{\pm}&\Pi^{x}\end{bmatrix}=\begin{bmatrix}\pi^{\pm}&\alpha_{x}\end{bmatrix}\begin{bmatrix}1&0\\ 0&\beta_{x}^{\mathsf{T}}\end{bmatrix}\eqqcolon\alpha^{\pm}\beta^{\mathsf{T}},$ (4.15)
where $\alpha_{x}\in\mathbb{R}^{p\times(r-1)}$ and $\beta_{x}\in\mathbb{R}^{(p-1)\times(r-1)}$ have rank r1r-1, and $\alpha^{\pm},\beta\in\mathbb{R}^{p\times r}$ have rank $r$. (Note that the dimensions of $\beta_{x}$ here differ from those in case (ii).) The model is thus one in which the positive and negative regimes share a common cointegrating space, which contains $e_{p,1}$. Thus we would expect $y_{t}$ to behave like a stationary process: and indeed, the model falls within the very general framework of Saikkonen (2008), whose results could be applied to establish the stationarity and ergodicity of the equilibrium errors
$\xi_{t}\coloneqq\beta^{\mathsf{T}}z_{t}=\begin{bmatrix}y_{t}\\ \beta_{x}^{\mathsf{T}}x_{t}\end{bmatrix}$,
and hence of $y_{t}$. Our technical contribution here is to exploit the structure of the CKSVAR so as to permit his conditions, which refer to the JSR of a collection of autoregressive matrices, to be relaxed to merely requiring the stability of a certain deterministic subsystem (for which control over the JSR is a sufficient but not necessary condition). [^12] It should be emphasised that we do not claim to be relaxing the conditions of Saikkonen (2008, Thm. 1) for the general class of regime-switching error correction models considered in that paper. Rather, we are able to exploit the fact that our model (the CKSVAR configured as per (4.15)) is a special case of that framework to obtain weaker sufficient conditions for the ergodicity of $\beta^{\mathsf{T}}z_{t}$, in our setting.
Our only regularity condition on the system, in this case, is a stability condition of this kind. To present the system to which this applies, define
$\phi_{i}(y)$ $\coloneqq\phi_{i}^{+}\mathbf{1}^{+}(y)+\phi_{i}^{-}\mathbf{1}^{-}(y)$ $\Phi_{i}(y)$ $\coloneqq\begin{bmatrix}\phi_{i}(y)&\Phi_{i}^{x}\end{bmatrix}$
for $i\in\{1,\ldots,k\}$, $\boldsymbol{y}_{t-1}\coloneqq(y_{t-1},\ldots,y_{t-k})^{\mathsf{T}}$, and recognise that the factorisation (4.15) applies more generally to
$\Pi(\boldsymbol{y}_{t-1})\coloneqq\sum_{i=1}^{k}\Phi(y_{t-i})-I_{p}=\begin{bmatrix}\sum_{i=1}^{k}\phi_{i}(y_{t-i})-e_{1}&\alpha_{x}\end{bmatrix}\beta^{\mathsf{T}}\eqqcolon\alpha(\boldsymbol{y}_{t-1})\beta^{\mathsf{T}}.$
Further, define $\boldsymbol{\alpha}(\boldsymbol{y}_{t-1}),\boldsymbol{\beta}\in\mathbb{R}^{kp\times[r+(k-1)p]}$ as
$\boldsymbol{\alpha}(\boldsymbol{y}_{t-1})$ $\coloneqq\begin{bmatrix}\alpha(\boldsymbol{y}_{t-1})&\Gamma_{1}(\boldsymbol{y}_{t-1})&\cdots&\Gamma_{k-1}(\boldsymbol{y}_{t-1})\\ &I_{p}\\ &&\ddots\\ &&&I_{p}\end{bmatrix}$, $\boldsymbol{\beta}^{\mathsf{T}}$ $\coloneqq\begin{bmatrix}\beta^{\mathsf{T}}\\ I_{p}&-I_{p}\\ &\ddots&\ddots\\ &&I_{p}&-I_{p}\end{bmatrix}$, (4.16)
for
$\Gamma_{i}(\boldsymbol{y}_{t-1})\coloneqq-\sum_{j=i+1}^{k}\Phi_{j}(y_{t-j})$ (4.17)
all of which depend only on the signs of the elements of $\boldsymbol{y}_{t-1}$. Collect the short memory components of the model as
$\chi_{t}\coloneqq\begin{bmatrix}\xi_{t}\\ \beta_{\perp}^{\mathsf{T}}\Delta z_{t}\end{bmatrix}=\begin{bmatrix}y_{t}\\ \beta_{x}^{\mathsf{T}}x_{t}\\ \beta_{x,\perp}^{\mathsf{T}}\Delta x_{t}\end{bmatrix}$
for $\beta_{x,\perp}\in\mathbb{R}^{p-1}$ having $\operatorname{rk}\beta_{x,\perp}=p-r$, such that $\beta_{x,\perp}^{\mathsf{T}}\beta_{x}=0$. Letting $\boldsymbol{\chi}_{t}\coloneqq(\chi_{t}^{\mathsf{T}},\ldots,\chi_{t-k+1}^{\mathsf{T}})^{\mathsf{T}}$, we show in the proof of Theorem 4.3 below that
$\chi_{t}=Bc+M[I_{p(k-1)+r}+\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{\alpha}(\boldsymbol{G}\boldsymbol{\chi}_{t-1})]\boldsymbol{H}\boldsymbol{\chi}_{t-1}+Bu_{t}$
for $B\in\mathbb{R}^{p\times p}$ invertible, $M\in\mathbb{R}^{p\times[p(k-1)+r]}$, $\boldsymbol{G}\in\mathbb{R}^{k\times pk}$ and $\boldsymbol{H}\in\mathbb{R}^{[p(k-1)+r]\times kp}$. $\{\boldsymbol{\chi}_{t}\}$ thus evolves according to a regime-switching VAR, a Markov process that will be stationary and geometrically ergodic under the conditions given below.
The first of these conditions relates to the innovations $\{u_{t}\}$: for technical reasons, requiring that $u_{t}$ have a (conditional) Lebesgue density that is bounded away from zero, an assumption that is common in the literature on ergodic Markov processes, greatly facilitates our analysis.
Assumption ERR.
$\{u_{t}\}$ is i.i.d. with a Lebesgue density that is bounded away from zero on compact subsets of $\mathbb{R}^{p}$, $\mathbb{E}u_{t}=0$, and $\mathbb{E}\lVert u_{t}\rVert^{m_{0}}<\infty$ for some $m_{0}\geq 1$.
Our main condition on the model parameters is the following.
Assumption CO(iii).
  1. 1.
    $r^{+}=r^{-}=r$ and $\operatorname{rk}\Pi^{x}=r-1$; and
  2. 2.
    the deterministic system:
    $\hat{\boldsymbol{\xi}}_{t}=(I_{p(k-1)+r}+\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{\alpha}(\boldsymbol{G}\hat{\boldsymbol{\xi}}_{t-1}))\hat{\boldsymbol{\xi}}_{t-1}$
    is stable in the sense that $\hat{\boldsymbol{\xi}}_{t}\rightarrow 0$ for every initialisation $\hat{\boldsymbol{\xi}}_{0}\in\mathbb{R}^{p(k-1)+r}$.
Finally, to state our main result for case (iii), we recall the following (cf. Liebscher, 2005, p. 671; Meitz and Saikkonen, 2008, pp. 460f.).
Definition 4.1.
Let $\{w_{t}\}_{t\in\mathbb{N}_{0}}$ be a Markov chain taking values in $\mathbb{R}^{d_{w}}$, with $m$-step transition kernel $P^{m}(w,A)\coloneqq\mathbb{P}\{w_{t+m}\in A\mid w_{t}=w\}$, and $\mathcal{Q}:\mathbb{R}^{d_{w}}\rightarrow\mathbb{R}_{+}$. We say that $\{w_{t}\}$ is $\mathcal{Q}$-geometrically ergodic, with stationary distribution $\pi$, if $\int_{\mathbb{R}^{d_{w}}}\mathcal{Q}(w)\pi(\mathrm{d}w)<\infty$, and there exist $a,b>0$ and $\gamma\in(0,1)$ such that
$\sup_{B\in\mathcal{B}}\lvert P^{m}(w,B)-\pi(B)\rvert\leq(a+b\mathcal{Q}(w))\gamma^{m}$
for all $w\in\mathbb{R}^{d_{w}}$, where $\mathcal{B}$ denotes the Borel sigma-field on $\mathbb{R}^{d_{w}}$.
If $\{w_{t}\}$ is $\mathcal{Q}$-geometrically ergodic, it will be stationary if given a stationary initialisation, i.e. if $w_{0}$ also has distribution $\pi$; moreover, it will have geometrically decaying $\beta$-mixing coefficients. For these and further properties, and a discussion of how this concept relates to other notions of ergodicity used in the literature, see Liebscher (2005, pp. 671–73).
Theorem 4.3.
Suppose DGP, ERR, CVAR and CO(iii) hold. Then $\{\boldsymbol{\chi}_{t}\}$ is $\mathcal{Q}$-geometrically ergodic, for $\mathcal{Q}(\boldsymbol{\chi})\coloneqq 1+\lVert\boldsymbol{\chi}\rVert^{m_{0}}$.
Remark 4.3.
(i). If DGP is replaced by DGP, then CO(iii).2 should be replaced by
  1. CO(iii).2
    the following deterministic system is stable:
    $\hat{\boldsymbol{\xi}}_{t}=(I_{p(k-1)+r}+\tilde{\boldsymbol{\beta}}^{\mathsf{T}}\tilde{\boldsymbol{\alpha}}(\boldsymbol{G}\hat{\boldsymbol{\xi}}_{t-1}))\hat{\boldsymbol{\xi}}_{t-1},$
where the tildes refer to the parameters of the canonical CKSVAR derived from the structural form via Proposition 2.1. The theorem then delivers the $\cal Q$}-geometric ergodicity of $\tilde{\boldsymbol{\chi}}_{t}=(\tilde{\chi}_{t}^{\mathsf{T}},\ldots,\tilde{\chi}_{t-k+1}^{\mathsf{T}})^{\mathsf{T}}$, where
$\tilde{\chi}_{t}\coloneqq\begin{bmatrix}\tilde{\beta}^{\mathsf{T}}\tilde{z}_{t}\\ \tilde{\beta}_{\perp}^{\mathsf{T}}\Delta\tilde{z}_{t}\end{bmatrix}$
is formed from the canonical parameters and variables. Since, as shown in Appendix D, we can write $\beta^{\mathsf{T}}z_{t}=(y_{t},(\beta_{x}^{\mathsf{T}}x_{t})^{\mathsf{T}})^{\mathsf{T}}$ and $\Delta x_{t}$ as measurable (indeed, Lipschitz continuous) functions of $\tilde{\beta}^{\mathsf{T}}\tilde{z}_{t}$ and $(\tilde{\chi}_{t},\tilde{\chi}_{t-1})$ respectively, these processes will inherit the stationarity and geometric $\beta$-mixing properties of $\boldsymbol{\chi}_{t}$ that are a corollary of $\mathcal{Q}$-geometric ergodicity.
(ii). Since $\{\Delta x_{t}\}$ is geometrically $\beta$-mixing, and $x_{t}=\sum_{s=1}^{t}\Delta x_{s}+x_{0}$, the preceding result may be used as a starting point for the derivation of the asymptotics of $n^{-1/2}x_{\lfloor n\lambda\rfloor}$, which should converge to a multivariate Brownian motion (possibly after linear detrending). However, as discussed in Saikkonen (2008, pp. 307f.), determining the rank of the long-run variance of $\{\Delta x_{t}\}$, and hence the rank of the limiting process, is non-trivial. While its rank is bounded above by q=prq=p-r, it need not be equal to $q$; guaranteeing the latter is likely to require further conditions on the model parameters. We leave this for future work.
4.4 Deterministic trends
To simplify the exposition of our results for cases (i) and (ii) above, we have so far maintained DET, which prevents the model from generating any common deterministic trends. Since this is likely to be restrictive in applications – there being many macroeconomic series that exhibit both stochastic and deterministic trends – it is important to clarify that this simplifying assumption may be almost entirely dispensed with.
The presence of deterministic trends complicates the analysis of cases (i) and (ii), because these deterministic trends will generally dominate any common stochastic trends, such that the limit theory developed in Theorems 4.1 and 4.2 no longer applies directly to the standardised process $Z_{n}(\lambda)\coloneqq n^{-1/2}z_{\lfloor n\lambda\rfloor}$. Nonetheless, as explained below, these results continue to provide an accurate description of the asymptotics of this process upon linear detrending. (Since we did not maintain DET in the context of case (iii), we have nothing to say about that case in this section.)
We therefore now contemplate a model in which DET is relaxed, i.e. in which it is no longer required that $c\in\operatorname{sp}\Pi^{+}\cap\Pi^{-}$. In a linear cointegrated VAR, an unrestricted intercept $c\in\mathbb{R}^{p}$ permits the model to impart deterministic trends to all elements of $z_{t}=(y_{t},x_{t}^{\mathsf{T}})^{\mathsf{T}}$, but in such a way that these deterministic trends are eliminated by the cointegrating relations (simultaneously with the common stochastic trends; see Johansen, 1995, Sec. 5.7). This remains true in the CKSVAR, with the caveat that if a deterministic trend is imparted to $y_{t}$, then this dominant drift component will asymptotically push $y_{t}$ so far into either the positive or negative region (depending on the sign of the drift) as to render the threshold nonlinearity at zero irrelevant – and so the series would come to be adequately described by a linear VAR.
Since we would only apply the CKSVAR when $\{y_{t}\}$ spent an appreciable portion of the sample on both sides of zero, it is appropriate to work within an asymptotic framework in which both regions continue to be visited with nonvanishing probability as $T\rightarrow\infty$. [^13] Here we maintain that the location of the threshold delimiting the regimes is normalised to zero; of course the argument continues to apply if that threshold were instead fixed at some other finite level. The cleanest way to ensure this is to restrict $c$ such that a deterministic trend may be imparted to $x_{t}$, but not to $y_{t}$. In the context of cases (i) and (ii), essentially the same condition is required, but is denoted slightly differently so as to be consistent with the notation used in the analysis of these two cases. (Recall the definitions of $P_{\beta_{\perp}^{+}}$ and $P_{\beta_{\perp}}(+1)$ appearing in (4.3) and (4.7) above.)
Assumption DET.
In case (i), $e_{1}^{\mathsf{T}}P_{\beta_{\perp}^{+}}c=0$; in case (ii), $e_{1}^{\mathsf{T}}P_{\beta_{\perp}}(+1)c=0$.
To state our results in terms of the structural form of the CKSVAR, we shall say that ‘CO(i) holds’ whenever CO(i) holds with CO(i).2 replaced by (ii), and mutatis mutandis for CO(ii) (see part (i) of Remarks 4.1 and 4.2 above). To deal with the possible presence of a deterministic trend in $x_{t}$, in case (i) we define the linearly detrended processes
$\begin{bmatrix}y_{t}^{d}\\ x_{t}^{d}\end{bmatrix}=z_{t}^{d}\coloneqq z_{t}-(P_{\beta_{\perp}^{+}}c)t$
noting that $y_{t}^{d}=y_{t}$ under DET; in case (ii), $z_{t}^{d}$ is defined similarly with $P_{\beta_{\perp}^{+}}$ replaced by $P_{\beta_{\perp}}(+1)$. Define $X_{n}^{d}(\lambda)\coloneqq n^{-1/2}x_{\lfloor n\lambda\rfloor}^{d}$ to be the standardised process corresponding to $x_{t}^{d}$. The only modification that then needs to be made to the conclusions of Theorems 4.1 and 4.2 above is that $X_{n}$ should be replaced by $X_{n}^{d}$.
Theorem 4.4.
Suppose that DGP, ERR, DET and CVAR hold. If additionally CO(i) (respectively CO(ii)) holds, then the conclusions of Theorem 4.1 (respectively Theorem 4.2) hold with with $X_{n}^{d}$ in place of $X_{n}$.
Remark 4.4.
(i). The preceding illustrates how common stochastic and deterministic trends may be simultaneously accomodated within a CKSVAR. In particular, because the deterministic trends enter $z_{t}$ through $(P_{\beta_{\perp}^{+}}c)t$ (or $P_{\beta_{\perp}}(+1)c]t$), it remains true that $\beta^{+\mathsf{T}}z_{t}\sim I^{\ast}(0)$ in case (i) and $\beta(y_{t})^{\mathsf{T}}z_{t}\sim I^{\ast}(0)$ in case (ii), since these transformations of $z_{t}$ now eliminate both its common stochastic and deterministic trends. These objects thus continue to describe the (nonlinear) cointegrating relations between the elements of $z_{t}$.
(ii). It would be possible to accommodate a deterministic trend in $y_{t}$ within our asymptotic framework, provided that this component is sufficiently small that it does not overwhelm the stochastic trend component in $y_{t}$, such that the vicinity of $y_{t}=0$ continues to be visited with non-negligible probability in the limit. Mathematically, this could be engineered by relaxing DET so as to permit $e_{1}^{\mathsf{T}}P_{\beta_{\perp}^{+}}c$ (or $e_{1}^{\mathsf{T}}P_{\beta_{\perp}}(+1)c$) to be nonzero but of order $n^{-1/2}$: in which case the conclusions of Theorem 4.4 would remain unaltered, except that the limiting processes would now also incorporate a drift term. (See Bykhovskaya and Duffy, 2024, for an illustration of the form that this takes in the univariate form of case (i).) Thus even if DET is not imposed on the model, so that $c$ is unrestricted, it is still possible to interpret $\beta^{+}$ and $\beta(y)$ in the manner suggested above, in cases (i) and (ii) respectively.
5 Conclusion
The CKSVAR provides a flexible yet tractable framework in which to structurally model vector time series subject to an occasionally binding constraint, such as the zero lower bound on interest rates, and more general threshold nonlinearities. Nonetheless, even that seemingly limited amount of nonlinearity radically changes the properties of the model relative to a linear VAR. When unit autoregressive roots are introduced into the model, it is able to accommodate varieties of long-run behaviour that cannot be generated within a linear VAR, such as nonlinear common stochastic trends (censored, regulated and kinked Brownian motions) and cointegrating relationships that may be regime-dependent. This is not merely a theoretical curiosity but rather something that, as our examples illustrate, allows the long-run properties of the model to carry useful identifying information on structural parameters, as might pertain e.g. to the relative effectiveness of unconventional monetary policy.
Our results provide a complete characterisation of the forms of nonlinear cointegration (between processes $\ast$-integrated of order one) generated by the CKSVAR. In deriving these, we have given the first treatment of how nonlinear cointegration, in the profound sense of nonlinear common stochastic trends and nonlinear cointegrating relations, may be systematically generated within a (nonlinear) VAR, and thus the first extension of the Granger–Johansen representation theorem to a nonlinearly cointegrated setting. The special structure of the CKSVAR makes this problem peculiarly tractable, while being flexible enough to generate interesting departures from linear cointegration. Our results indicate how progress may now be made in the analysis of more general nonlinear VARs with unit roots, while our representation theory provides the foundations for inference on cointegrating relations in the CKSVAR. Our findings with respect to these problems will be reported elsewhere; some initial results regarding inference in this setting, in the univariate case, are given in Bykhovskaya and Duffy (2024) and Duffy and Jiao (2025).
Appendix A Auxiliary lemmas
We here collect three elementary lemmas, whose proofs appear in Appendix C for completeness. The third records, for convenience, some properties of the linear cointegrated VAR which are closely related to Lemmas A.1 and A.2 in Hansen (2005).
Lemma A.1.
Let $\{v_{t}\}$, $\{A_{t}\}$, $\{B_{t}\}$ and $\{c_{t}\}$ be random sequences, respectively taking values in $\mathbb{R}^{d_{v}}$, $\mathbb{R}^{d_{w}\times d_{w}}$, $\mathbb{R}^{d_{w}\times d_{v}}$ and $\mathbb{R}^{d_{w}}$, where $t\in\mathbb{N}$. Suppose $\{w_{t}\}$ satisfies
$w_{t}=c_{t}+A_{t}w_{t-1}+B_{t}v_{t}$
for some given (random) $w_{0}$, and:
  1. (i)
    $A_{t}\in\mathcal{A}$, $B_{t}\in\mathcal{B}$ and $c_{t}\in{\cal C}$ for all $t\in\mathbb{N}$, where $\mathcal{A}$, $\mathcal{B}$ and $\cal C$} are bounded subsets of $\mathbb{R}^{d_{w}\times d_{w}}$, $\mathbb{R}^{d_{w}\times d_{v}}$ and $\mathbb{R}^{d_{w}}$ respectively, and $\rho_{{\scriptstyle\mathrm{JSR}}}(\mathcal{A})<1$;
  2. (ii)
    $m_{0}\geq 1$ is such that $\lVert w_{0}\rVert_{m_{0}}+\sup_{t\in\mathbb{N}}\lVert v_{t}\rVert_{m_{0}}<\infty$.
Then for each $\gamma\in(\rho_{{\scriptstyle\mathrm{JSR}}}(\mathcal{A}),1)$, there exists a $C<\infty$ such that for all $t\in\mathbb{N}$,
$\lVert w_{t}\rVert\leq C\left[\sum_{s=0}^{t-1}\gamma^{s}(1+\lVert v_{t-s}\rVert)+\gamma^{t}\lVert w_{0}\rVert\right],$
and for all $n\in\mathbb{N}$ and $m\in[1,m_{0}]$,
$\max_{1\leq t\leq n}\lVert w_{t}\rVert_{m}\leq C\left(1+\lVert w_{0}\rVert_{m}+\max_{1\leq t\leq n}\lVert v_{t}\rVert_{m}\right).$
Lemma A.2.
Suppose $c\in\mathbb{R}^{m}$, $d,v\in\mathbb{R}^{n}$, and $A,B_{1},B_{2}\in\mathbb{R}^{m\times n}$ have full column rank and are such that $B_{1}-B_{2}=cd^{\mathsf{T}}$ and $\det(A^{\mathsf{T}}B_{1})\cdot\det(A^{\mathsf{T}}B_{2})>0$. Then
  1. (i)
    there exists a $\mu>0$ such that $d^{\mathsf{T}}(A^{\mathsf{T}}B_{1})^{-1}=\mu d^{\mathsf{T}}(A^{\mathsf{T}}B_{2})^{-1}$; and
  2. (ii)
    if $d^{\mathsf{T}}(A^{\mathsf{T}}B_{1})^{-1}v=0$, then $(A^{\mathsf{T}}B_{1})^{-1}v=(A^{\mathsf{T}}B_{2})^{-1}v$.
Lemma A.3.
Suppose $\Phi(\lambda)\coloneqq I_{p}-\sum_{i=1}^{k}\Phi_{i}\lambda^{i}$ and $\Pi\coloneqq-\Phi(1)$ satisfy CVAR (i.e. without the $\pm$’ superscripts). Define
$\boldsymbol{\Pi}\coloneqq\begin{bmatrix}\Pi+\Gamma_{1}&-\Gamma_{1}+\Gamma_{2}&\cdots&-\Gamma_{k-1}\\ I_{p}&-I_{p}\\ &\ddots&\ddots\\ &&I_{p}&-I_{p}\end{bmatrix}$,
and let $\Gamma(\lambda)\coloneqq I_{p}-\sum_{i=1}^{k-1}\Gamma_{i}\lambda^{i}$ be such that $\Phi(\lambda)=\Phi(1)\lambda+\Gamma(\lambda)(1-\lambda)$. [^14] As per the footnote to (4.2), the definition of $\boldsymbol{\Pi}$ here should be read ‘row-wise’, so that for $m\in\{2,\ldots,k\}$, rows (m1)p+1(m-1)p+1 to $mp$ of $\boldsymbol{\Pi}$ are given by $0_{p\times(m-2)p},I_{p},-I_{p},0_{p\times(k-m)p}]$. The matrix definitions appearing subsequently in these appendices should be interpreted similarly. Then:
  1. (i)
    there exist $\alpha,\beta\in\mathbb{R}^{p\times r}$ with rank $r$, such that $\Pi=\alpha\beta^{\mathsf{T}}$ and $\boldsymbol{\Pi}=\boldsymbol{\alpha}\boldsymbol{\beta}^{\mathsf{T}}$, where
    $\boldsymbol{\alpha}$ $\coloneqq\begin{bmatrix}\alpha&\Gamma_{1}&\cdots&\Gamma_{k-1}\\ &I_{p}\\ &&\ddots\\ &&&I_{p}\end{bmatrix}$ $\boldsymbol{\beta}^{\mathsf{T}}$ $\coloneqq\begin{bmatrix}\beta^{\mathsf{T}}\\ I_{p}&-I_{p}\\ &\ddots&\ddots\\ &&I_{p}&-I_{p}\end{bmatrix}$;
  2. (ii)
    $I_{p(k-1)+r}+\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{\alpha}$ has its eigenvalues strictly inside the unit circle, and $\boldsymbol{\alpha}_{\perp}^{\mathsf{T}}\boldsymbol{\beta}_{\perp}=\alpha_{\perp}^{\mathsf{T}}\Gamma(1)\beta_{\perp}$ and $\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{\alpha}$ are full rank, where
    $\boldsymbol{\alpha}_{\perp}^{\mathsf{T}}$ $\coloneqq\alpha_{\perp}^{\mathsf{T}}\begin{bmatrix}I_{p}&-\Gamma_{1}&\cdots&-\Gamma_{k-1}\end{bmatrix}$ $\boldsymbol{\beta}_{\perp}^{\mathsf{T}}$ $\coloneqq\beta_{\perp}^{\mathsf{T}}\begin{bmatrix}I_{p}&I_{p}&\cdots&I_{p}\end{bmatrix}$; (A.1)
  3. (iii)
    $\boldsymbol{P}_{\boldsymbol{\beta}_{\perp}}\coloneqq\boldsymbol{\beta}_{\perp}[\boldsymbol{\alpha}_{\perp}^{\mathsf{T}}\boldsymbol{\beta}_{\perp}]^{-1}\boldsymbol{\alpha}_{\perp}^{\mathsf{T}}$ and $\boldsymbol{P}_{\boldsymbol{\alpha}}\coloneqq\boldsymbol{\alpha}[\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{\alpha}]^{-1}\boldsymbol{\beta}^{\mathsf{T}}$ are complementary projections, and the upper left $p\times p$ block of $\boldsymbol{P}_{\boldsymbol{\beta}_{\perp}}$ is $P_{\beta_{\perp}}\coloneqq\beta_{\perp}[\alpha_{\perp}^{\mathsf{T}}\Gamma(1)\beta_{\perp}]^{-1}\alpha_{\perp}^{\mathsf{T}}$.
Appendix B Proofs of theorems
B.1 Proof of Theorem 4.1
Defining
$v_{t}$ $\coloneqq u_{t}+(\pi^{-}-\pi^{+})y_{t-1}^{-}+\sum_{i=1}^{k-1}(\gamma_{i}^{-}-\gamma_{i}^{+})\Delta y_{t-i}^{-}=u_{t}+\sum_{i=1}^{k}(\phi_{i}^{-}-\phi_{i}^{+})y_{t-i}^{-}$ (B.1)
and recalling $z_{t}=(y_{t},x_{t}^{\mathsf{T}})^{\mathsf{T}}$, we may rewrite the model (3.2) as
$\Delta z_{t}$ $=c+\Pi^{+}z_{t-1}+\sum_{i=1}^{k-1}\Gamma_{i}^{+}\Delta z_{t-i}+v_{t}$
where $\Pi^{+}=[\pi^{+},\Pi^{x}]$ and $\Gamma_{i}^{+}=[\gamma_{i}^{+},\Gamma^{x}]$. Conformably defining
$\boldsymbol{\Pi}^{+}$ $\coloneqq\begin{bmatrix}\Pi^{+}+\Gamma_{1}^{+}&-\Gamma_{1}^{+}+\Gamma_{2}^{+}&\cdots&-\Gamma_{k-1}^{+}\\ I_{p}&-I_{p}\\ &\ddots&\ddots\\ &&I_{p}&-I_{p}\end{bmatrix}$ $\boldsymbol{c}$ $\coloneqq\begin{bmatrix}c\\ 0_{p}\\ \vdots\\ 0_{p}\end{bmatrix}$ $\boldsymbol{v}_{t}$ $\coloneqq\begin{bmatrix}v_{t}\\ 0_{p}\\ \vdots\\ 0_{p}\end{bmatrix}$ $\boldsymbol{z}_{t}$ $\coloneqq\begin{bmatrix}z_{t}\\ z_{t-1}\\ \vdots\\ z_{t-k+1}\end{bmatrix}$
so that $\boldsymbol{\Pi}^{+}=\boldsymbol{\alpha}^{+}\boldsymbol{\beta}^{+\mathsf{T}}$, where $\boldsymbol{\alpha}^{+}$ and $\boldsymbol{\beta}^{+}$ are as in (4.2), we render the system in companion form as
$\Delta\boldsymbol{z}_{t}=\boldsymbol{c}+\boldsymbol{\Pi}^{+}\boldsymbol{z}_{t-1}+\boldsymbol{v}_{t}.$ (B.2)
The next lemma provides a (nonlinear) VAR representation for the short-memory components, which comprise: (a) the equilibrium errors (using the cointegrating vectors $\beta^{+}$ from the positive regime) and (lagged) differences,
$\boldsymbol{\xi}_{t}^{+}\coloneqq\boldsymbol{\beta}^{+\mathsf{T}}\boldsymbol{z}_{t}=(\xi_{t}^{+\mathsf{T}},\Delta z_{t}^{\mathsf{T}},\ldots,\Delta z_{t-k+2}^{\mathsf{T}})^{\mathsf{T}}$
where $\xi_{t}^{+}\coloneqq\beta^{+\mathsf{T}}z_{t}$; (b) the lagged levels $\boldsymbol{y}_{t-1}^{-}\coloneqq(y_{t-1}^{-},\ldots,y_{t-k+1}^{-})^{\mathsf{T}}$; and (c) an auxiliary series $\text{$\overline{y}$}_{t}$. Collect these in $\boldsymbol{\zeta}_{t}\in\mathbb{R}^{p(k-1)+k+r}$, and conformably define
$\boldsymbol{\zeta}_{t}$ $=\begin{bmatrix}\boldsymbol{\xi}_{t}^{+}\\ \text{$\overline{y}$}_{t}\\ \boldsymbol{y}_{t-1}^{-}\end{bmatrix}$ $\boldsymbol{\varepsilon}_{t}$ $\coloneqq\begin{bmatrix}\boldsymbol{\beta}_{1:p}^{+\mathsf{T}}u_{t}\\ e_{1}^{\mathsf{T}}u_{t}\\ 0_{k-1}\end{bmatrix}$ $\boldsymbol{\mu}$ $\coloneqq\begin{bmatrix}\boldsymbol{\beta}_{1:p}^{+\mathsf{T}}c\\ e_{1}^{\mathsf{T}}c\\ 0_{k-1}\end{bmatrix}$
where $\boldsymbol{\beta}_{1:p}^{+}$ denotes the first $p$ rows of $\boldsymbol{\beta}^{+}$. Recall the definition of $\boldsymbol{F}_{\delta}$ given in (4.1) above.
Lemma B.1.
Suppose CVAR and CO(i).1 hold. Set $\delta_{0}\coloneqq 1$ and $\overline{y}_{0}\coloneqq y_{0}^{-}$. Then $\{\boldsymbol{\zeta}_{t}\}$ follows
$\boldsymbol{\zeta}_{t}$ $=\boldsymbol{\mu}+\boldsymbol{F}_{\delta_{t-1}}\boldsymbol{\zeta}_{t-1}+\boldsymbol{\varepsilon}_{t}$ (B.3)
$\delta_{t}$ $=\begin{cases}(y_{t-1}^{+}+\text{$\overline{y}$}_{t})/\text{$\overline{y}$}_{t}&\text{if }y_{t-1}^{+}+\text{$\overline{y}$}_{t}<0,\\ 0&\text{otherwise};\end{cases}$ (B.4)
which implies, in particular, that $y_{t}^{-}=\delta_{t}\text{$\overline{y}$}_{t}$ for all $t\in\mathbb{N}$.
Proof.
Premultiply (B.2) by $\boldsymbol{\beta}^{+\mathsf{T}}$ and then use (B.1) to obtain
$\boldsymbol{\xi}_{t}^{+}$ $=\boldsymbol{\beta}_{1:p}^{+\mathsf{T}}c+(I+\boldsymbol{\beta}^{+\mathsf{T}}\boldsymbol{\alpha}^{+})\boldsymbol{\xi}_{t-1}^{+}+\boldsymbol{\beta}_{1:p}^{+\mathsf{T}}v_{t}$
$=\boldsymbol{\beta}_{1:p}^{+\mathsf{T}}c+(I+\boldsymbol{\beta}^{+\mathsf{T}}\boldsymbol{\alpha}^{+})\boldsymbol{\xi}_{t-1}^{+}$
$\qquad\qquad+\boldsymbol{\beta}_{1:p}^{+\mathsf{T}}(\phi_{1}^{-}-\phi_{1}^{+})y_{t-1}^{-}+\boldsymbol{\beta}_{1:p}^{+\mathsf{T}}\sum_{i=2}^{k}(\phi_{i}^{-}-\phi_{i}^{+})y_{t-i}^{-}+\boldsymbol{\beta}_{1:p}^{+\mathsf{T}}u_{t}.$ (B.5)
To obtain the law of motion for $y_{t}^{-}$, we take the first equation in (B.2)
$\Delta y_{t}=e_{1}^{\mathsf{T}}c+\boldsymbol{e}_{1}^{\mathsf{T}}\boldsymbol{\alpha}^{+}\boldsymbol{\xi}_{t-1}^{+}+e_{1}^{\mathsf{T}}v_{t}$
(where $e_{1}$ and $\boldsymbol{e}_{1}$ denote the first column of $I_{p}$ and $I_{kp}$ respectively), which using (B.1) can be rearranged as
$y_{t}-y_{t-1}^{+}$ $=e_{1}^{\mathsf{T}}c+[1+e_{1}^{\mathsf{T}}(\phi_{1}^{-}-\phi_{1}^{+})]y_{t-1}^{-}$ (B.6)
$\qquad\qquad+\sum_{i=2}^{k}e_{1}^{\mathsf{T}}(\phi_{i}^{-}-\phi_{i}^{+})y_{t-i}^{-}+\boldsymbol{e}_{1}^{\mathsf{T}}\boldsymbol{\alpha}^{+}\boldsymbol{\xi}_{t-1}^{+}+e_{1}^{\mathsf{T}}u_{t}.$
It follows that if we define $\text{$\overline{y}$}_{t}$ to equal the r.h.s., then $y_{t}=y_{t-1}^{+}+\text{$\overline{y}$}_{t}$, and so
$y_{t}^{-}=[y_{t}]_{-}=[y_{t-1}^{+}+\text{$\overline{y}$}_{t}]_{-}=\delta_{t}\text{$\overline{y}$}_{t}$
where $\delta_{t}\in[0,1]$ is defined in (B.4). (Observe that if $y_{t-1}^{+}+\text{$\overline{y}$}_{t}<0$, then $\overline{y}_{t}<-y_{t-1}^{+}\leq 0$, so this object is well defined.) Making the substitution $y_{t-1}^{-}=\delta_{t-1}\text{$\overline{y}$}_{t-1}$ on the r.h.s. of (B.5) and (B.6), we see that the trajectories of $\{\boldsymbol{\xi}_{t}^{+}\}$ and $\{y_{t}^{-}\}$ for $t\geq 1$ (conditional on the initial values $\{z_{t}\}_{t=-k+1}^{0}$) will be reproduced exactly by
$\boldsymbol{\xi}_{t}^{+}$ $=\boldsymbol{\beta}_{1:p}^{+\mathsf{T}}c+(I+\boldsymbol{\beta}^{+\mathsf{T}}\boldsymbol{\alpha}^{+})\boldsymbol{\xi}_{t-1}^{+}$ (B.7a)
$\qquad\qquad\qquad+\boldsymbol{\beta}_{1:p}^{+\mathsf{T}}(\phi_{1}^{-}-\phi_{1}^{+})\delta_{t-1}\text{$\overline{y}$}_{t-1}+\boldsymbol{\beta}_{1:p}^{+\mathsf{T}}\sum_{i=2}^{k}(\phi_{i}^{-}-\phi_{i}^{+})y_{t-i}^{-}+\boldsymbol{\beta}_{1:p}^{+\mathsf{T}}u_{t},$
$\text{$\overline{y}$}_{t}$ $=e_{1}^{\mathsf{T}}c+[1+e_{1}^{\mathsf{T}}(\phi_{1}^{-}-\phi_{1}^{+})]\delta_{t-1}\text{$\overline{y}$}_{t-1}$
$\qquad\qquad\qquad+\sum_{i=2}^{k}e_{1}^{\mathsf{T}}(\phi_{i}^{-}-\phi_{i}^{+})y_{t-i}^{-}+\boldsymbol{e}_{1}^{\mathsf{T}}\boldsymbol{\alpha}^{+}\boldsymbol{\xi}_{t-1}^{+}+e_{1}^{\mathsf{T}}u_{t},$ (B.7b)
$y_{t-1}^{-}$ $=\delta_{t-1}\text{$\overline{y}$}_{t-1},$ (B.7c)
with $\delta_{0}\coloneqq 1$ and $\overline{y}_{0}\coloneqq y_{0}^{-}$. Finally, we note that (B.3) provides the companion form representation of this system. ∎
By Jungers (2009, Prop. 1.8), CO(i).2 implies that $\rho_{{\scriptstyle\mathrm{JSR}}}(\{\boldsymbol{F}_{\delta}\mid\delta\in[0,1]\})=\rho_{{\scriptstyle\mathrm{JSR}}}(\{\boldsymbol{F}_{0},\boldsymbol{F}_{1}\})<1$. In view of ERR and CO(i).4, applying Lemma A.1 to the representation (B.3) then yields that $\sup_{t\in\mathbb{N}}\lVert\boldsymbol{\zeta}_{t}\rVert_{2+\delta_{u}}<\infty$, and hence $\boldsymbol{\zeta}_{t}\sim I^{\ast}(0)$. In particular, $\boldsymbol{\xi}_{t}^{+}\sim I^{\ast}(0)$ and $y_{t}^{-}\sim I^{\ast}(0)$, which gives part (i) of the theorem.
For the purposes of proving part (ii), we shall for the moment we replace DET by the weaker condition
$e_{1}^{\mathsf{T}}P_{\beta_{\perp}^{+}}c=0,$ (B.8)
for $P_{\beta_{\perp}^{+}}$ as defined in (4.3), noting only at the very end of the proof how the result simplifies when DET holds. As discussed in Section 4.4, the condition (B.8) (which appears there as DET) permits the model to impart a deterministic trend to $x_{t}$, but not to $y_{t}$.
Our first step is to extract the common trend components. Define the $q\times kp$ matrices
$\boldsymbol{\alpha}_{\perp}^{+\mathsf{T}}$ $\coloneqq\alpha_{\perp}^{+\mathsf{T}}\begin{bmatrix}I_{p}&-\Gamma_{1}^{+}&\cdots&-\Gamma_{k-1}^{+}\end{bmatrix}$ $\boldsymbol{\beta}_{\perp}^{+\mathsf{T}}$ $\coloneqq\beta_{\perp}^{+\mathsf{T}}\begin{bmatrix}I_{p}&I_{p}&\cdots&I_{p}\end{bmatrix}$,
which are orthogonal to $\boldsymbol{\alpha}^{+}$ and $\boldsymbol{\beta}^{+}$, and the complementary projections
$\boldsymbol{P}_{\boldsymbol{\beta}_{\perp}^{+}}$ $\coloneqq\boldsymbol{\beta}_{\perp}^{+}[\boldsymbol{\alpha}_{\perp}^{+\mathsf{T}}\boldsymbol{\beta}_{\perp}^{+}]^{-1}\boldsymbol{\alpha}_{\perp}^{+\mathsf{T}}$ $\boldsymbol{P}_{\boldsymbol{\alpha}^{+}}$ $\coloneqq\boldsymbol{\alpha}^{+}[\boldsymbol{\beta}^{+\mathsf{T}}\boldsymbol{\alpha}^{+}]^{-1}\boldsymbol{\beta}^{+\mathsf{T}}$
(see Lemma A.3). Premutiplying (B.2) by $\boldsymbol{\alpha}_{\perp}^{+\mathsf{T}}$ and cumulating yields
$\boldsymbol{\alpha}_{\perp}^{+\mathsf{T}}\boldsymbol{z}_{t}$ $=\boldsymbol{\alpha}_{\perp}^{+\mathsf{T}}\boldsymbol{z}_{0}+\boldsymbol{\alpha}_{\perp}^{+\mathsf{T}}\sum_{s=1}^{t}\boldsymbol{v}_{s}+(\boldsymbol{\alpha}_{\perp}^{+\mathsf{T}}\boldsymbol{c})t.$
Hence
$\boldsymbol{z}_{t}=(\boldsymbol{P}_{\boldsymbol{\beta}_{\perp}^{+}}+\boldsymbol{P}_{\boldsymbol{\alpha}^{+}})\boldsymbol{z}_{t}=\boldsymbol{P}_{\boldsymbol{\beta}_{\perp}^{+}}\left(\boldsymbol{z}_{0}+\sum_{s=1}^{t}\boldsymbol{v}_{s}+\boldsymbol{c}t\right)+\boldsymbol{z}_{\xi,t}$
where $\boldsymbol{z}_{\xi,t}\coloneqq\boldsymbol{P}_{\boldsymbol{\alpha}^{+}}\boldsymbol{z}_{t}=\boldsymbol{\alpha}^{+}[\boldsymbol{\beta}^{+\mathsf{T}}\boldsymbol{\alpha}^{+}]^{-1}\boldsymbol{\xi}_{t}^{+}\sim I^{\ast}(0)$.
Since the top left $p\times p$ block of $\boldsymbol{P}_{\boldsymbol{\beta}_{\perp}^{+}}$ is $P_{\beta_{\perp}^{+}}=\beta_{\perp}^{+}[\alpha_{\perp}^{+\mathsf{T}}\Gamma^{+}(1)\beta_{\perp}^{+}]^{-1}\alpha_{\perp}^{+\mathsf{T}}$ (Lemma A.3), the preceding implies
$z_{t}$ $=[\boldsymbol{P}_{\boldsymbol{\beta}_{\perp}^{+}}\boldsymbol{z}_{0}]_{1:p}+P_{\beta_{\perp}^{+}}\sum_{s=1}^{t}v_{s}+(P_{\beta_{\perp}^{+}}c)t+z_{\xi,t}$ (B.9)
where $z_{\xi,t}\coloneqq[\boldsymbol{z}_{\xi,t}]_{1:p}$. Under CO(i).4, $n^{-1/2}z_{t}\overset{p}{\rightarrow}{\cal Z}_{0}$ for each $t\in\{-k+1,\ldots,0\}$; hence
$n^{-1/2}[\boldsymbol{P}_{\boldsymbol{\beta}_{\perp}^{+}}\boldsymbol{z}_{0}]_{1:p}\overset{p}{\rightarrow}\beta_{\perp}^{+}[\alpha_{\perp}^{+\mathsf{T}}\Gamma^{+}(1)\beta_{\perp}^{+}]^{-1}\alpha_{\perp}^{+\mathsf{T}}\Gamma^{+}(1){\cal Z}_{0}={\cal Z}_{0}$ (B.10)
where the final equality follows since ${\cal Z}_{0}\in\mathscr{M}\subset\operatorname{sp}\beta_{\perp}^{+}$. Now recalling (B.1),
$\sum_{s=1}^{t}v_{s}$ $=\sum_{s=1}^{t}u_{s}+(\pi^{-}-\pi^{+})\sum_{s=0}^{t-1}y_{s}^{-}+\sum_{i=1}^{k-1}(\gamma_{i}^{-}-\gamma_{i}^{+})[y_{t-i}^{-}-y_{-i}^{-}],$
we have from (B.9) that
$z_{t}=[\boldsymbol{P}_{\boldsymbol{\beta}_{\perp}^{+}}\boldsymbol{z}_{0}]_{1:p}+(P_{\beta_{\perp}^{+}}c)t+P_{\beta_{\perp}^{+}}\sum_{s=1}^{t}u_{s}+\kappa\sum_{s=0}^{t-1}y_{s}^{-}+P_{\beta_{\perp}^{+}}\sum_{i=1}^{k-1}(\gamma_{i}^{-}-\gamma_{i}^{+})[y_{t-i}^{-}-y_{-i}^{-}]+z_{\xi,t},$
where $\kappa\coloneqq P_{\beta_{\perp}^{+}}(\pi^{-}-\pi^{+})=P_{\beta_{\perp}^{+}}\pi^{-}$. Using that $z_{t}=[\begin{smallmatrix}y_{t}\\ x_{t}\end{smallmatrix}]=[\begin{smallmatrix}y_{t}^{+}\\ x_{t}\end{smallmatrix}]+e_{1}y_{t}^{-}$, and defining
$\eta_{t}\coloneqq-(e_{1}+\kappa)y_{t}^{-}+P_{\beta_{\perp}^{+}}\sum_{i=1}^{k-1}(\gamma_{i}^{-}-\gamma_{i}^{+})[y_{t-i}^{-}-y_{-i}^{-}]+z_{\xi,t},$
which is $I^{\ast}(0)$ by part (i) of the theorem, we may rewrite the preceding as
$\begin{bmatrix}y_{t}^{+}\\ x_{t}\end{bmatrix}$ $=[\boldsymbol{P}_{\boldsymbol{\beta}_{\perp}^{+}}\boldsymbol{z}_{0}]_{1:p}+(P_{\beta_{\perp}^{+}}c)t+\kappa\sum_{s=0}^{t}y_{s}^{-}+P_{\beta_{\perp}^{+}}\sum_{s=1}^{t}u_{s}+\eta_{t}$ (B.11)
Under (B.8), the first equation in (B.11) is
$y_{t}^{+}$ $=e_{1}^{\mathsf{T}}[\boldsymbol{P}_{\boldsymbol{\beta}_{\perp}^{+}}\boldsymbol{z}_{0}]_{1:p}+\kappa_{1}\sum_{s=0}^{t}y_{s}^{-}+e_{1}^{\mathsf{T}}P_{\beta_{\perp}^{+}}\sum_{s=1}^{t}u_{s}+\eta_{1,t}$ (B.12)
where $\eta_{1,t}=e_{1}^{\mathsf{T}}\eta_{t}$. Taking first differences yields
$y_{t}^{+}-\kappa_{1}y_{t}^{-}=y_{t-1}^{+}+e_{1}^{\mathsf{T}}P_{\beta_{\perp}^{+}}u_{t}+\Delta\eta_{1,t}.$
Since $\kappa_{1}<0$ by CO(i).3, only one of $y_{t}^{+}$ and $\kappa y_{t}^{-}$ can be nonzero, and must have opposite signs; hence
$y_{t}^{+}=[y_{t-1}^{+}+e_{1}^{\mathsf{T}}P_{\beta_{\perp}^{+}}u_{t}+\Delta\eta_{1,t}]_{+}.$
As noted above $\eta_{t}\sim I^{\ast}(0)$, while $n^{-1/2}y_{0}\overset{p}{\rightarrow}{\cal Y}_{0}$ by CO(i).4. Hence by ERR,
$n^{-1/2}y_{0}+n^{-1/2}\sum_{t=1}^{\lfloor n\lambda\rfloor}(e_{1}^{\mathsf{T}}P_{\beta_{\perp}^{+}}u_{t}+\Delta\eta_{t})$
$\qquad\qquad={\cal Y}_{0}+n^{-1/2}e_{1}^{\mathsf{T}}P_{\beta_{\perp}^{+}}\sum_{t=1}^{\lfloor n\lambda\rfloor}u_{t}+n^{-1/2}(\eta_{1,\lfloor n\lambda\rfloor}-\eta_{1,0})+o_{p}(1)$
$\qquad\qquad\rightsquigarrow{\cal Y}_{0}+e_{1}^{\mathsf{T}}P_{\beta_{\perp}^{+}}U(\lambda)=e_{1}^{\mathsf{T}}P_{\beta_{\perp}^{+}}[\Gamma^{+}(1){\cal Z}_{0}+U(\lambda)]=e_{1}^{\mathsf{T}}P_{\beta_{\perp}^{+}}U_{0}(\lambda).$
on $D[0,1]$, where the penultimate equality follows from $P_{\beta_{\perp}^{+}}\Gamma^{+}(1){\cal Z}_{0}={\cal Z}_{0}$, as per (B.10) above. Therefore by Bykhovskaya and Duffy (2024, Lem. A.1), and $y_{t}^{-}\sim I^{\ast}(0)$,
$n^{-1/2}y_{\lfloor n\lambda\rfloor}$ $=n^{-1/2}y_{\lfloor n\lambda\rfloor}^{+}+o_{p}(1)$
$\rightsquigarrow e_{1}^{\mathsf{T}}P_{\beta_{\perp}^{+}}U_{0}(\lambda)+\sup_{\lambda^{\prime}\leq\lambda}[-e_{1}^{\mathsf{T}}P_{\beta_{\perp}^{+}}U_{0}(\lambda^{\prime})]_{+}=Y(\lambda).$ (B.13)
Since (B.8) permits a deterministic trend to be present in $x_{t}$, as evident from (B.11) above, we consider the linearly detrended process
$x_{t}^{d}\coloneqq x_{t}-(E_{-1}^{\mathsf{T}}P_{\beta_{\perp}^{+}}c)t$
where $E_{-1}\in\mathbb{R}^{p\times(p-1)}$ collects the final p1p-1 columns of $I_{p}$. To obtain the weak limit of $n^{-1/2}x_{\lfloor n\lambda\rfloor}^{d},$ we substitute
$\sum_{s=0}^{t}y_{s}^{-}=\kappa_{1}^{-1}y_{t}^{+}-\kappa_{1}^{-1}e_{1}^{\mathsf{T}}\left([\boldsymbol{P}_{\boldsymbol{\beta}_{\perp}^{+}}\boldsymbol{z}_{0}]_{1:p}+P_{\beta_{\perp}^{+}}\sum_{s=1}^{t}u_{s}+\eta_{t}\right)$
from (B.12) into the final p1p-1 equations from (B.11), to obtain
$x_{t}^{d}$ $=x_{t}-(E_{-1}^{\mathsf{T}}P_{\beta_{\perp}^{+}}c)t$
$=E_{-1}^{\mathsf{T}}\left\{[\boldsymbol{P}_{\boldsymbol{\beta}_{\perp}^{+}}\boldsymbol{z}_{0}]_{1:p}+\kappa\sum_{s=0}^{t}y_{s}^{-}+P_{\beta_{\perp}^{+}}\sum_{s=1}^{t}u_{s}+\eta_{t}\right\}$
$=E_{-1}^{\mathsf{T}}\left\{(I_{p}-\kappa_{1}^{-1}\kappa e_{1}^{\mathsf{T}})\left[[\boldsymbol{P}_{\boldsymbol{\beta}_{\perp}^{+}}\boldsymbol{z}_{0}]_{1:p}+P_{\beta_{\perp}^{+}}\sum_{s=1}^{t}u_{s}+\eta_{t}\right]+\kappa_{1}^{-1}\kappa y_{t}^{+}\right\}.$
for $E_{-1}$ the final p1p-1 columns of $I_{p}$. Hence, by $\eta_{t}\sim I^{\ast}(0)$, ERR, (B.10) and (B.13),
$n^{-1/2}x_{\lfloor n\lambda\rfloor}^{d}$ $=E_{-1}^{\mathsf{T}}\left\{(I_{p}-\kappa_{1}^{-1}\kappa e_{1}^{\mathsf{T}})\left[{\cal Z}_{0}+P_{\beta_{\perp}^{+}}n^{-1/2}\sum_{s=1}^{\lfloor n\lambda\rfloor}u_{s}\right]+\kappa_{1}^{-1}\kappa n^{-1/2}y_{\lfloor n\lambda\rfloor}^{+}\right\}+o_{p}(1)$
$\rightsquigarrow E_{-1}^{\mathsf{T}}\left\{(I_{p}-\kappa_{1}^{-1}\kappa e_{1}^{\mathsf{T}})P_{\beta_{\perp}^{+}}\left[\Gamma^{+}(1){\cal Z}_{0}+U(\lambda)\right]+\kappa_{1}^{-1}\kappa Y(\lambda)\right\}$
$=E_{-1}^{\mathsf{T}}\left\{P_{\beta_{\perp}^{+}}U_{0}(\lambda)+\kappa_{1}^{-1}\kappa\sup_{\lambda^{\prime}\leq\lambda}[-e_{1}^{\mathsf{T}}P_{\beta_{\perp}^{+}}U_{0}(\lambda^{\prime})]_{+}\right\}$ (B.14)
on $D[0,1]$.
Finally, we suppose that DET holds in place of (B.8). In this case, $P_{\beta_{\perp}^{+}}c=0$ and therefore $x_{t}^{d}=x_{t}$ for all $t\in\{1,\ldots,T\}$. Thus part (ii) of the theorem follows from (B.13) and (B.14). ∎
B.2 Proof of Theorem 4.2
Because $\{y_{t}\}$ now behaves like an integrated process in both the positive and negative regimes, we have to take quite a different approach from that utilised in case (i). Recall $z_{t}^{\mathsf{T}}=(y_{t},x_{t}^{\mathsf{T}})$, but now also define $z_{t}^{\ast\mathsf{T}}\coloneqq(y_{t}^{+},y_{t}^{-},x_{t}^{\mathsf{T}})$, so that (3.2) may be written as
$\Delta z_{t}=\begin{bmatrix}\Delta y_{t}\\ \Delta x_{t}\end{bmatrix}$ $=c+\begin{bmatrix}\pi^{+}&\pi^{-}&\Pi^{x}\end{bmatrix}\begin{bmatrix}y_{t-1}^{+}\\ y_{t-1}^{-}\\ x_{t-1}\end{bmatrix}+\sum_{i=1}^{k-1}\begin{bmatrix}\gamma_{i}^{+}&\gamma_{i}^{-}&\Gamma_{i}^{x}\end{bmatrix}\begin{bmatrix}\Delta y_{t-i}^{+}\\ \Delta y_{t-i}^{-}\\ \Delta x_{t-i}\end{bmatrix}+u_{t}$
$=c+\Pi(y_{t-1})z_{t-1}+\sum_{i=1}^{k-1}\Gamma_{i}\Delta z_{t-i}^{\ast}+u_{t},$
where $\Pi(y)\coloneqq\Pi^{+}\mathbf{1}^{+}(y)+\Pi^{-}\mathbf{1}^{-}(y)$. Taking $\boldsymbol{z}_{t}^{\mathsf{T}}\coloneqq(z_{t}^{\mathsf{T}},z_{t-1}^{\ast\mathsf{T}},\ldots,z_{t-k+1}^{\ast\mathsf{T}})$ – the first $p$ elements of which are $z_{t}$, not $z_{t}^{\ast}$ – as the state vector, and conformably defining
$\boldsymbol{\Pi}(y)$ $\coloneqq\begin{bmatrix}\Pi(y)+\Gamma_{1}S(y)&-\Gamma_{1}+\Gamma_{2}&-\Gamma_{2}+\Gamma_{3}&\cdots&-\Gamma_{k-1}\\ S(y)&-I_{p+1}\\ &I_{p+1}&-I_{p+1}\\ &&\ddots&\ddots\\ &&&I_{p+1}&-I_{p+1}\end{bmatrix}$ $\begin{aligned} \boldsymbol{c}&\coloneqq\begin{bmatrix}c\\ 0_{(p+1)(k-1)}\end{bmatrix}\\ \\ \boldsymbol{u}_{t}&\coloneqq\begin{bmatrix}u_{t}\\ 0_{(p+1)(k-1)}\end{bmatrix}\end{aligned}$
so that $\boldsymbol{\Pi}(y)=\boldsymbol{\alpha}\boldsymbol{\beta}(y)^{\mathsf{T}}$, where $\boldsymbol{\alpha}$, $\boldsymbol{\beta}(y)$ and $S(y)$ are as defined in (4.9)–(4.10), we can rewrite the system in companion form as
$\Delta\boldsymbol{z}_{t}=\boldsymbol{c}+\boldsymbol{\Pi}(y_{t-1})\boldsymbol{z}_{t-1}+\boldsymbol{u}_{t}.$ (B.15)
Note the definitions of $\boldsymbol{z}_{t}$, $\boldsymbol{c}$ and $\boldsymbol{u}_{t}$ here differ from those in case (i).
Our next result is the counterpart of Lemma B.1 for the present case, in that it provides a nonlinear VAR representation for the short-memory components of the model: the equilibrium errors $\xi_{t}\coloneqq\beta(y_{t})^{\mathsf{T}}z_{t}$ and the differences $\Delta z_{t}^{\ast}$, as collected in
$\boldsymbol{\xi}_{t}\coloneqq\boldsymbol{\beta}(y_{t})^{\mathsf{T}}\boldsymbol{z}_{t}=(\xi_{t}^{\mathsf{T}},\Delta z_{t}^{\ast\mathsf{T}},\ldots,\Delta z_{t-k+2}^{\ast\mathsf{T}})^{\mathsf{T}}.$
Lemma B.2.
Suppose that DGP, CVAR and CO(ii).1 hold. Let
$\delta_{t}\coloneqq\begin{cases}0&\text{if }y_{t-1}\cdot y_{t}\geq 0,\\ y_{t}/\Delta y_{t}&\text{if }y_{t-1}\cdot y_{t}<0\end{cases}$ (B.16)
for $t\in\mathbb{N}$, which takes values only in $[0,1]$, and define $\overline{\boldsymbol{\beta}}_{t}\coloneqq(1-\delta_{t})\boldsymbol{\beta}(y_{t-1})+\delta_{t}\boldsymbol{\beta}(y_{t})$. Then $\{\boldsymbol{\xi}_{t}\}_{t\in\mathbb{N}}$ satisfies
$\boldsymbol{\xi}_{t}=\overline{\boldsymbol{\beta}}_{t}^{\mathsf{T}}\boldsymbol{c}+[I_{r+(k-1)(p+1)}+\overline{\boldsymbol{\beta}}_{t}^{\mathsf{T}}\boldsymbol{\alpha}]\boldsymbol{\xi}_{t-1}+\overline{\boldsymbol{\beta}}_{t}^{\mathsf{T}}\boldsymbol{u}_{t}.$ (B.17)
Proof.
Using the definition of $\boldsymbol{\xi}_{t}$, we may write (B.15) as
$\Delta\boldsymbol{z}_{t}=\boldsymbol{c}+\boldsymbol{\alpha}\boldsymbol{\xi}_{t-1}+\boldsymbol{u}_{t}.$ (B.18)
Let $\Delta\boldsymbol{\beta}(y_{t})\coloneqq\boldsymbol{\beta}(y_{t})-\boldsymbol{\beta}(y_{t-1})$, and note that $\Delta\boldsymbol{\beta}(y_{t})]^{\mathsf{T}}\boldsymbol{z}_{t}=[\Delta\boldsymbol{\beta}(y_{t})]_{1:1}^{\mathsf{T}}y_{t}$, for $\Delta\boldsymbol{\beta}(y_{t})]_{1:1}$ the first row of $\Delta\boldsymbol{\beta}(y_{t})$, since all its other rows are zero. Thus with the aid of (B.18) we obtain
$\boldsymbol{\xi}_{t}$ $=\boldsymbol{\beta}(y_{t})^{\mathsf{T}}\boldsymbol{z}_{t}=\boldsymbol{\beta}(y_{t-1})^{\mathsf{T}}(\boldsymbol{z}_{t-1}+\Delta\boldsymbol{z}_{t})+[\Delta\boldsymbol{\beta}(y_{t})]^{\mathsf{T}}\boldsymbol{z}_{t}$
$=\boldsymbol{\xi}_{t-1}+\boldsymbol{\beta}(y_{t-1})^{\mathsf{T}}\Delta\boldsymbol{z}_{t}+[\Delta\boldsymbol{\beta}(y_{t})]^{\mathsf{T}}\boldsymbol{z}_{t}$
$=\boldsymbol{\beta}(y_{t-1})^{\mathsf{T}}\boldsymbol{c}+[I+\boldsymbol{\beta}(y_{t-1})^{\mathsf{T}}\boldsymbol{\alpha}]\boldsymbol{\xi}_{t-1}+\boldsymbol{\beta}(y_{t-1})^{\mathsf{T}}\boldsymbol{u}_{t}+[\Delta\boldsymbol{\beta}(y_{t})]_{1:1}^{\mathsf{T}}y_{t}.$ (B.19)
By construction, for $\delta_{t}$ as defined in (B.16), we have that $y_{t}=\delta_{t}\Delta y_{t}$ whenever $y_{t-1}$ and $y_{t}$ have opposite signs. Since this is also the only case in which $\Delta\boldsymbol{\beta}(y_{t})\neq 0$, it follows that
$[\Delta\boldsymbol{\beta}(y_{t})]_{1:1}^{\mathsf{T}}y_{t}=\delta_{t}[\Delta\boldsymbol{\beta}(y_{t})]_{1:1}^{\mathsf{T}}\Delta y_{t}$ $=\delta_{t}[\Delta\boldsymbol{\beta}(y_{t})]^{\mathsf{T}}\Delta\boldsymbol{z}_{t}$
$=\delta_{t}[\boldsymbol{\beta}(y_{t})-\boldsymbol{\beta}(y_{t-1})]^{\mathsf{T}}(\boldsymbol{c}+\boldsymbol{\alpha}\boldsymbol{\xi}_{t-1}+\boldsymbol{u}_{t}).$
Substituting this into (B.19) and recalling the definition of $\overline{\boldsymbol{\beta}}_{t}$ thus yields (B.17).
Finally, observe that $\delta_{t}$ is nonzero only if $y_{t-1}$ and $y_{t}$ have opposite signs: in which case $\lvert y_{t}\rvert\leq\lvert\Delta y_{t}\rvert$, and so $\lvert\delta_{t}\rvert\leq 1$. Since $y_{t}$ and $\Delta y_{t}$ must also have the same sign in this case, we deduce that $\delta_{t}\in[0,1]$, as claimed. ∎
Since $\overline{\boldsymbol{\beta}}_{t}$ is always a convex combination of $\boldsymbol{\beta}(-1)$ and $\boldsymbol{\beta}(+1)$, the autoregressive matrices in the representation (B.17) are contained in the set
$\mathcal{A}\coloneqq\{I+((1-\delta)\boldsymbol{\beta}(+1)+\delta\boldsymbol{\beta}(-1))^{\mathsf{T}}\boldsymbol{\alpha}\mid\delta\in[0,1]\}.$
By Jungers (2009, Prop. 1.8), $\rho_{{\scriptstyle\mathrm{JSR}}}(\mathcal{A})=\rho_{{\scriptstyle\mathrm{JSR}}}(\{I+\boldsymbol{\beta}(+1)^{\mathsf{T}}\boldsymbol{\alpha},I+\boldsymbol{\beta}(-1)^{\mathsf{T}}\boldsymbol{\alpha}\})$, which is strictly less than unity by CO(ii).2. That $\boldsymbol{\xi}_{t}\sim I^{\ast}(0)$ now follows from ERR, CO(ii).4 and Lemma A.1. In particular, $\xi_{t}=\beta(y_{t})^{\mathsf{T}}z_{t}$, $\Delta y_{t}^{+}$, $\Delta y_{t}^{-}$ and $\Delta x_{t}$ are all $I^{\ast}(0)$, and hence so too is
$\Delta y_{t}=(y_{t}-y_{t-1})=(y_{t}^{+}-y_{t-1}^{+})+(y_{t}^{-}-y_{t-1}^{-})=\Delta y_{t}^{+}+\Delta y_{t}^{-},$
as claimed in part (i) of the theorem.
For the purposes of proving part (ii) of the theorem, we shall initially suppose (as we did in the proof of Theorem 4.1) that DET is replaced by the weaker condition
$e_{1}^{\mathsf{T}}P_{\beta_{\perp}}(+1)c=0,$ (B.20)
for $P_{\beta_{\perp}}(+1)$ as defined in (4.7), noting only at the very end of the proof how the result simplifies when DET holds.
Our next step is to extract the common trend components. Recall the definitions of $\beta_{\perp}(y)$ and $\Gamma(1;y)$ in (4.8), and note that we may take
$\boldsymbol{\alpha}_{\perp}^{\mathsf{T}}$ $=\alpha_{\perp}^{\mathsf{T}}\begin{bmatrix}I_{p}&-\Gamma_{1}&\cdots&-\Gamma_{k-1}\end{bmatrix}$, $\boldsymbol{\beta}_{\perp}(y)^{\mathsf{T}}$ $\coloneqq\beta_{\perp}(y)^{\mathsf{T}}\begin{bmatrix}I_{p}&S(y)^{\mathsf{T}}&\cdots&S(y)^{\mathsf{T}}\end{bmatrix}$,
as the orthocomplements of $\boldsymbol{\alpha}$ and $\boldsymbol{\beta}(y)$ respectively. Observe that
$\boldsymbol{\alpha}_{\perp}^{\mathsf{T}}\boldsymbol{\beta}_{\perp}(y)=\alpha_{\perp}^{\mathsf{T}}\left[I_{p}-\sum_{i=1}^{k-1}\Gamma_{i}S(y)\right]\beta_{\perp}(y)=\alpha_{\perp}^{\mathsf{T}}\Gamma(1;y)\beta_{\perp}(y)$ (B.21)
where the r.h.s. is invertible for each $y\in\mathbb{R}$ by CO(ii).3. By Lemma A.1 in Hansen (2005), $\boldsymbol{\beta}(y)^{\mathsf{T}}\boldsymbol{\alpha}$ is therefore also invertible, whence it follows that the complementary projections
$\boldsymbol{P}_{\boldsymbol{\beta}_{\perp}}(y)$ $\coloneqq\boldsymbol{\beta}_{\perp}(y)[\boldsymbol{\alpha}_{\perp}^{\mathsf{T}}\boldsymbol{\beta}_{\perp}(y)]^{-1}\boldsymbol{\alpha}_{\perp}^{\mathsf{T}}$ $\boldsymbol{P}_{\boldsymbol{\alpha}}(y)$ $\coloneqq\boldsymbol{\alpha}[\boldsymbol{\beta}(y)^{\mathsf{T}}\boldsymbol{\alpha}]^{-1}\boldsymbol{\beta}(y)^{\mathsf{T}}$
are well defined for each $y\in\mathbb{R}$.
Now premultiplying (B.15) by $\boldsymbol{\alpha}_{\perp}^{\mathsf{T}}$ and cumulating yields
$\boldsymbol{\alpha}_{\perp}^{\mathsf{T}}\boldsymbol{z}_{t}=\boldsymbol{\alpha}_{\perp}^{\mathsf{T}}\boldsymbol{z}_{0}+\boldsymbol{\alpha}_{\perp}^{\mathsf{T}}\sum_{s=1}^{t}\boldsymbol{u}_{s}+(\boldsymbol{\alpha}_{\perp}^{\mathsf{T}}\boldsymbol{c})t.$
Hence
$\boldsymbol{z}_{t}=[\boldsymbol{P}_{\boldsymbol{\beta}_{\perp}}(y_{t})+\boldsymbol{P}_{\boldsymbol{\alpha}}(y_{t})]\boldsymbol{z}_{t}=\boldsymbol{P}_{\boldsymbol{\beta}_{\perp}}(y_{t})\left(\boldsymbol{z}_{0}+\sum_{s=1}^{t}\boldsymbol{u}_{s}+\boldsymbol{c}t\right)+\boldsymbol{\alpha}(\boldsymbol{\beta}(y_{t})^{\mathsf{T}}\boldsymbol{\alpha})^{-1}\boldsymbol{\xi}_{t}.$ (B.22)
In view of (B.21) and the definitions of $\boldsymbol{\alpha}_{\perp}$ and $\boldsymbol{\beta}_{\perp}(y)$, the upper left $p\times p$ block of $\boldsymbol{P}_{\boldsymbol{\beta}_{\perp}}(y)$ is $P_{\beta_{\perp}}(y)$, where the latter is as defined in (4.7). Under CO(ii).4, $n^{-1/2}z_{t}\overset{p}{\rightarrow}{\cal Z}_{0}$ for all $t\in\{-k+1,\ldots,0\}$, and so
$n^{-1/2}\boldsymbol{\alpha}_{\perp}^{\mathsf{T}}\boldsymbol{z}_{0}=n^{-1/2}\alpha_{\perp}^{\mathsf{T}}\left[z_{0}-\sum_{i=1}^{k-1}\Gamma_{i}z_{-i}^{\ast}\right]\overset{p}{\rightarrow}\alpha_{\perp}^{\mathsf{T}}\Gamma(1;{\cal Y}_{0}){\cal Z}_{0}.$
Hence, using the result of part (i), the first $p$ rows of (B.22) give
$\begin{bmatrix}y_{t}\\ x_{t}\end{bmatrix}=P_{\beta_{\perp}}(y_{t})\left[n^{1/2}\Gamma(1;{\cal Y}_{0}){\cal Z}_{0}+\sum_{s=1}^{t}u_{s}+ct\right]+o_{p}(n^{1/2})$ (B.23)
uniformly in $t\in\{1,\ldots,n\}$. To determine the weak limit of the preceding upon standardisation (and possibly linear detrending), we first provide an auxiliary result on the mapping
$g(y,u)\coloneqq P_{\beta_{\perp}}(y)u.$ (B.24)
Lemma B.3.
Suppose CVAR and CO(ii) hold. Then
  1. (i)
    taking $\vartheta^{\mathsf{T}}\coloneqq e_{1}^{\mathsf{T}}P_{\beta_{\perp}}(+1)\neq 0$, there exists a $\mu>0$ such that
    $e_{1}^{\mathsf{T}}P_{\beta_{\perp}}(y)=[\mathbf{1}^{+}(y)+\mu\mathbf{1}^{-}(y)]\vartheta^{\mathsf{T}}\eqqcolon h(y)\vartheta^{\mathsf{T}};$
  2. (ii)
    if $u\in\mathbb{R}^{p}$ is such that $\vartheta^{\mathsf{T}}u=0$, then $P_{\beta_{\perp}}(+1)u=P_{\beta_{\perp}}(-1)u$; and
  3. (iii)
    $g$ is Lipschitz continuous at every point in $D_{g}\coloneqq\{(y,u)\in\mathbb{R}^{p+1}\mid y=h(y)\vartheta^{\mathsf{T}}u\}$, in the sense that for every $(y,u)\in D_{g}$ and $(y^{\prime},u^{\prime})\in\mathbb{R}\times\mathbb{R}^{p}$,
    $\lvert g(y,u)-g(y^{\prime},u^{\prime})\rvert\leq C(\lvert y-y^{\prime}\rvert+\lVert u-u^{\prime}\rVert).$ (B.25)
Proof.
(i). Recall from (4.6) that $\beta^{\pm\mathsf{T}}=[\beta_{y}^{\pm},\beta_{x}^{\mathsf{T}}]=\beta_{x}^{\mathsf{T}}[\theta^{\pm},I_{p-1}]$ for some $\theta^{\pm}\in\mathbb{R}^{p-1}$, and take
$\beta_{\perp}^{\pm}=\begin{bmatrix}1&0\\ -\theta^{\pm}&\beta_{x,\perp}\end{bmatrix}=\beta_{\perp}(\pm 1)$ (B.26)
as per (4.8), so that $e_{1}^{\mathsf{T}}\beta_{\perp}^{\pm}=e_{1}^{\mathsf{T}}$; it follows that $e_{1}^{\mathsf{T}}P_{\beta_{\perp}}(+1)\neq 0$. Since
$\Gamma^{\pm}(1)\beta_{\perp}^{\pm}=\begin{bmatrix}\gamma^{\pm}(1)&\Gamma^{x}(1)\end{bmatrix}\begin{bmatrix}1&0\\ -\theta^{\pm}&\beta_{x,\perp}\end{bmatrix}=\begin{bmatrix}\gamma^{\pm}(1)-\Gamma^{x}(1)\theta^{\pm}&\Gamma^{x}(1)\beta_{x,\perp}\end{bmatrix}$
we have
$\Gamma^{-}(1)\beta_{\perp}^{-}-\Gamma^{+}(1)\beta_{\perp}^{+}=\delta_{\gamma}e_{1}^{\mathsf{T}},$
where $\delta_{\gamma}\coloneqq[\gamma^{-}(1)-\gamma^{+}(1)-\Gamma^{x}(1)(\theta^{-}-\theta^{+})]$. Under CO(ii).3, we may apply Lemma A.2 with $A=\alpha_{\perp}^{\mathsf{T}}$, $B_{1}=\Gamma^{-}(1)\beta_{\perp}^{-}$, $B_{2}=\Gamma^{+}(1)\beta_{\perp}^{+}$ and $d=e_{1}$, to obtain
$e_{1}^{\mathsf{T}}[\alpha_{\perp}^{\mathsf{T}}\Gamma^{-}(1)\beta_{\perp}^{-}]^{-1}=\mu e_{1}^{\mathsf{T}}[\alpha_{\perp}^{\mathsf{T}}\Gamma^{+}(1)\beta_{\perp}^{+}]^{-1}$
for some $\mu>0$. Hence, for $\vartheta^{\mathsf{T}}=e_{1}^{\mathsf{T}}P_{\beta_{\perp}}(+1)=e_{1}^{\mathsf{T}}[\alpha_{\perp}^{\mathsf{T}}\Gamma^{+}(1)\beta_{\perp}^{+}]^{-1}\alpha_{\perp}^{\mathsf{T}}$
$e_{1}^{\mathsf{T}}P_{\beta_{\perp}}(y)$ $=e_{1}^{\mathsf{T}}\beta_{\perp}(y)[\alpha_{\perp}^{\mathsf{T}}\Gamma(1;y)\beta_{\perp}(y)]^{-1}\alpha_{\perp}^{\mathsf{T}}=e_{1}^{\mathsf{T}}[\alpha_{\perp}^{\mathsf{T}}\Gamma(1;y)\beta_{\perp}(y)]^{-1}\alpha_{\perp}^{\mathsf{T}}$
$=[\mathbf{1}^{+}(y)+\mu\mathbf{1}^{-}(y)]\vartheta^{\mathsf{T}}=h(y)\vartheta^{\mathsf{T}}$
(ii). Let $u\in\mathbb{R}^{p}$ be such that $\vartheta^{\mathsf{T}}u=0$. Then
$0=\mu\vartheta^{\mathsf{T}}u=e_{1}^{\mathsf{T}}[\alpha_{\perp}^{\mathsf{T}}\Gamma^{-}(1)\beta_{\perp}^{-}]^{-1}\alpha_{\perp}^{\mathsf{T}}u=d^{\mathsf{T}}(A^{\mathsf{T}}B_{1})^{-1}v$
where $v\coloneqq\alpha_{\perp}^{\mathsf{T}}u$, and $A$, $B_{1}$, $B_{2}$ and $d$ are as above. Hence by Lemma A.2,
$0$ $=[(A^{\mathsf{T}}B_{2})^{-1}-(A^{\mathsf{T}}B_{1})^{-1}]v=[(\alpha_{\perp}^{\mathsf{T}}\Gamma^{+}(1)\beta_{\perp}^{+})^{-1}-(\alpha_{\perp}^{\mathsf{T}}\Gamma^{-}(1)\beta_{\perp}^{-})^{-1}]\alpha_{\perp}^{\mathsf{T}}u.$
Finally, noting that (B.26) implies $\beta_{\perp}^{+}-\beta_{\perp}^{-}=\delta_{\beta}e_{1}^{\mathsf{T}}$ for $\delta_{\beta}\coloneqq(0,(\theta^{-}-\theta^{+})^{\mathsf{T}})^{\mathsf{T}}$, we have
$[P_{\beta_{\perp}}(+1)-P_{\beta_{\perp}}(-1)]u$ $=\beta_{\perp}^{+}[(\alpha_{\perp}^{\mathsf{T}}\Gamma^{+}(1)\beta_{\perp}^{+})^{-1}-(\alpha_{\perp}^{\mathsf{T}}\Gamma^{-}(1)\beta_{\perp}^{-})^{-1}]\alpha_{\perp}u$
$\qquad\qquad-(\beta_{\perp}^{-}-\beta_{\perp}^{+})(\alpha_{\perp}^{\mathsf{T}}\Gamma^{-}(1)\beta_{\perp}^{-})^{-1}\alpha_{\perp}u$
$=\delta_{\beta}e_{1}^{\mathsf{T}}(\alpha_{\perp}^{\mathsf{T}}\Gamma^{-}(1)\beta_{\perp}^{-})^{-1}\alpha_{\perp}^{\mathsf{T}}u$
$=_{(1)}\delta_{\beta}e_{1}^{\mathsf{T}}P_{\beta_{\perp}}(-1)u=_{(2)}\mu\delta_{\beta}\vartheta^{\mathsf{T}}u=_{(3)}0,$
where $_{(1)}$ holds by $e_{1}^{\mathsf{T}}\beta_{\perp}^{-}=e_{1}^{\mathsf{T}}$, $_{(2)}$ by the result of part (i), and $_{(3)}$ by our choice of $u$.
(iii). Now let $(y,u)\in D_{g}$ and $(y^{\prime},u^{\prime})\in\mathbb{R}\times\mathbb{R}^{p}$. Suppose first that $h(y)=h(y^{\prime})$: then
$\lVert g(y,u)-g(y^{\prime},u^{\prime})\rVert\leq\max\{\lVert P_{\beta_{\perp}}(+1)\rVert,\lVert P_{\beta_{\perp}}(-1)\rVert\}\lVert u-u^{\prime}\rVert\leq C\lVert u-u^{\prime}\rVert.$
Suppose next that $h(y)\neq h(y^{\prime})$; then either: (a) $y\geq 0$ and $y^{\prime}<0$; or (b) $y<0$ and $y^{\prime}\geq 0$. Suppose (a) holds. Then $\vartheta^{\mathsf{T}}u\geq 0$, and so
$\lvert y-y^{\prime}\rvert=y-y^{\prime}=h(y)\vartheta^{\mathsf{T}}u-y^{\prime}\geq h(y)\vartheta^{\mathsf{T}}u\geq 0$
whence
$\lvert\vartheta^{\mathsf{T}}u\rvert\leq C_{0}\lvert y-y^{\prime}\rvert$ (B.27)
where $C_{0}\coloneqq\max\{1,\mu^{-1}\}\in(0,\infty)$. (If (b) holds instead, then (B.27) follows by an analogous argument.) Now let $u_{\vartheta_{\perp}}\coloneqq u-\vartheta(\vartheta^{\mathsf{T}}\vartheta)^{-1}\vartheta^{\mathsf{T}}u$. Then $\vartheta^{\mathsf{T}}u_{\vartheta_{\perp}}=0$, and so
$g(y,u)-g(y^{\prime},u^{\prime})$ $=P_{\beta_{\perp}}(+1)(u-u_{\vartheta_{\perp}}+u_{\vartheta_{\perp}})-P_{\beta_{\perp}}(-1)u^{\prime}$
$=P_{\beta_{\perp}}(+1)(u-u_{\vartheta_{\perp}})+P_{\beta_{\perp}}(-1)(u_{\vartheta_{\perp}}-u^{\prime}),$ (B.28)
where the second equality follows by the result of part (ii). By (B.27),
$\lVert u-u_{\vartheta_{\perp}}\rVert\leq\lVert\vartheta(\vartheta^{\mathsf{T}}\vartheta)^{-1}\rVert\lvert\vartheta^{\mathsf{T}}u\rvert\leq C_{1}\lvert y-y^{\prime}\rvert$ (B.29)
where $C_{1}\coloneqq\lVert\vartheta(\vartheta^{\mathsf{T}}\vartheta)^{-1}\rVert C_{0}$. Further,
$\lVert u_{\vartheta_{\perp}}-u^{\prime}\rVert\leq\lVert u_{\vartheta_{\perp}}-u\rVert+\lVert u^{\prime}-u\rVert\leq C_{1}\lvert y-y^{\prime}\rvert+\lVert u^{\prime}-u\rVert.$ (B.30)
Thus (B.25) follows from (B.28)–(B.30) ∎
Returning now to (B.23), we note that (B.20) and Lemma B.3 imply that $P_{\beta_{\perp}}(y)c=P_{\beta_{\perp}}(+1)c$ for all $y\in\mathbb{R}$. Therefore by defining the detrended process
$x_{t}^{d}\coloneqq x_{t}-[E_{-1}^{\mathsf{T}}P_{\beta_{\perp}}(+1)c]t,$
for $E_{-1}$ the final p1p-1 columns of $I_{p}$, we may rewrite (B.23) as
$n^{-1/2}\begin{bmatrix}y_{t}\\ x_{t}^{d}\end{bmatrix}=P_{\beta_{\perp}}(y_{t})\left[\Gamma(1;{\cal Y}_{0}){\cal Z}_{0}+n^{-1/2}\sum_{s=1}^{t}u_{s}\right]+o_{p}(1)$
uniformly in $t\in\{1,\ldots,n\}$, and so for $\lambda\in[0,1]$,
$\begin{bmatrix}Y_{n}(\lambda)\\ X_{n}^{d}(\lambda)\end{bmatrix}=P_{\beta_{\perp}}[Y_{n}(\lambda)]U_{n,0}(\lambda)+\overline{o}_{p}(1),$ (B.31)
where $X_{n}^{d}(\lambda)\coloneqq n^{-1/2}x_{\lfloor n\lambda\rfloor}^{d}$, $U_{n,0}(\lambda)\coloneqq\Gamma(1;{\cal Y}_{0}){\cal Z}_{0}+U_{n}(\lambda)$, $\overline{o}_{p}(1)$ denotes a term that converges in probability to zero uniformly over $\lambda\in[0,1]$, and we have used the fact that $P_{\beta_{\perp}}(y)$ depends only on the sign of $y$.
It remains to determine the weak limit of the r.h.s. of (B.31). By Lemma B.3, $e_{1}^{\mathsf{T}}P_{\beta_{\perp}}(y)=h(y)\vartheta^{\mathsf{T}}$, and the first equation of (B.31) becomes
$Y_{n}(\lambda)=h[Y_{n}(\lambda)]\vartheta^{\mathsf{T}}U_{n,0}(\lambda)+\overline{o}_{p}(1).$
Now let $f(y)\coloneqq h(y)^{-1}y$. This is Lipschitz, and because $h(y)>0$ is bounded away from zero and infinity, it has an inverse $f^{-1}(y^{\ast})=h(y^{\ast})y^{\ast}$ that is also Lipschitz. Thus
$Y_{n}^{\ast}(\lambda)\coloneqq f[Y_{n}(\lambda)]=\vartheta^{\mathsf{T}}U_{n,0}(\lambda)+h[Y_{n}(\lambda)]^{-1}\overline{o}_{p}(1)\rightsquigarrow\vartheta^{\mathsf{T}}U_{0}(\lambda)$
on $D[0,1]$, since $\sup_{\lambda\in[0,1]}h[Y_{n}(\lambda)]^{-1}\leq\max\{1,\mu^{-1}\}$. Hence, by the CMT,
$Y_{n}(\lambda)=f^{-1}[Y_{n}^{\ast}(\lambda)]\rightsquigarrow f^{-1}[\vartheta^{\mathsf{T}}U_{0}(\lambda)]\\ =h[\vartheta^{\mathsf{T}}U_{0}(\lambda)]\vartheta^{\mathsf{T}}U_{0}(\lambda)=e_{1}^{\mathsf{T}}P_{\beta_{\perp}}[\vartheta^{\mathsf{T}}U_{0}(\lambda)]U_{0}(\lambda)=Y(\lambda)$ (B.32)
on $D[0,1]$. Now consider the remaining p1p-1 equations in (B.31). Recalling the definition of $g$ in (B.24) above, we have
$X_{n}^{d}(\lambda)=E_{-1}g[Y_{n}(\lambda),U_{n,0}(\lambda)]+\overline{o}_{p}(1).$
By (B.32), $(Y_{n},U_{n,0})\rightsquigarrow(Y,U_{0})$, which have continuous paths and concentrate in the set
$\overline{D}\coloneqq\{(y,u)\in D_{\mathbb{R}^{p+1}}[0,1]\mid[y(\lambda),u(\lambda)]\in D_{g}\text{ for all }\lambda\in[0,1]\}.$
By Lemma B.3, if $(y_{n},u_{n})\in D_{\mathbb{R}^{p+1}}[0,1]$ converge uniformly to $(y,u)\in\overline{D}$, then $g[y_{n}(\lambda),u_{n}(\lambda)]\rightarrow g[y(\lambda),u(\lambda)]$ uniformly over $\lambda\in[0,1]$. Hence by the CMT,
$X_{n}^{d}(\lambda)\rightsquigarrow E_{-1}g[Y(\lambda),U_{0}(\lambda)]$ $=E_{-1}P_{\beta_{\perp}}[Y(\lambda)]U_{0}(\lambda)=E_{-1}P_{\beta_{\perp}}[\vartheta^{\mathsf{T}}U_{0}(\lambda)]U_{0}(\lambda)$ (B.33)
on $D_{\mathbb{R}^{p-1}}[0,1]$.
Finally, we suppose that DET holds in place of (B.20). In this case, $P_{\beta_{\perp}}(+1)c=0$ and therefore $x_{t}^{d}=x_{t}$ for all $t\in\{1,\ldots,T\}$. Thus part (ii) of the theorem follows from (B.32) and (B.33). ∎
B.3 Proof of Theorem 4.3
We first restate a simplified version of Lemma B.2 from Duffy et al. (2023).
Lemma B.4.
Let $\{w_{t}\}_{t\in\mathbb{N}}$ be a random sequence in $\mathbb{R}^{p}$, and $\boldsymbol{w}_{t-1}\coloneqq(w_{t-1}^{\mathsf{T}},\ldots,w_{t-k}^{\mathsf{T}})^{\mathsf{T}}$. Suppose
$w_{t}=\mu(\boldsymbol{w}_{t-1})+\boldsymbol{A}(\boldsymbol{w}_{t-1})\boldsymbol{w}_{t-1}+u_{t}$ (B.34)
where $\boldsymbol{A}:\mathbb{R}^{kp}\rightarrow\mathbb{R}^{p}$ is such that $\boldsymbol{w}\mapsto\boldsymbol{A}(\boldsymbol{w})\boldsymbol{w}$ is continuous and homogeneous of degree one, $\mu(\boldsymbol{w})=o(\lVert\boldsymbol{w}\rVert)$ as $\lVert\boldsymbol{w}\rVert\rightarrow\infty$, and that $\{u_{t}\}$ satisfies ERR. If the associated deterministic system
$\hat{w}_{t}=\boldsymbol{A}(\hat{\boldsymbol{w}}_{t-1})\hat{\boldsymbol{w}}_{t-1},$
is stable (in the sense that $\hat{w}_{t}\rightarrow 0$ for every initialisation $\hat{\boldsymbol{w}}_{0}\in\mathbb{R}^{kp}$), then $\{w_{t}\}_{t\in\mathbb{N}}$ is $\mathcal{Q}$-geometrically ergodic, for $\mathcal{Q}(w)\coloneqq(1+\lVert w\rVert^{m_{0}})$, with a stationary distribution that is absolutely continuous with respect to Lebesgue measure, and which has finite $m_{0}$th moment.
To prove Theorem 4.3, we need to put the model into a VECM form different again from that used in the analysis of the preceding cases. Recalling $z_{t}^{\mathsf{T}}=(y_{t},x_{t}^{\mathsf{T}})$, the canonical CKSVAR (2.17) may be written as
$z_{t}=c+\sum_{i=1}^{k}\Phi_{i}(y_{t-i})z_{t-i},$
where $\Phi_{i}(y)\coloneqq[\phi_{i}(y),\Phi^{x}]$ for $\phi_{i}(y)\coloneqq\phi_{i}^{+}\mathbf{1}^{+}(y)+\phi_{i}^{-}\mathbf{1}^{-}(y)$, and hence
$\Delta z_{t}=c+\Pi(\boldsymbol{y}_{t-1})z_{t-1}+\sum_{i=1}^{k-1}\Gamma_{i}(\boldsymbol{y}_{t-1})\Delta z_{t-i}+u_{t}.$
where $\Pi(\boldsymbol{y}_{t-i})\coloneqq\sum_{i=1}^{k}\Phi_{i}(y_{t-i})-I_{p}$, and as per (4.17) above,
$\Gamma_{i}(\boldsymbol{y}_{t-1})=-\sum_{j=i+1}^{k}\Phi_{j}(y_{t-j})=\begin{bmatrix}-\sum_{j=i+1}^{k}\phi_{j}(y_{t-j}),&-\sum_{j=i+1}^{k}\Phi_{j}^{x}\end{bmatrix}$.
Taking $\boldsymbol{z}_{t}\coloneqq(z_{t}^{\mathsf{T}},\ldots,z_{t-k+1}^{\mathsf{T}})^{\mathsf{T}}$, we may write the system in companion form as
$\Delta\boldsymbol{z}_{t}=\boldsymbol{c}+\boldsymbol{\Pi}(\boldsymbol{y}_{t-1})\boldsymbol{z}_{t-1}+\boldsymbol{u}_{t}$ (B.35)
where for $\boldsymbol{\alpha}(\boldsymbol{y}_{t-1})$ and $\boldsymbol{\beta}$ as defined in (4.16),
$\boldsymbol{\Pi}(\boldsymbol{y})$ $=\begin{bmatrix}\Pi(\boldsymbol{y})+\Gamma_{1}(\boldsymbol{y})&-\Gamma_{1}(\boldsymbol{y})+\Gamma_{2}(\boldsymbol{y})&\cdots&\Gamma_{k-1}(\boldsymbol{y})\\ I_{p}&-I_{p}\\ &\ddots&\ddots\\ &&I_{p}&-I_{p}\end{bmatrix}$ $\boldsymbol{c}$ $\coloneqq\begin{bmatrix}c\\ 0_{p}\\ \vdots\\ 0_{p}\end{bmatrix}$ $\boldsymbol{u}_{t}$ $\coloneqq\begin{bmatrix}u_{t}\\ 0_{p}\\ \vdots\\ 0_{p}\end{bmatrix}$
$=\boldsymbol{\alpha}(\boldsymbol{y})\boldsymbol{\beta}^{\mathsf{T}}$
Hence for $\boldsymbol{\xi}_{t}\coloneqq\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{z}_{t}=(\xi_{t}^{\mathsf{T}},\Delta z_{t}^{\mathsf{T}},\ldots,\Delta z_{t-k+2}^{\mathsf{T}})^{\mathsf{T}}$, where $\xi_{t}=\beta^{\mathsf{T}}z_{t}$, it follows from (B.35) that
$\boldsymbol{\xi}_{t}=\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{c}+(I_{p(k-1)+r}+\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{\alpha}(\boldsymbol{y}_{t-1}))\boldsymbol{\xi}_{t-1}+\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{u}_{t}.$ (B.36)
To ‘close’ (B.36), we need only to recognise that with the ordering of the cointegrating vectors given in (4.15), $y_{t-1}=e_{r,1}^{\mathsf{T}}\xi_{t-1}$ and
$y_{t-i}=y_{t-1}-\sum_{j=1}^{i-1}\Delta y_{t-j}=e_{r,1}^{\mathsf{T}}\xi_{t-1}-\sum_{j=1}^{i-1}e_{p,1}^{\mathsf{T}}\Delta z_{t-j}$
so that $\boldsymbol{y}_{t-1}$ can be extracted from $\boldsymbol{\xi}_{t-1}$ via
$\boldsymbol{y}_{t-1}=\begin{bmatrix}y_{t-1}\\ y_{t-2}\\ \vdots\\ y_{t-k}\end{bmatrix}=\begin{bmatrix}e_{r,1}^{\mathsf{T}}\\ e_{r,1}^{\mathsf{T}}&-e_{p,1}^{\mathsf{T}}\\ \vdots&\vdots&\ddots\\ e_{r,1}^{\mathsf{T}}&-e_{p,1}^{\mathsf{T}}&\cdots&-e_{p,1}^{\mathsf{T}}\end{bmatrix}\begin{bmatrix}\xi_{t-1}\\ \Delta z_{t-1}\\ \vdots\\ \Delta z_{t-k+1}\end{bmatrix}\eqqcolon\boldsymbol{G}_{\xi}\boldsymbol{\xi}_{t-1}.$ (B.37)
Hence
$\boldsymbol{\xi}_{t}$ $=\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{c}+[I_{p(k-1)+r}+\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{\alpha}(\boldsymbol{G}_{\xi}\boldsymbol{\xi}_{t-1})]\boldsymbol{\xi}_{t-1}+\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{u}_{t}.$ (B.38)
We would like to apply Lemma B.4 to this system, but we are prevented from doing so because $\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{u}_{t}$ does not satisfy the required condition on the innovations. To remedy this, we modify the state vector of the system as follows. Let $\overline{\beta}\coloneqq\beta(\beta^{\mathsf{T}}\beta)^{-1}$, $\overline{\beta}_{\perp}\coloneqq\beta_{\perp}(\beta_{\perp}^{\mathsf{T}}\beta_{\perp})^{-1}$, and recall $\chi_{t}=(\xi_{t}^{\mathsf{T}},(\beta_{\perp}^{\mathsf{T}}\Delta z_{t})^{\mathsf{T}})^{\mathsf{T}}$ and $\boldsymbol{\chi}_{t}=(\chi_{t}^{\mathsf{T}},\ldots,\chi_{t-k+1}^{\mathsf{T}})^{\mathsf{T}}$. Since the r.h.s. of
$\Delta z_{t}=(\overline{\beta}_{\perp}\beta_{\perp}^{\mathsf{T}}+\overline{\beta}\beta^{\mathsf{T}})\Delta z_{t}=\overline{\beta}_{\perp}(\beta_{\perp}^{\mathsf{T}}\Delta z_{t})+\overline{\beta}(\xi_{t}-\xi_{t-1})$
is a linear function of $\chi_{t}$ and $\chi_{t-1}$, there is a matrix $\boldsymbol{H}$ such that $\boldsymbol{\xi}_{t}=\boldsymbol{H}\boldsymbol{\chi}_{t}$. Defining
$M$ $\coloneqq\begin{bmatrix}I_{r}&0&0_{r\times p(k-2)}\\ 0&\beta_{\perp}^{\mathsf{T}}&0_{q\times p(k-2)}\end{bmatrix}$ $B$ $\coloneqq\begin{bmatrix}\beta^{\mathsf{T}}\\ \beta_{\perp}^{\mathsf{T}}\end{bmatrix}$
we have $M\boldsymbol{\xi}_{t}=\chi_{t}$, $M\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{c}=Bc$ and $M\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{u}_{t}=Bu_{t}$, and so premultiplying (B.38) by $M$ yields
$\chi_{t}=Bc+M[I_{p(k-1)+r}+\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{\alpha}(\boldsymbol{G}\boldsymbol{\chi}_{t-1})]\boldsymbol{H}\boldsymbol{\chi}_{t-1}+Bu_{t}$ (B.39)
where $\boldsymbol{G}\coloneqq\boldsymbol{G}_{\xi}\boldsymbol{H}$.
We may now apply Lemma B.4 to (B.39). Since $B$ has full rank, ERR is satisfied with $Bu_{t}$ taking the place of $u_{t}$. Regarding the other conditions of the lemma, let $\boldsymbol{F}$ be such that $\boldsymbol{F}\boldsymbol{z}_{t-1}=\boldsymbol{y}_{t-1}$. By the continuity of the r.h.s. of the CKSVAR, the map
$\boldsymbol{z}_{t-1}\mapsto[I_{kp}+\boldsymbol{\Pi}(\boldsymbol{F}\boldsymbol{z}_{t-1})]\boldsymbol{z}_{t-1}$
which simply replicates the autoregressive part of the companion form (B.35) of the CKSVAR (in levels), is continuous. Hence so too is
$\boldsymbol{z}_{t-1}\mapsto[I_{p(k-1)+r}+\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{\alpha}(\boldsymbol{F}\boldsymbol{z}_{t-1})]\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{z}_{t-1}.$
$\boldsymbol{\xi}_{t-1}=\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{z}_{t-1}$ varies freely, and by our arguments following (B.38), $\boldsymbol{F}\boldsymbol{z}_{t-1}$ depends (continuously) only on the elements of $\boldsymbol{\xi}_{t-1}$. Indeed, $\boldsymbol{F}\boldsymbol{z}_{t-1}=\boldsymbol{y}_{t-1}=\boldsymbol{G}_{\xi}\boldsymbol{\xi}_{t-1}$ where $\boldsymbol{G}_{\xi}$ is as defined in (B.37). Hence the mapping
$\boldsymbol{\xi}_{t-1}\mapsto[I_{p(k-1)+r}+\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{\alpha}(\boldsymbol{G}_{\xi}\boldsymbol{\xi}_{t-1})]\boldsymbol{\xi}_{t-1},$
that appears on the r.h.s. of (B.38), is also continuous. Thus
$\boldsymbol{\chi}_{t-1}\mapsto M[I_{p(k-1)+r}+\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{\alpha}(\boldsymbol{G}\boldsymbol{\chi}_{t-1})]\boldsymbol{H}\boldsymbol{\chi}_{t-1}$
is continuous by the composition of continuous maps. Since $\boldsymbol{\alpha}(\boldsymbol{y})$ depends only on the signs of the elements of $\boldsymbol{y}$, the preceding is also homogeneous of degree one. Finally, since it is possible to write $\boldsymbol{\chi}_{t+1}$ as a (deterministic) function of $\xi_{t+1}$ and $\boldsymbol{\xi}_{t}$, the deterministic system associated to (B.39) must be stable if that associated to (B.38) is stable, where the latter is maintained under CO(iii).2. The result now follows by Lemma B.4. ∎
B.4 Proof of Theorem 4.4
Suppose initially that DGP holds. With the exception of the final claim made in each, the proofs of Theorems 4.1 and 4.2 proceed under conditions (B.8) and (B.20) respectively, which correspond to DET rather than DET, and yield the claimed limit weak limits for $X_{n}^{d}$. Thus for $z_{t}=(y_{t},x_{t}^{\mathsf{T}})^{\mathsf{T}}$ generated by a canonical CKSVAR, the theorem is proved.
Now suppose that merely DGP holds (along with CO(i) or CO(ii)). It is shown in Appendix D that, in these cases, the derived canonical form (denoted by tildes) satisfies $e_{1}^{\mathsf{T}}\tilde{P}_{\beta_{\perp}^{+}}\tilde{c}=0$ in case (i), and $e_{1}^{\mathsf{T}}\tilde{P}_{\beta_{\perp}}(+1)\tilde{c}=0$ in case (ii), i.e. DET holds in the canonical form; the argument subsequently given in that appendix shows that the conclusions of the theorem may then be transposed back to the structural form. ∎
Appendix C Proofs of auxiliary lemmas
Proof of Lemma A.1.
Since $\rho_{{\scriptstyle\mathrm{JSR}}}(\mathcal{A})<1$, it follows by Jungers (2009, Prop. 1.4) that for any given $\gamma\in(\rho_{{\scriptstyle\mathrm{JSR}}}(\mathcal{A}),1)$, there exists a norm $\lVert\cdot\rVert_{\ast}$ on $\mathbb{R}^{d_{w}}$ such that
$\lVert w_{t}\rVert_{\ast}\leq\lVert c_{t}\rVert_{\ast}+\lVert A_{t}\rVert_{\ast}\lVert w_{t-1}\rVert_{\ast}+\lVert B_{t}\rVert_{\ast}\lVert\varepsilon_{t}\rVert\leq\gamma\lVert w_{t-1}\rVert_{\ast}+C(1+\lVert v_{t}\rVert_{\ast})$
where $C\coloneqq\max\{\sup_{B\in\mathcal{B}}\lVert B\rVert_{\ast},\sup_{c\in{\cal C}}\lVert c\rVert_{\ast}\}<\infty$, by the boundedness of $\mathcal{B}$ and $\cal C$}. Hence by backward substitution,
$\lVert w_{t}\rVert_{\ast}\leq C\sum_{s=0}^{t-1}\gamma^{s}(1+\lVert v_{t-s}\rVert_{\ast})+\gamma^{t}\lVert w_{0}\rVert_{\ast}.$
By the equivalence of norms on finite-dimensional spaces, there exists a $C^{\prime}<\infty$ such that
$\lVert w_{t}\rVert\leq C^{\prime}\left[\sum_{s=0}^{t-1}\gamma^{s}(1+\lVert v_{t-s}\rVert)+\gamma^{t}\lVert w_{0}\rVert\right],$
where $\lVert\cdot\rVert$ denotes the Euclidean norm on $\mathbb{R}^{d_{w}}$. Deduce that for any $m\in[1,m_{0}]$,
$\lVert w_{t}\rVert_{m}$ $\leq C^{\prime}\left[\sum_{s=0}^{t-1}\gamma^{s}(1+\lVert v_{t-s}\rVert_{m})+\gamma^{t}\lVert w_{0}\rVert_{m}\right]$
$\leq C^{\prime}\left[\frac{1}{1-\gamma}\left(1+\max_{1\leq s\leq t}\lVert v_{s}\rVert_{m}\right)+\lVert w_{0}\rVert_{m}\right],$
where $\lVert w_{t}\rVert_{m}=\mathbb{E}(\lVert w_{t}\rVert^{m})^{1/m}$, etc., and hence there exists a $C^{\prime\prime}<\infty$ such that
$\max_{1\leq t\leq n}\lVert w_{t}\rVert_{m}\leq C^{\prime\prime}\left[1+\max_{1\leq t\leq n}\lVert v_{s}\rVert_{m}+\lVert w_{0}\rVert_{m}\right].\qed$
Proof of Lemma A.2.
Since $A^{\mathsf{T}}B_{1}=A^{\mathsf{T}}B_{2}+(A^{\mathsf{T}}c)d^{\mathsf{T}}$, it follows by Cauchy’s formula for a rank-one perturbation (Horn and Johnson, 2013, (0.8.5.11)) that
$\det(A^{\mathsf{T}}B_{1})$ $=\det(A^{\mathsf{T}}B_{2})+d^{\mathsf{T}}(\operatorname{adj}A^{\mathsf{T}}B_{2})A^{\mathsf{T}}c$
$=\det(A^{\mathsf{T}}B_{2})\{1+d^{\mathsf{T}}(A^{\mathsf{T}}B_{2})^{-1}A^{\mathsf{T}}c\}$
whence $\chi\coloneqq 1+d^{\mathsf{T}}(A^{\mathsf{T}}B_{2})^{-1}A^{\mathsf{T}}c=\det(A^{\mathsf{T}}B_{2})^{-1}\det(A^{\mathsf{T}}B_{1})>0$. Hence, by the Sherman–Morrison–Woodbury formula (Horn and Johnson, 2013, (0.7.4.2))
$(A^{\mathsf{T}}B_{1})^{-1}=(A^{\mathsf{T}}B_{2})^{-1}-\frac{(A^{\mathsf{T}}B_{2})^{-1}A^{\mathsf{T}}cd^{\mathsf{T}}(A^{\mathsf{T}}B_{2})^{-1}}{1+d^{\mathsf{T}}(A^{\mathsf{T}}B_{2})^{-1}A^{\mathsf{T}}c}.$ (C.1)
Premultiplying the preceding by $d^{\mathsf{T}}$ and rearranging yields
$d^{\mathsf{T}}(A^{\mathsf{T}}B_{1})^{-1}$ $=\left\{1-\frac{d^{\mathsf{T}}(A^{\mathsf{T}}B_{2})^{-1}A^{\mathsf{T}}c}{1+d^{\mathsf{T}}(A^{\mathsf{T}}B_{2})^{-1}A^{\mathsf{T}}c}\right\}d^{\mathsf{T}}(A^{\mathsf{T}}B_{2})^{-1}=\chi^{-1}d^{\mathsf{T}}(A^{\mathsf{T}}B_{2})^{-1}$
and thus part (i) holds with $\mu=\chi^{-1}$. Finally, let $v\in\mathbb{R}^{n}$ be such that $d^{\mathsf{T}}(A^{\mathsf{T}}B_{1})^{-1}v=0$. By part (i) of the lemma, we must have that $d^{\mathsf{T}}(A^{\mathsf{T}}B_{2})^{-1}v=0$; hence postmultiplying (C.1) by $v$ immediately yields the result of part (ii). ∎
Proof of Lemma A.3.
The factorisation of $\Pi$ is immediate from the definition of matrix rank; that of $\boldsymbol{\Pi}$ then follows from direct calculation. Part (iii) will also follow from direct calculation, once we have verified (ii).
It remains to prove part (ii), which is the converse to Lemma A.2 in Hansen (2005). We first observe that $I_{kp}+\boldsymbol{\Pi}$ gives the companion form matrix associated with the autoregressive polynomial $\Phi(\lambda)$, in the sense that the roots of the latter coincide with the reciprocals of the eigenvalues of the former (e.g. Lütkepohl, 2007, Sec. 2.1.1), of which $q$ lie on the unit circle by assumption, and the remaining kpqkp-q lie strictly inside. It is readily verified that $\boldsymbol{\alpha}_{\perp}^{\mathsf{T}}\boldsymbol{\alpha}=0$ and $\boldsymbol{\beta}_{\perp}^{\mathsf{T}}\boldsymbol{\beta}=0$, when $\boldsymbol{\alpha}_{\perp}$ and $\boldsymbol{\beta}_{\perp}$ have the form given in (A.1). Moreover, since $\beta_{\perp}$ is determined only up to its column span, we are free to normalise it such that $\beta_{\perp}^{\mathsf{T}}\beta_{\perp}=k^{-1}I_{q}$, whence $\boldsymbol{\beta}_{\perp}^{\mathsf{T}}\boldsymbol{\beta}_{\perp}=I_{q}$. Letting $Q\coloneqq(\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{\beta})^{-1/2}$ denote the positive definite square root of $(\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{\beta})^{-1}$, we have that $\boldsymbol{\beta}Q,\boldsymbol{\beta}_{\perp}]$ is a (full rank) orthogonal matrix. Since
$\boldsymbol{\beta}^{\mathsf{T}}(I_{kp}+\boldsymbol{\Pi})\boldsymbol{\beta}_{\perp}=\boldsymbol{\beta}^{\mathsf{T}}(I_{kp}+\boldsymbol{\alpha}\boldsymbol{\beta}^{\mathsf{T}})\boldsymbol{\beta}_{\perp}=\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{\beta}_{\perp}=0$
and, noting $\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{\beta}Q=\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{\beta}(\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{\beta})^{-1/2}=(\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{\beta})^{-1/2}=Q^{-1}$,
$Q\boldsymbol{\beta}^{\mathsf{T}}(I_{kp}+\boldsymbol{\Pi})\boldsymbol{\beta}Q$ $=Q\boldsymbol{\beta}^{\mathsf{T}}(I_{kp}+\boldsymbol{\alpha}\boldsymbol{\beta}^{\mathsf{T}})\boldsymbol{\beta}Q$
$=Q(I_{kp-q}+\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{\alpha})\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{\beta}Q=Q(I_{kp-q}+\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{\alpha})Q^{-1}.$
It follows that
$\begin{bmatrix}\boldsymbol{\beta}_{\perp}^{\mathsf{T}}\\ Q\boldsymbol{\beta}^{\mathsf{T}}\end{bmatrix}[I_{kp}+\boldsymbol{\Pi}]\begin{bmatrix}\boldsymbol{\beta}_{\perp}&\boldsymbol{\beta}Q\end{bmatrix}=\begin{bmatrix}I_{q}&\boldsymbol{\beta}_{\perp}^{\mathsf{T}}\boldsymbol{\Pi}\boldsymbol{\beta}Q\\ 0&Q(I_{kp-q}+\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{\alpha})Q^{-1}\end{bmatrix}$.
Hence the eigenvalues of $I_{kp-q}+\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{\alpha}$ are the kpqkp-q eigenvalues of $I_{kp}+\boldsymbol{\Pi}$ that are strictly inside the unit circle. It then follows by Lemma A.1 in Hansen (2005) that $\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{\alpha}$ and $\boldsymbol{\alpha}_{\perp}^{\mathsf{T}}\boldsymbol{\beta}_{\perp}=\alpha_{\perp}^{\mathsf{T}}\Gamma(1)\beta_{\perp}$ have full rank. ∎
Appendix D Transposition to the structural form
Here we verify the claims made in part (i) of each of Remarks 4.14.3, and complete the proof of Theorem 4.4. We suppose that DGP holds, so that $(y_{t},x_{t})$ are generated by the CKSVAR
$\phi^{+}(L)y_{t}^{+}+\phi^{-}(L)y_{t}^{-}+\Phi^{x}(L)x_{t}=c+u_{t}.$ (D.1)
By Proposition 2.1, if $(\tilde{y}_{t},\tilde{x}_{t})$ are constructed from $(y_{t},x_{t})$ via (2.13), then $(\tilde{y}_{t},\tilde{x}_{t})$ follow a canonical CKSVAR, which we denote here as
$\tilde{\phi}^{+}(L)\tilde{y}_{t}^{+}+\tilde{\phi}^{-}(L)\tilde{y}_{t}^{-}+\tilde{\Phi}^{x}(L)\tilde{x}_{t}=\tilde{c}+\tilde{u}_{t}.$ (D.2)
We use tildes to distinguish the series and parameters of the canonical form (D.1), from the original CKSVAR (D.2) from which they were derived; the mapping between these is given in (2.13)–(2.15).
Our approach here is as follows. We show that if DGP, CVAR, DET (or DET) and a suitably modified form of CO(i) (resp. CO(ii), CO(iii)) hold for $(y_{t},x_{t})$, then DGP, CVAR, DET (or DET) and CO(i) (resp. CO(ii), CO(iii)) also hold for $(\tilde{y}_{t},\tilde{x}_{t})$. Since ERR (or ERR) trivially carry over from $u_{t}$ to $\tilde{u}_{t}$, this permits Theorems 4.14.3 to be applied to $(\tilde{y}_{t},\tilde{x}_{t})$: their implications for $(y_{t},x_{t})$ are then derived by inverting the mapping (2.13)–(2.15).
Preliminary to these arguments, we define
$(P^{\pm})^{-1}\coloneqq\begin{bmatrix}\bar{\phi}_{0,yy}^{\pm}&0\\ \phi_{0,xy}^{\pm}&\Phi_{0,xx}\end{bmatrix}$, (D.3)
and prove the following auxiliary result.
Proposition D.1.
Suppose DGP holds. Then
  1. (i)
    $\tilde{\Phi}^{\pm}(\lambda)=Q\Phi^{\pm}(\lambda)P^{\pm}$;
  2. (ii)
    the roots of $\det\tilde{\Phi}^{(i)}(\lambda)$ and $\det\Phi^{(i)}(\lambda)$ are identical, for $(i)\in\{+,-\}$;
  3. (iii)
    $\Pi^{\pm}=Q^{-1}\tilde{\Pi}^{\pm}(P^{\pm})^{-1}$ and $\Pi^{x}=Q^{-1}\tilde{\Pi}^{x}\Phi_{0,xx}$; and
  4. (iv)
    $c\in\operatorname{sp}\Pi^{+}\cap\operatorname{sp}\Pi^{-}$ if and only if $\tilde{c}=Qc\in\operatorname{sp}\tilde{\Pi}^{+}\cap\tilde{\Pi}^{-}$.
Proof.
By Proposition 2.1, a CKSVAR satisfying DGP has a canonical form, which relates to the original model via (2.13)–(2.14). Recall the definition of $(P^{\pm})^{-1}$ given in (D.3). Since
$P$ $=\begin{bmatrix}(\bar{\phi}_{0,yy}^{+})^{-1}&0&0\\ 0&(\bar{\phi}_{0,yy}^{-})^{-1}&0\\ -\Phi_{0,xx}^{-1}\phi_{0,xy}^{+}(\bar{\phi}_{0,yy}^{+})^{-1}&-\Phi_{0,xx}^{-1}\phi_{0,xy}^{-}(\bar{\phi}_{0,yy}^{-})^{-1}&\Phi_{0,xx}^{-1}\end{bmatrix}$
and
$P^{\pm}=\begin{bmatrix}(\bar{\phi}_{0,yy}^{\pm})^{-1}&0\\ -\Phi_{0,xx}^{-1}\phi_{0,xy}^{\pm}(\bar{\phi}_{0,yy}^{\pm})^{-1}&\Phi_{0,xx}^{-1}\end{bmatrix}$, (D.4)
as may be verified using the partitioned inverse formula, (2.14) may be equivalently stated as
$\tilde{\Phi}^{\pm}(\lambda)=Q\Phi^{\pm}(\lambda)P^{\pm},$ (D.5)
and thus (i) and (ii) hold. Since $\Pi^{\pm}=-\Phi^{\pm}(1)$ and $\tilde{\Pi}^{\pm}=-\tilde{\Phi}^{\pm}(1)$, (iii) follows immediately upon taking $\lambda=1$ in (D.5), and the fact that
$\Pi^{\pm}=Q^{-1}\begin{bmatrix}\tilde{\pi}^{\pm}&\tilde{\Pi}^{x}\end{bmatrix}(P^{\pm})^{-1}=Q^{-1}\begin{bmatrix}\tilde{\pi}^{\pm}&\tilde{\Pi}^{x}\end{bmatrix}\begin{bmatrix}\bar{\phi}_{0,yy}^{\pm}&0\\ \phi_{0,xy}^{\pm}&\Phi_{0,xx}\end{bmatrix}=\begin{bmatrix}\pi^{\pm}&Q^{-1}\tilde{\Pi}^{x}\Phi_{0xx}\end{bmatrix}$.
Finally, (iv) follows from $\operatorname{sp}\tilde{\Pi}^{+}\cap\tilde{\Pi}^{-}=\operatorname{sp}(Q\Pi^{+})\cap(Q\Pi^{-})$. ∎
We suppose henceforth that $(y_{t},x_{t})$ satisfies DGP and CVAR, and for the moment also DET. Then $(\tilde{y}_{t},\tilde{x}_{t})$ satisfies DGP, CVAR and DET by Propositions 2.1 and D.1. We now consider cases (i)–(iii) in turn, before turning to the case where DET is relaxed to DET.
(i). Suppose $(y_{t},x_{t})$ additionally satisfies ERR and CO(i), except with CO(i).2 replaced by (ii). We must verify that $(\tilde{y}_{t},\tilde{x}_{t})$ satisfies CO(i). $(\tilde{y}_{t},\tilde{x}_{t})$ satisfies CO(i).1 by Proposition D.1(iii); CO(i).2 since $(y_{t},x_{t})$ satisfies (ii); and CO(i).4 via Proposition 2.1.
It remains therefore to verify CO(i).3. We note that as a further consequence of Proposition D.1(iii),
$\tilde{\alpha}^{+}\tilde{\beta}^{+\mathsf{T}}=\tilde{\Pi}^{+}=Q\Pi^{+}P^{+}=Q\alpha^{+}\beta^{+\mathsf{T}}P^{+}$ (D.6)
so that $\tilde{\alpha}^{+}=Q\alpha^{+}$, $\tilde{\beta}^{+}=(P^{+})^{\mathsf{T}}\beta^{+}$, and we may take $\tilde{\alpha}_{\perp}^{+}=(Q^{-1})^{\mathsf{T}}\alpha_{\perp}^{+}$ and $\tilde{\beta}_{\perp}^{+}=(P^{+})^{-1}\beta_{\perp}^{+}$. Further, $\tilde{\Gamma}^{+}(\lambda)=Q\Gamma^{+}(\lambda)P^{+}$, and so
$\tilde{P}_{\beta_{\perp}^{+}}$ $=\tilde{\beta}_{\perp}^{+}[\tilde{\alpha}_{\perp}^{+\mathsf{T}}\tilde{\Gamma}^{+}(1)\tilde{\beta}_{\perp}^{+}]^{-1}\tilde{\alpha}_{\perp}^{+\mathsf{T}}=(P^{+})^{-1}\beta_{\perp}^{+}[\alpha_{\perp}^{+\mathsf{T}}\Gamma^{+}(1)\beta_{\perp}^{+}]^{-1}\alpha_{\perp}^{+\mathsf{T}}Q^{-1}.$
$=(P^{+})^{-1}P_{\beta_{\perp}^{+}}Q^{-1}.$ (D.7)
Hence, using Proposition D.1(iii) again,
$\tilde{\kappa}=\tilde{P}_{\beta_{\perp}^{+}}\tilde{\pi}^{-}=\tilde{P}_{\beta_{\perp}^{+}}\tilde{\Pi}^{-}e_{1}$ $=(P^{+})^{-1}P_{\beta_{\perp}^{+}}\Pi^{-}P^{-}e_{1},$
where
$\Pi^{-}P^{-}e_{1}$ $=\begin{bmatrix}\pi^{-}&\Pi^{x}\end{bmatrix}\begin{bmatrix}(\bar{\phi}_{0,yy}^{-})^{-1}&0\\ -\Phi_{0,xx}^{-1}\phi_{0,xy}^{-}(\bar{\phi}_{0,yy}^{-})^{-1}&\Phi_{0,xx}^{-1}\end{bmatrix}\begin{bmatrix}1\\ 0\end{bmatrix}$
$=(\bar{\phi}_{0,yy}^{-})^{-1}[\pi^{-}-\Pi^{x}\Phi_{0,xx}^{-1}\phi_{0,xy}^{-}],$
and hence, using that $P_{\beta_{\perp}^{+}}\Pi^{x}=0$,
$\tilde{\kappa}=(\bar{\phi}_{0,yy}^{-})^{-1}(P^{+})^{-1}P_{\beta_{\perp}^{+}}[\pi^{-}-\Pi^{x}\Phi_{0,xx}^{-1}\phi_{0,xy}^{-}]=(\bar{\phi}_{0,yy}^{-})^{-1}(P^{+})^{-1}\kappa$ (D.8)
where $\kappa=P_{\beta_{\perp}^{+}}\pi^{-}$. Since $e_{1}^{\mathsf{T}}(P^{+})^{-1}=\bar{\phi}_{0,yy}^{+}e_{1}^{\mathsf{T}}$, it follows that
$\tilde{\kappa}_{1}=e_{1}^{\mathsf{T}}\tilde{\kappa}=(\bar{\phi}_{0,yy}^{-})^{-1}\bar{\phi}_{0,yy}^{+}\kappa_{1},$ (D.9)
and so $\operatorname{sgn}\tilde{\kappa}_{1}=\operatorname{sgn}\kappa_{1}$, since $\bar{\phi}_{0,yy}^{\pm}>0$ by DGP.3. Thus $(\tilde{y}_{t},\tilde{x}_{t})$ satisfies CO(i).3, since $(y_{t},x_{t})$ does.
It follows that Theorem 4.1 applies to $(\tilde{y}_{t},\tilde{x}_{t})$. Regarding the conclusions of that theorem, note that in this case $\tilde{{\cal Y}}_{0}$ and ${\cal Y}_{0}$ must be non-negative. Hence $\tilde{{\cal Z}}_{0}=(P^{+})^{-1}{\cal Z}_{0}$ by (2.13), and so by Proposition 2.1
$\tilde{U}_{0}(\lambda)=\tilde{\Gamma}^{+}(1)\tilde{{\cal Z}}_{0}+\tilde{U}(\lambda)=Q[\Gamma^{+}(1){\cal Z}_{0}+U(\lambda)]=QU_{0}(\lambda)$
whence $P_{\tilde{\beta}_{\perp}^{+}}\tilde{U}_{0}(\lambda)=(P^{+})^{-1}P_{\beta_{\perp}^{+}}U_{0}(\lambda)$. Since $\tilde{Y}=\tilde{Y}^{+}$ in this case, we have
$Z(\lambda)=P^{+}\tilde{Z}(\lambda)$ $=P^{+}\left\{P_{\tilde{\beta}_{\perp}^{+}}\tilde{U}_{0}(\lambda)+\tilde{\kappa}_{1}^{-1}\tilde{\kappa}\sup_{\lambda^{\prime}\leq\lambda}[-e_{1}^{\mathsf{T}}P_{\tilde{\beta}_{\perp}^{+}}\tilde{U}_{0}(\lambda)]_{+}\right\}$
$=_{(1)}P_{\beta_{\perp}^{+}}U_{0}(\lambda)+(\bar{\phi}_{0,yy}^{-})^{-1}\tilde{\kappa}_{1}^{-1}\kappa\sup_{\lambda^{\prime}\leq\lambda}[-e_{1}^{\mathsf{T}}(P^{+})^{-1}P_{\beta_{\perp}^{+}}U_{0}(\lambda)]_{+}$
$=_{(2)}P_{\beta_{\perp}^{+}}U_{0}(\lambda)+\kappa_{1}^{-1}\kappa\sup_{\lambda^{\prime}\leq\lambda}[-e_{1}^{\mathsf{T}}P_{\beta_{\perp}^{+}}U_{0}(\lambda)]_{+},$ (D.10)
where $_{(1)}$ follows by (D.8), and $_{(2)}$ by (D.9) and $e_{1}^{\mathsf{T}}(P^{+})^{-1}=\bar{\phi}_{0,yy}^{+}e_{1}^{\mathsf{T}}$. It follows immediately that $\beta^{+\mathsf{T}}z_{t}$ and $y_{t}^{-}$ are $I^{\ast}(0)$. Since $\tilde{y}_{t}^{-}\sim I^{\ast}(0)$, it follows that $\Delta\tilde{y}_{t}^{-}\sim I^{\ast}(0)$, and therefore so too is $\Delta\tilde{y}_{t}^{+}=\Delta\tilde{y}_{t}-\Delta\tilde{y}_{t}^{-}$. Hence
$\begin{bmatrix}\Delta y_{t}^{+}\\ \Delta y_{t}^{-}\\ \Delta x_{t}\end{bmatrix}=P\begin{bmatrix}\Delta\tilde{y}_{t}^{+}\\ \Delta\tilde{y}_{t}^{-}\\ \Delta\tilde{x}_{t}\end{bmatrix}\sim I^{\ast}(0)$ (D.11)
implies that $\Delta z_{t}\sim I^{\ast}(0)$. Thus the conclusions of the theorem hold also for $(y_{t},x_{t})$, exactly as stated.
(ii). Suppose $(y_{t},x_{t})$ additionally satisfies ERR and CO(ii), except with CO(ii).2 replaced by (ii). Analogously to case (i), $(\tilde{y}_{t},\tilde{x}_{t})$ satisfies CO(ii).1 by Proposition D.1(iii); CO(ii).2 since $(y_{t},x_{t})$ satisfies (ii); and CO(ii).4 via Proposition 2.1.
Regarding CO(ii).3, we first note that similarly to (D.6),
$\tilde{\alpha}\tilde{\beta}^{\pm\mathsf{T}}=\tilde{\Pi}^{\pm}=Q\Pi^{\pm}P^{\pm}=Q\alpha\beta^{\pm\mathsf{T}}P^{\pm}$
so that $\tilde{\alpha}=Q\alpha$ and $\tilde{\beta}^{\pm}=(P^{\pm})^{\mathsf{T}}\beta^{\pm}$, and we may take $\tilde{\alpha}_{\perp}=(Q^{-1})^{\mathsf{T}}\alpha_{\perp}$. With respect to $\tilde{\beta}_{\perp}^{\pm}=\tilde{\beta}_{\perp}(\pm 1)$, for the purposes of verifying CO(ii).3, we need to choose this such that its first row is equal to $e_{q,1}^{\mathsf{T}}$, and its final columns are regime-invariant, as per (4.8) (see Remark 4.24.2). This may be engineered by defining
$P(y)$ $\coloneqq P^{+}\mathbf{1}^{+}(y)+P^{-}\mathbf{1}^{-}(y)$ $\bar{\phi}_{0,yy}(y)$ $\coloneqq\bar{\phi}_{0,yy}^{+}\mathbf{1}^{+}(y)+\bar{\phi}_{0,yy}^{-}P^{-}\mathbf{1}^{-}(y)$ (D.12)
and $M(y)=\operatorname{diag}\{\bar{\phi}_{0,yy}^{-1}(y),I_{p-1}\}$, and then taking
$\tilde{\beta}_{\perp}(y)\coloneqq P(y)^{-1}\beta_{\perp}(y)M(y)$ $=\begin{bmatrix}\bar{\phi}_{0,yy}(y)&0\\ \phi_{0,xy}(y)&\Phi_{0,xx}\end{bmatrix}\begin{bmatrix}1&0\\ -\theta(y)&\beta_{x,\perp}\end{bmatrix}\begin{bmatrix}\bar{\phi}_{0,yy}^{-1}(y)&0\\ 0&I_{p-1}\end{bmatrix}$
$\eqqcolon\begin{bmatrix}1\\ -\tilde{\theta}(y)&\tilde{\beta}_{x,\perp}\end{bmatrix}$
where $\tilde{\beta}_{x,\perp}=\Phi_{0,xx}\beta_{x,\perp}$. Since $\tilde{\Gamma}(1;y)=Q\Gamma(1;y)P(y)$ by Proposition D.1(iii), we have
$\det\tilde{\alpha}_{\perp}^{\mathsf{T}}\tilde{\Gamma}(1;y)\tilde{\beta}_{\perp}(y)=\det\alpha_{\perp}^{\mathsf{T}}\Gamma(1;y)\beta_{\perp}(y)\det M(y).$
Noting that $\operatorname{sgn}\det M(y)=\operatorname{sgn}\bar{\phi}_{0,yy}(y)>0$, it follows that so that $(\tilde{y}_{t},\tilde{x}_{t})$ satisfies CO(ii).3, since $(y_{t},x_{t})$ does.
Hence Theorem 4.2 applies to $(\tilde{y}_{t},\tilde{x}_{t})$. Regarding the conclusions of that theorem, we first note that $\operatorname{sgn}y_{t}=\operatorname{sgn}\tilde{y}_{t}$, $\operatorname{sgn}{\cal Y}_{0}=\operatorname{sgn}{\cal\tilde{Y}}_{0}$,
$\tilde{U}_{0}(\lambda)=\tilde{\Gamma}(1;{\cal\tilde{Y}}_{0})\tilde{{\cal Z}}_{0}+\tilde{U}(\lambda)=Q[\Gamma(1;{\cal Y}_{0}){\cal Z}_{0}+U(\lambda)],$
and
$P_{\tilde{\beta}_{\perp}}(y)=\tilde{\beta}_{\perp}(y)[\tilde{\alpha}_{\perp}^{\mathsf{T}}\tilde{\Gamma}(1;y)\tilde{\beta}_{\perp}(y)]^{-1}\tilde{\alpha}_{\perp}^{\mathsf{T}}=P(y)^{-1}P_{\beta_{\perp}}(y)Q^{-1}.$
Hence
$Z(\lambda)=P[\tilde{Y}(\lambda)]\tilde{Z}(\lambda)$ $=P[\tilde{Y}(\lambda)]P_{\tilde{\beta}_{\perp}}[\tilde{Y}(\lambda)]\tilde{U}_{0}(\lambda)=P_{\beta_{\perp}}[Y(\lambda)]U_{0}(\lambda).$ (D.13)
Further,
$\beta(y_{t})^{\mathsf{T}}z_{t}=\beta(y_{t})^{\mathsf{T}}z_{t}=\beta(y_{t})^{\mathsf{T}}P(\tilde{y}_{t})\tilde{z}_{t}=\tilde{\beta}(\tilde{y}_{t})^{\mathsf{T}}\tilde{z}_{t}\sim I^{\ast}(0).$
Since $\tilde{y}_{t}\sim I^{\ast}(1)$, it follows that $\tilde{y}_{t}^{\pm}\sim I^{\ast}(1)$, and hence $\Delta\tilde{y}_{t}^{\pm}\sim I^{\ast}(0)$. That $\Delta z_{t}\sim I^{\ast}(0)$ then follows as in (D.11). Deduce that the conclusions of the theorem hold also for $(y_{t},x_{t})$.
(iii). Suppose $(y_{t},x_{t})$ additionally satisfies ERR and CO(iii), except with CO(iii).2 replaced by (ii). Since $(\tilde{y}_{t},\tilde{x}_{t})$ satisfies CO(iii).1 by Proposition D.1(iii), it follows from Theorem 4.3 that $\tilde{\boldsymbol{\chi}}_{t}\coloneqq(\tilde{\chi}_{t}^{\mathsf{T}},\ldots,\tilde{\chi}_{t-k+1}^{\mathsf{T}})^{\mathsf{T}}$ is $\cal Q$}-geometrically ergodic.
Recalling the definitions in (D.12) above, we may rewrite (2.13) as
$\begin{bmatrix}\tilde{y}_{t}\\ \tilde{x}_{t}\end{bmatrix}=\tilde{z}_{t}=P(y_{t})^{-1}z_{t}=\begin{bmatrix}\bar{\phi}_{0,yy}(y_{t})&0\\ \phi_{0,xy}(y_{t})&\Phi_{0,xx}\end{bmatrix}\begin{bmatrix}y_{t}\\ x_{t}\end{bmatrix}$ (D.14)
and using the fact that $\operatorname{sgn}y_{t}=\operatorname{sgn}\tilde{y}_{t}$,
$\begin{bmatrix}y_{t}\\ x_{t}\end{bmatrix}=\begin{bmatrix}\bar{\phi}_{0,yy}^{-1}(\tilde{y}_{t})&0\\ \bar{\phi}_{0,xy}(\tilde{y}_{t})&\Phi_{0,xx}^{-1}\end{bmatrix}\begin{bmatrix}\tilde{y}_{t}\\ \tilde{x}_{t}\end{bmatrix}$ (D.15)
where $\bar{\phi}_{0,xy}(y)\coloneqq-\Phi_{0,xx}^{-1}\phi_{0,xy}(y)\bar{\phi}_{0,yy}^{-1}(y)$. Further, we have by Proposition D.1(iii) that
$\tilde{\alpha}_{x}\tilde{\beta}_{x}^{\mathsf{T}}=\tilde{\Pi}^{x}=Q\Pi^{x}\Phi_{0,xx}^{-1}=Q\alpha_{x}\beta_{x}^{\mathsf{T}}\Phi_{0,xx}^{-1}.$ (D.16)
Hence by (D.14) and (D.16)
$\begin{bmatrix}\tilde{y}_{t}\\ \tilde{\beta}_{x}^{\mathsf{T}}\tilde{x}_{t}\end{bmatrix}=\tilde{\beta}^{\mathsf{T}}\tilde{z}_{t}=\begin{bmatrix}1&0\\ 0&\tilde{\beta}_{x}^{\mathsf{T}}\end{bmatrix}P(y_{t})^{-1}z_{t}=\begin{bmatrix}1&0\\ 0&\beta_{x}^{\mathsf{T}}\Phi_{0,xx}^{-1}\end{bmatrix}\begin{bmatrix}\bar{\phi}_{0,yy}(y_{t})&0\\ \phi_{0,xy}(y_{t})&\Phi_{0,xx}\end{bmatrix}\begin{bmatrix}y_{t}\\ x_{t}\end{bmatrix}\\ =\begin{bmatrix}\bar{\phi}_{0,yy}(y_{t})&0\\ \beta_{x}^{\mathsf{T}}\Phi_{0,xx}^{-1}\phi_{0,xy}(y_{t})&I_{r-1}\end{bmatrix}\begin{bmatrix}y_{t}\\ \beta_{x}^{\mathsf{T}}x_{t}\end{bmatrix}$,
and so, since $\operatorname{sgn}y_{t}=\operatorname{sgn}\tilde{y}_{t}$,
$\beta^{\mathsf{T}}z_{t}=\begin{bmatrix}y_{t}\\ \beta_{x}^{\mathsf{T}}x_{t}\end{bmatrix}=\begin{bmatrix}\bar{\phi}_{0,yy}^{-1}(\tilde{y}_{t})&0\\ -\beta_{x}^{\mathsf{T}}\bar{\phi}_{0,xy}(\tilde{y}_{t})&I_{r-1}\end{bmatrix}\begin{bmatrix}\tilde{y}_{t}\\ \tilde{\beta}_{x}^{\mathsf{T}}\tilde{x}_{t}\end{bmatrix}$
so that the l.h.s. is a Lipschitz continuous function of $\tilde{\beta}^{\mathsf{T}}\tilde{z}_{t}$, as claimed. Further, we have from (D.15) that
$\Delta x_{t}=\Delta[\bar{\phi}_{0,xy}(\tilde{y}_{t})\tilde{y}_{t}]+\Phi_{0,xx}^{-1}\Delta\tilde{x}_{t}$
where $\tilde{y}_{t}\mapsto\bar{\phi}_{0,xy}(\tilde{y}_{t})\tilde{y}_{t}$ is a Lipschitz continuous function of (the first element of) $\tilde{\chi}_{t}$, and since
$\Delta\tilde{x}_{t}=(\overline{\beta}_{x,\perp}\tilde{\beta}_{x,\perp}^{\mathsf{T}}+\overline{\beta}_{x}\tilde{\beta}_{x}^{\mathsf{T}})\Delta\tilde{x}_{t}=\overline{\beta}_{x,\perp}(\tilde{\beta}_{x,\perp}^{\mathsf{T}}\Delta\tilde{x}_{t})+\overline{\beta}_{x}\Delta(\tilde{\beta}_{x}^{\mathsf{T}}\tilde{x}_{t})$
for $\overline{\beta}_{x,\perp}\coloneqq\tilde{\beta}_{x,\perp}(\tilde{\beta}_{x,\perp}^{\mathsf{T}}\tilde{\beta}_{x,\perp})^{-1}$ and $\overline{\beta}_{x}\coloneqq\tilde{\beta}_{x}(\tilde{\beta}_{x}^{\mathsf{T}}\tilde{\beta}_{x})^{-1}$, it follows that $\Delta\tilde{x}_{t}$ is a linear function of $(\tilde{\chi}_{t},\tilde{\chi}_{t-1})$. Thus $\Delta x_{t}$ is a Lipschitz continuous function of $(\tilde{\chi}_{t},\tilde{\chi}_{t-1})$.
Deterministics in cases (i) and (ii). Now suppose that $z_{t}=(y_{t},x_{t}^{\mathsf{T}})^{\mathsf{T}}$ satisfies DET rather than DET. It follows from (D.7) that $e_{1}^{\mathsf{T}}\tilde{P}_{\beta_{\perp}^{+}}\tilde{c}=0$; thus DET also holds for the derived canonical process $\tilde{z}_{t}=(\tilde{y}_{t},\tilde{x}_{t}^{\mathsf{T}})^{\mathsf{T}}$. We then have from the proof of Theorem 4.4 (under DGP) that the conclusions of part (ii) of Theorems 4.1 and 4.2 apply to the detrended canonical processes
$\begin{bmatrix}\tilde{y}_{t}^{d}\\ \tilde{x}_{t}\end{bmatrix}=\tilde{z}_{t}^{d}=\tilde{z}_{t}-(\tilde{P}_{\beta_{\perp}^{+}}\tilde{c})t,$
where $\tilde{y}_{t}^{d}=\tilde{y}_{t}$ because $e_{1}^{\mathsf{T}}\tilde{P}_{\beta_{\perp}^{+}}c=0$. Using $P(y)$ as defined in (D.12), and noting that $\operatorname{sgn}y_{t}=\operatorname{sgn}\tilde{y}_{t}$, we may write
$z_{t}=P(\tilde{y}_{t})\tilde{z}_{t}$ (D.17)
or equivalently $P(y_{t})^{-1}z_{t}=\tilde{z}_{t}$, and so
$\tilde{z}_{t}^{d}=\tilde{z}_{t}-(\tilde{P}_{\beta_{\perp}^{+}}\tilde{c})t$ $=P(y_{t})^{-1}z_{t}-[(P^{+})^{-1}P_{\beta_{\perp}^{+}}c]t$
$=_{(1)}P(y_{t})^{-1}[z_{t}-(P_{\beta_{\perp}^{+}}c)t]=P(y_{t})^{-1}z_{t}^{d}$
where $_{(1)}$ holds because only the first column of $P(y)$ depends on $y$, and $e_{1}^{\mathsf{T}}P_{\beta_{\perp}^{+}}c=0$. Hence
$z_{t}^{d}=P(\tilde{y}_{t}^{d})\tilde{z}_{t}^{d},$ (D.18)
where $\tilde{y}_{t}=\tilde{y}_{t}^{d}$. Because the mapping between the detrended processes in (D.18) is identical to that for the original processes in (D.17), the same arguments as given in cases (i) and (ii) immediately above will now transpose the conclusions of Theorem 4.4 from $\tilde{z}_{t}^{d}$ to $z_{t}^{d}$.
Appendix E Verification of Remarks 4.14.1 and 4.14.1
We first note the following corollary to Lemma A.3. We say a VAR is stationary if all its autoregressive roots lie outside the unit circle (i.e. irrespective of whether the series is given a stationary initialisation).
Lemma E.1.
Suppose that the assumptions of Lemma A.3 hold, and $\{z_{t}\}$ is generated according to
$z_{t}=c+\Phi(L)z_{t-1}+u_{t}.$
Let $\xi_{t}\coloneqq\beta^{\mathsf{T}}z_{t}$. Then $\boldsymbol{\xi}_{t}\coloneqq(\xi_{t}^{\mathsf{T}},\Delta z_{t}^{\mathsf{T}},\ldots,\Delta z_{t-(k-2)}^{\mathsf{T}})^{\mathsf{T}}$ follows the stationary VAR given by
$\boldsymbol{\xi}_{t}=\boldsymbol{\beta}_{1:p}^{\mathsf{T}}c+(I_{p(k-1)+r}+\boldsymbol{\beta}^{\mathsf{T}}\boldsymbol{\alpha})\boldsymbol{\xi}_{t-1}+\boldsymbol{\beta}_{1:p}^{\mathsf{T}}u_{t}.$
Proof.
The claim follows from part (ii) of Lemma A.3, and premultiplying the companion form $\Delta\boldsymbol{z}_{t}=\boldsymbol{c}+\boldsymbol{\Pi}\boldsymbol{z}_{t-1}+\boldsymbol{u}_{t}$ by $\boldsymbol{\beta}^{\mathsf{T}}$, where $\boldsymbol{z}_{t}=(z_{t}^{\mathsf{T}},z_{t-1}^{\mathsf{T}},\ldots,z_{t-k+1}^{\mathsf{T}})^{\mathsf{T}}$, $\boldsymbol{c}=(c^{\mathsf{T}},0_{p(k-1)}^{\mathsf{T}})^{\mathsf{T}}$ and $\boldsymbol{u}_{t}=(u_{t}^{\mathsf{T}},0_{p(k-1)}^{\mathsf{T}})^{\mathsf{T}}$. ∎
Proof of Remark 4.14.1.
Rather than working with the matrices $\boldsymbol{F}_{0}$ and $\boldsymbol{F}_{1}$ directly, it is easier if we consider the autoregressive systems that they correspond to, as given in (B.7), recognising that those systems are dynamically stable if and only if the eigenvalues of their companion form matrices lie strictly inside the unit circle (e.g. Lütkepohl, 2007, Sec. 2.1.1).
Let $t\in\mathbb{N}$ be given. We first consider $\boldsymbol{F}_{0}$: this is operative if $\delta_{s}=0$ for all $s\in\{t-k,\ldots,t-1\}$. In this case, $\text{$\overline{y}$}_{s}=y_{s}^{-}=0$ for all such $s$, and (B.7) reduces to
$\boldsymbol{\xi}_{t}^{+}=\boldsymbol{\beta}_{1:p}^{+\mathsf{T}}c+(I+\boldsymbol{\beta}^{+\mathsf{T}}\boldsymbol{\alpha}^{+})\boldsymbol{\xi}_{t-1}^{+}+\boldsymbol{\beta}_{1:p}^{+\mathsf{T}}u_{t},$
By Lemma A.3, the eigenvalues of $I+\boldsymbol{\beta}^{+\mathsf{T}}\boldsymbol{\alpha}^{+}$ lie inside the unit circle; hence so too do those of $\boldsymbol{F}_{0}$. (Indeed, it may be verified that the only nonzero eigenvalues of $\boldsymbol{F}_{0}$ are those of $I+\boldsymbol{\beta}^{+\mathsf{T}}\boldsymbol{\alpha}^{+}$.)
We next consider $\boldsymbol{F}_{1}$, which is operative if $\delta_{s}=1$ for all $s\in\{t-k,\ldots,t-1\}$. Then $y_{s}=\text{$\overline{y}$}_{s}=y_{s}^{-}<0$ for all such $s$, and (B.7) becomes
$\boldsymbol{\xi}_{t}^{+}$ $=\boldsymbol{\beta}_{1:p}^{+\mathsf{T}}c+(I+\boldsymbol{\beta}^{+\mathsf{T}}\boldsymbol{\alpha}^{+})\boldsymbol{\xi}_{t-1}^{+}$ (E.1a)
$\qquad\qquad+\boldsymbol{\beta}_{1:p}^{+\mathsf{T}}(\phi_{1}^{-}-\phi_{1}^{+})y_{t-1}+\boldsymbol{\beta}_{1:p}^{+\mathsf{T}}\sum_{i=2}^{k}(\phi_{i}^{-}-\phi_{i}^{+})y_{t-i}+\boldsymbol{\beta}_{1:p}^{+\mathsf{T}}u_{t}$
$y_{t}$ $=e_{1}^{\mathsf{T}}c+\boldsymbol{e}_{1}^{\mathsf{T}}\boldsymbol{\alpha}^{+}\boldsymbol{\xi}_{t-1}^{+}$
$\qquad\qquad+[1+e_{1}^{\mathsf{T}}(\phi_{1}^{-}-\phi_{1}^{+})]y_{t-1}+\sum_{i=2}^{k}e_{1}^{\mathsf{T}}(\phi_{i}^{-}-\phi_{i}^{+})y_{t-i}+e_{1}^{\mathsf{T}}u_{t}.$ (E.1b)
Now the preceding must agree with the $(\boldsymbol{\xi}_{t}^{+},y_{t})$ generated by the VAR when $y_{s}<0$ for all $s\in\{t-k,\ldots,t-1\}$, i.e. generated according to
$z_{t}=\Phi^{-}(L)z_{t-1}+u_{t}.$
Under CVAR and CO(i).1, the preceding is a linear cointegrated VAR with
$\Pi^{-}=\begin{bmatrix}\pi^{-}&\Pi^{x}\end{bmatrix}=\Pi^{+}+\begin{bmatrix}\pi^{-}-\pi^{+}&0\end{bmatrix}=\begin{bmatrix}\alpha^{+}&0\\ 0&\pi^{-}-\pi^{+}\end{bmatrix}\begin{bmatrix}\beta^{+\mathsf{T}}\\ e_{1}^{\mathsf{T}}\end{bmatrix}\eqqcolon\alpha^{-}\beta^{-\mathsf{T}},$
and equilibrium errors $\xi_{t}^{-}\coloneqq\beta^{-\mathsf{T}}z_{t}=[\begin{smallmatrix}\beta^{+\mathsf{T}}z_{t}\\ y_{t}\end{smallmatrix}]=[\begin{smallmatrix}\xi_{t}^{+}\\ y_{t}\end{smallmatrix}]$. It follows from Lemma E.1 that
$\boldsymbol{\xi}_{t}^{-}\coloneqq(\xi_{t}^{-\mathsf{T}},\Delta z_{t}^{\mathsf{T}},\ldots,\Delta z_{t-(k-2)}^{\mathsf{T}})^{\mathsf{T}}=(\xi_{t}^{+\mathsf{T}},y_{t},\Delta z_{t}^{\mathsf{T}},\ldots,\Delta z_{t-(k-2)}^{\mathsf{T}})^{\mathsf{T}}$
evolves according to a stationary VAR. Since $(\boldsymbol{\xi}_{t}^{+},y_{t})$ is simply a reordering of $\boldsymbol{\xi}_{t}^{-}$, it follows that the system described by (E.1), i.e. by $\boldsymbol{F}_{1}$, must also be stationary. Hence $\boldsymbol{F}_{1}$ must have all its eigenvalues strictly inside the unit circle. ∎
Proof of Remark 4.14.1.
Suppose $k=1$. In this case, $\pi^{-}-\pi^{+}=\phi_{1}^{-}-\phi_{1}^{+}$, and so by Remark 4.14.1, the matrix
$\boldsymbol{F}_{1}=\begin{bmatrix}I_{r}+\beta^{+\mathsf{T}}\alpha^{+}&\beta^{+\mathsf{T}}(\pi^{-}-\pi^{+})\\ e_{1}^{\mathsf{T}}\alpha^{+}&1+e_{1}^{\mathsf{T}}(\pi^{-}-\pi^{+})\end{bmatrix}$
must have all its eigenvalues inside the unit circle; hence $\det(I_{r+1}-\boldsymbol{F}_{1})=\prod_{i=1}^{r+1}(1-\rho_{i}(\boldsymbol{F}_{1}))>0$, where $\rho_{i}(\boldsymbol{F}_{1})$ denotes the $i$th eigenvalue of $\boldsymbol{F}_{1}$. Note that
$e_{1}^{\mathsf{T}}(\pi^{-}-\pi^{+})-e_{1}^{\mathsf{T}}\alpha^{+}(\beta^{+\mathsf{T}}\alpha^{+})^{-1}\beta^{+\mathsf{T}}(\pi^{-}-\pi^{+})$ $=e_{1}^{\mathsf{T}}[I_{p}-\alpha^{+}(\beta^{+\mathsf{T}}\alpha^{+})^{-1}\beta^{+\mathsf{T}}](\pi^{-}-\pi^{+})$
$=e_{1}^{\mathsf{T}}\beta_{\perp}^{+}(\alpha_{\perp}^{+\mathsf{T}}\beta_{\perp}^{+})^{-1}\alpha_{\perp}^{+\mathsf{T}}(\pi^{-}-\pi^{+})$
$=e_{1}^{\mathsf{T}}P_{\beta_{\perp}^{+}}\pi^{-}=\kappa_{1},$
where the penultimate equality holds since $\alpha_{\perp}^{+}\pi^{+}=0$, and $\Gamma^{+}(1)=I_{p}$ when $k=1$. Hence
$\begin{bmatrix}I_{r}&0\\ -e_{1}^{\mathsf{T}}\alpha^{+}(\beta^{+\mathsf{T}}\alpha^{+})^{-1}&1\end{bmatrix}(\boldsymbol{F}_{1}-I_{r+1})=\begin{bmatrix}\beta^{+\mathsf{T}}\alpha^{+}&\beta^{+\mathsf{T}}(\pi^{-}-\pi^{+})\\ 0&\kappa_{1}\end{bmatrix}$,
from which it follows that
$(-1)^{r+1}\det(I_{r+1}-\boldsymbol{F}_{1})$ $=\det(\boldsymbol{F}_{1}-I_{r+1})$
$=\kappa_{1}\det(\beta^{+\mathsf{T}}\alpha^{+})=\kappa_{1}(-1)^{r}\det[I_{r}-(I_{r}+\beta^{+\mathsf{T}}\alpha^{+})].$
By Lemma A.3, $I_{r}+\beta^{+\mathsf{T}}\alpha^{+}$ has all its eigenvalues inside the unit circle, and so $\det[I_{r}-(I_{r}+\beta^{+\mathsf{T}}\alpha^{+})]>0$. Hence $\kappa_{1}<0$.
Next suppose $p=1$, while allowing for general $k\in\mathbb{N}$. Then
$\kappa_{1}=\kappa=P_{\beta_{\perp}^{+}}\pi^{-}=-\gamma^{+}(1)^{-1}\phi^{-}(1)$
where $\gamma^{+}(1)=1-\sum_{i=1}^{k-1}\gamma_{i}^{+}$ and $\phi^{-}(1)=1-\sum_{i=1}^{k}\phi_{i}^{-}$. Both $\gamma^{+}(\lambda)$ and $\phi^{-}(\lambda)$ have all their roots outside the unit circle, and hence both $\gamma^{+}(1)$ and $\phi^{-}(1)$ are strictly positive. It follows that $\kappa_{1}<0$ as required. ∎
Appendix F Detailed calculations for Example 4.2
Here we provide some further details of the calculations underlying Example 4.2. For convenience, we reproduce the system (4.12) in CKSVAR form as
$\begin{bmatrix}1\\ -\delta^{+}\\ 0\end{bmatrix}\bar{\pi}_{t}^{+}+\begin{bmatrix}1\\ -\delta^{-}\\ 0\end{bmatrix}\bar{\pi}_{t}^{-}+\begin{bmatrix}0&0\\ 1&0\\ 0&1\end{bmatrix}\begin{bmatrix}\Delta\bar{y}_{t}\\ g_{t}\end{bmatrix}\\ =\begin{bmatrix}1\\ -\delta^{+}\\ 0\end{bmatrix}\bar{\pi}_{t-1}^{+}+\begin{bmatrix}1\\ -\delta^{-}\\ 0\end{bmatrix}\bar{\pi}_{t-1}^{-}+\begin{bmatrix}0&0\\ 0&1\\ 0&1\end{bmatrix}\begin{bmatrix}\Delta\bar{y}_{t-1}\\ g_{t-1}\end{bmatrix}+\begin{bmatrix}u_{t}^{\pi}\\ u_{t}^{y}\\ u_{t}^{g}\end{bmatrix}$. (F.1)
We thus have first-order ($k=1$) CKSVAR with
$\Phi_{0}^{\pm}$ $=\begin{bmatrix}1&0&0\\ -\delta^{\pm}&1&0\\ 0&0&1\end{bmatrix}$ $\Phi_{1}^{\pm}$ $=\begin{bmatrix}1&0&0\\ -\delta^{\pm}&0&1\\ 0&0&1\end{bmatrix}$. (F.2)
Hence
$\Pi^{\pm}=\begin{bmatrix}1&0&0\\ -\delta^{\pm}&0&1\\ 0&0&1\end{bmatrix}-\begin{bmatrix}1&0&0\\ -\delta^{\pm}&1&0\\ 0&0&1\end{bmatrix}=\begin{bmatrix}0&0&0\\ 0&-1&1\\ 0&0&0\end{bmatrix}=\begin{bmatrix}0\\ 1\\ 0\end{bmatrix}\begin{bmatrix}0&-1&1\end{bmatrix}=\alpha\beta^{\pm\mathsf{T}}$ (F.3)
and thus $\Pi^{\pm}$ and $\beta^{\pm}=(0,-1,1)^{\mathsf{T}}$ are regime-invariant. $\Pi^{x}$ is obtained from the final p1=2p-1=2 columns of $\Pi^{+}$ (or $\Pi^{-}$), and so
$\Pi^{x}=\begin{bmatrix}0&0\\ -1&1\\ 0&0\end{bmatrix}$
whence $\operatorname{rk}\Pi^{\pm}=\operatorname{rk}\Pi^{x}=1$, showing that CO(ii).1 holds with $r=1$.
In view of (F.3), we can construct orthocomplement matrices as
$\alpha_{\perp}$ $=\begin{bmatrix}1&0\\ 0&0\\ 0&1\end{bmatrix}$ $\beta_{\perp}^{\pm}$ $=\begin{bmatrix}1&0\\ 0&1\\ 0&1\end{bmatrix}$,
where it will be observed that the $\beta_{\perp}^{\pm}$ indeed have the form required by (4.8). Since $\Gamma_{i}=0$ for all $i\geq 1$ in this first-order model, we have $\Gamma^{\pm}(1)=\Phi_{0}^{\pm}$, and thus from computing
$\alpha_{\perp}^{\mathsf{T}}\Gamma^{\pm}(1)\beta_{\perp}^{\pm}=\begin{bmatrix}1&0&0\\ 0&0&1\end{bmatrix}\begin{bmatrix}1&0&0\\ -\delta^{\pm}&1&0\\ 0&0&1\end{bmatrix}\begin{bmatrix}1&0\\ 0&1\\ 0&1\end{bmatrix}=\begin{bmatrix}1&0\\ 0&1\end{bmatrix}=I_{2}$
we see immediately that $\det\alpha_{\perp}^{\mathsf{T}}\Gamma^{+}(1)\beta_{\perp}^{+}=1=\det\alpha_{\perp}^{\mathsf{T}}\Gamma^{-}(1)\beta_{\perp}^{-}$, whence CO(ii).3 holds.
The final condition to verify is CO(ii).2 – or rather (ii), because the CSKVAR (F.1) is in structural form. We must therefore compute $\tilde{\Pi}^{\pm}$ for the associated canonical form, the parameters can be obtained via the mapping (2.13)–(2.15). From (F.2) we have (noting that here $y_{t}=\bar{\pi}_{t}$ and $x_{t}=(\Delta y_{t},g_{t})^{\mathsf{T}}$),
$\phi_{0,yy}^{\pm}$ $=1$ $\phi_{0,xy}^{\pm}$ $=\begin{bmatrix}-\delta^{\pm}\\ 0\end{bmatrix}$ $\phi_{0,yx}^{\mathsf{T}}$ $=\begin{bmatrix}0&0\end{bmatrix}$ $\Phi_{0,xx}=$ $\begin{bmatrix}1&0\\ 0&1\end{bmatrix}$
whence
$\bar{\phi}_{0,yy}^{\pm}=\phi_{0,yy}^{\pm}-\phi_{0,yx}^{\mathsf{T}}\Phi_{0,xx}^{-1}\phi_{0,xy}^{\pm}=1$
and so
$P^{-1}=\begin{bmatrix}\bar{\phi}_{0,yy}^{+}&0&0\\ 0&\bar{\phi}_{0,yy}^{-}&0\\ \phi_{0,xy}^{+}&\phi_{0,xy}^{-}&\Phi_{0,xx}\end{bmatrix}=\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ -\delta^{+}&-\delta^{-}&1&0\\ 0&0&0&1\end{bmatrix}$
and
$Q=\begin{bmatrix}1&-\phi_{0,yx}^{\mathsf{T}}\Phi_{0,xx}^{-1}\\ 0&I_{p-1}\end{bmatrix}=\begin{bmatrix}1&0\\ 0&I_{2}\end{bmatrix}=I_{3}.$
Hence, $\tilde{\Pi}=[\tilde{\pi}^{+},\tilde{\pi}^{-},\Pi^{x}]=-\tilde{\Phi}(1)$ is given by
$\tilde{\Pi}=Q\begin{bmatrix}\pi^{+}&\pi^{-}&\Pi^{x}\end{bmatrix}P=\begin{bmatrix}0&0&0&0\\ 0&0&-1&1\\ 0&0&0&0\end{bmatrix}\begin{bmatrix}1&0&0&0\\ 0&1&0&0\\ \delta^{+}&\delta^{-}&1&0\\ 0&0&0&1\end{bmatrix}=\begin{bmatrix}0&0&0&0\\ -\delta^{+}&-\delta^{-}&-1&1\\ 0&0&0&0\end{bmatrix}$.
It follows that the matrices $\tilde{\Pi}^{\pm}=[\tilde{\pi}^{\pm},\Pi^{x}]$ are regime dependent, and factorise as
$\tilde{\Pi}^{\pm}=\begin{bmatrix}0&0&0\\ -\delta^{\pm}&-1&1\\ 0&0&0\end{bmatrix}=\begin{bmatrix}0\\ 1\\ 0\end{bmatrix}\begin{bmatrix}-\delta^{\pm}&-1&1\end{bmatrix}\eqqcolon\tilde{\alpha}\tilde{\beta}^{\pm\mathsf{T}}$
whence
$I_{r}+\tilde{\beta}^{\pm\mathsf{T}}\tilde{\alpha}=1+\begin{bmatrix}-\delta^{\pm}&-1&1\end{bmatrix}\begin{bmatrix}0\\ 1\\ 0\end{bmatrix}=0$
and thus (ii) is satisfied.